### v3

# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

Abstract
We propose a general approach to optimize noise schedules for training and sampling in diffusion models. Our approach optimizes the noise schedules to ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process. Any distance metric for measuring the probability-distributional change is applicable to our approach, and we introduce three distance metrics. We evaluated the effectiveness of our approach on unconditional and class-conditional image-generation tasks using the LSUN (Horse, Bedroom, Church), ImageNet, FFHQ, and CIFAR10 datasets. Through extensive experiments, we confirmed that our approach broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations ranging from 5 to 250. Notably, by using our approach for optimizing both training and sampling schedules, we achieved a state-of-the-art FID score of 2.03 without sacrificing mode coverage on LSUN Horse 256 $\times$ 256.
1 Introduction
Diffusion models are a type of probabilistic generative model that have gained significant attention for their expressive ability to generate high-quality samples across various domains, such as image generation image_generation1 ; image_generation2 ; image_generation3 , image super-resolution image_super_resolution1 ; image_super_resolution2 , video synthesis video_generation1 ; video_generation2 , and audio generation audio_generation1 ; audio_generation2 .
Diffusion models were originally proposed in the previous study nonequiliburium_thermodynamics , and denoising diffusion probabilistic models (DDPMs) ddpm_original demonstrate that diffusion models can achieve state-of-the-art performance in image generation. In diffusion models, two diffusion processes are defined using Markov chains, referred to as the forward process and reverse process. In the forward process, data are diffused by gradually adding Gaussian noise, until the probability distribution of diffused data becomes a Gaussian distribution. In the reverse process, data are generated from Gaussian noise by tracing the forward process in the reverse direction. Noise levels added to data at each step of the diffusion processes are hyperparameters and referred to as the noise schedule. We do not necessarily have to use a same noise schedule for training and sampling, and recent studies edm ; vdmpp have shown that using different noise schedules can lead to improved performance. Although diffusion models tend to perform well in terms of sample fidelity and mode coverage, they face the challenge of longer generation times due to the iterative function evaluations during sampling diffusion_gans .
To improve the performance of diffusion models, we propose general approach for optimizing training and sampling schedules that ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process. We call our approach constant rate scheduling (CRS). As previously stated, the reverse process needs to accurately trace the forward process in the reverse direction to restore the data from pure noise. This motivates us to determine noise schedules that can enhance the traceability of the forward process. Figure 1 illustrates a toy example of diffused data distributions with three data points in one-dimensional data space (see Appendix A for details). Here, $\alpha$ represents the decay rate of data at each step of the diffusion process. In this example, the diffusion process can be divided into three regions. The first is an area where the probability distributions barely change, and the noise level can be changed rapidly ($\alpha\lesssim 0.6$). The second is an area where three modes corresponding to the data points emerge, and careful adjustment of the noise level is necessary to enhance the mode coverage ($0.6\lesssim\alpha\lesssim 0.97$). The third is an area where the peaks of each mode become distinct, and the noise level should be changed slowly for sample fidelity ($\alpha\gtrsim 0.97$). As evident from this example, the diffusion process exhibits regions with rapid and slow changes in the probability distributions. We assume that the probability-distributional change of diffused data represents the traceability of the diffusion process and determines the noise schedule to ensure a constant rate of change in the probability distributions.
CRS can be considered general from the following perspectives: 1) CRS can optimize both the training and sampling schedules, 2) any distance metric for measuring probability-distributional change is applicable, allowing the introduction of domain specific distance metrics and the use of different noise schedules for training and sampling, 3) any noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
Our contributions are summarized as follows:
  1. 1.
    We propose a general approach to optimize noise schedules for both training and sampling. We can use any distance metric for measuring probability-distributional change and introduce three distance metrics. One is specialized for image generation, while the other two are expected to be applicable across a wide range of domains beyond image generation.
  2. 2.
    Through extensive experiments, we confirm that CRS broadly improves the performance of image generation in pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations (NFEs) from 5 to 250. CRS contributes to the performance improvement at both training and sampling, achieving a state-of-the-art Fréchet inception distance (FID) fid_score score of 2.03 on LSUN Horse 256 $\times$ 256 dataset. CRS is fundamentally different from conventional approaches, such as align your steps (AYS) noise_ays and geometry-inspired time scheduling (GITS) noise_gits , in that it can optimize the training schedule as well.
Figure 1: Toy example of diffused data distributions with three data points in one-dimensional data space. Probability distributions barely change when $\alpha\lesssim 0.6$, and we can rapidly change noise level. Three modes corresponding to data points emerge when $0.6\lesssim\alpha\lesssim 0.97$, and noise level should be changed slowly for mode coverage. Three modes become distinct when $\alpha\gtrsim 0.97$, requiring careful control of noise level for sample fidelity.
2 Related Work
Recent studies regarding fast sampling can be mainly categorized into the following three strategies.
Samplers: This strategy devises update rules of diffused data in the reverse process. Denoising diffusion implicit models (DDIMs) ddim_original introduce a non-Markov chain to generalize DDPMs and propose a deterministic sampler. In the continuous-time diffusion models, the forward and reverse processes are formulated using a stochastic differential equation (SDE) and ordinary differential equation (ODE) score_matching3 . Numerous studies have been conducted to develop efficient solvers tailored for the reverse SDE and ODE in diffusion models sampler_adaptive_step ; sampler_pndm ; sampler_dpm_solver ; sampler_dpmpp ; sampler_ipndm ; sampler_unipc ; sampler_dpm-v3 ; sampler_amed .
Noise schedules: Noise schedules significantly affect the performance of diffusion models noise_effect , and many noise schedules have been proposed ddpm_original ; improved_diffusion ; noise_effect ; shifted_cosine . However, the functional form in these studies, such as the linear ddim_original and cosine improved_diffusion functions, is determined somewhat arbitrarily. Recently, systematic optimization of noise schedules have begun to be actively studied noise_gits ; noise_jys . The most relevant approach to CRS is AYS noise_ays , which determines a noise schedule for sampling to minimize the discretization errors when solving the reverse SDE. However, noise schedules aimed at improving the training efficiency have been largely overlooked compared with the sampling efficiency, though several studies, such as elucidating diffusion models (EDMs) edm and variational diffusion models++ (VDM++) vdmpp , have investigated them.
Distillations: The distillation-based acceleration is one of the most important direction for fast sampling distill_org1 ; distill_org2 ; distill_consistency ; distill_dmd ; distill_dmd2 ; distill_s4s . CRS is expected to provide better teacher models for distillation methods because it can optimize the training schedule as well.
3 Background
We briefly introduce DDPMs.
In DDPMs ddpm_original , the forward process, which starts from a data point $\bm{x}_{0}\sim q(\bm{x})$, is defined as
$q(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})\right)q(\bm{x}_{0})\ \ \mathrm{and}\ \ q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\beta_{t}\bm{x}_{t-1},\delta_{t}^{2}\bm{I}),$ (1)
where $\beta_{t}$ is the pre-defined decay rate of the data, $\delta_{t}$ represents the noise strength added to $\bm{x}_{t-1}$, and $T$ is the number of timesteps in the forward process. It is straightforward to verify that
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I}),$ (2)
where $\alpha_{t}=\beta_{t}\alpha_{t-1}$ and $\sigma_{t}^{2}=\delta_{t}^{2}+\beta_{t}^{2}\sigma_{t-1}^{2}$. We focus on the variance-preserving process, where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ and $\beta_{t}^{2}+\delta_{t}^{2}=1$ are satisfied.
The reverse process starts from Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. In the limit of $T\to\infty$, the probability distribution of the reverse process is well approximated by the Gaussian distribution ddpm_approx :
$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\nu_{t}^{2}\bm{I}).$ (3)
Using this approximation, the noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},t)$:
$\bm{\mu}_{\theta}(\bm{x},t)=\frac{1}{\beta_{t}}\left(\bm{x}_{t}-\frac{\delta_{t}^{2}}{\sigma_{t}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)\right),$ (4)
is optimized using a simplified loss function derived from the variational lower bound.
4 General Approach for Optimizing Noise Schedule
We explain our motivation and present CRS in more detail.
4.1 Motivation
As stated in Sec. 3, when $T$ is sufficiently large, it is a good approximation to assume the probability distribution of the reverse process $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ as a Gaussian distribution, as expressed in Eq. (3). In other words, the distributional change between adjacent timesteps must be small enough to be well approximated by a Gaussian distribution. This observation motivates us to minimize the maximum distributional change in the diffusion process for holding the approximation in Eq. (3) throughout the diffusion process.
We can understand the importance of the probability-distributional change in the diffusion process from the perspective of noise conditional score networks (NCSNs) score_matching1 ; score_matching2 ; score_matching3 , which are also the origin of diffusion models. According to the manifold hypothesis, natural data typically lie on a low-dimensional manifold embedded within a high-dimensional data space. Based on this property, there exist two main challenges in generating samples using only the score of the data distribution $\nabla_{\bm{x}}\log q(\bm{x})$ starting from a random initialization $\bm{x}_{\mathrm{rnd}}$ score_matching1 . First, $\bm{x}_{\mathrm{rnd}}$ is likely to be located in regions with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, where it is difficult to learn the score accurately. Second, even if the score can be learned correctly, the score tends to be small in the region with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, leading to slow mixing. To address these challenges, NCSNs interpolate between the data distribution and a Gaussian distribution, the support of which is the whole space. If the distributional change between adjacent timesteps is sufficiently small during the interpolation process, a sample generated at a certain timestep is likely to fall into a region where the probability density of the adjacent timestep is high. Therefore, by minimizing the maximum distributional change between adjacent timesteps, it becomes possible to generate samples using the reliable score at all timesteps of the interpolation process.
4.2 Constant Rate Scheduling
As described above, our criterion to optimize noise schedules is to minimize the maximum distance between consecutive probability distributions in the diffusion process. Using the distance $D(t,t+\Delta t)$ between the probability distributions of diffused data $p(\bm{x}_{t})$ and $p(\bm{x}_{t+\Delta t})$, our criterion is expressed as
$\underset{\alpha(t)}{\operatorname{argmin}}\left(\max_{t}D(t,t+\Delta t)\right).$ (5)
This optimization problem has a straightforward solution as follows:
$D(t,t+\Delta t)=\frac{D(t,t+\Delta t)-D(t,t)}{\Delta t}\Delta t\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=\mathrm{const.},$ (6)
where we use $D(t,t)=0$.
To explicitly show the dependence on the noise schedule, we rewrite $D(t,t^{\prime})$ as $\tilde{D}(\alpha(t),\alpha(t^{\prime}))\equiv\tilde{D}(\alpha,\alpha^{\prime})$. Here, $\alpha(t)$ is the noise-schedule function to be optimized. This results in
$D(t,t+\Delta t)\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=-v(\alpha)\frac{d\alpha(t)}{dt}\Delta t=\mathrm{const.},$ (7)
where $v(\alpha)=\left.\frac{\partial\tilde{D}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}$ represents the rate of change in the probability distributions of diffused data. Therefore, the noise-schedule function is determined to satisfy $-\frac{d\alpha(t)}{dt}\propto v(\alpha)^{-1}$. The proportionality coefficient can be calculated using the boundary conditions: $\alpha(0)=\alpha_{\max}$ and $\alpha(1)=\alpha_{\min}$. Finally, we obtain the noise-schedule function as follows:
[…]
Here, we introduce a hyperparameter $\xi>0$ to control the dependence of $\alpha(t)$ on $v(\alpha)$. As $\xi$ is increased, more timesteps are allocated to the region where $v(\alpha)$ is large.
Similar to NCSN (score_matching1, ) inspired by simulated annealing (sa_original, ), our formulation draws inspiration from the adiabatic theorem (adiabatic_theorem, ) in quantum annealing (qa_original, ). Quantum annealing is an algorithm for solving combinatorial optimization problems, where the target cost function to be optimized is interpolated with an initial cost function with a trivial optimal solution. The initial state of the optimization is set to the optimal solution of the initial cost function. By tracing the optimal solutions of the instantaneous cost functions during the interpolation, the optimal solution of the target cost function is then obtained. In quantum annealing, accurately tracing the optimal solution of an instantaneous cost function is crucial, and the adiabatic theorem derives the traceability of the optimal solution at each step. Let $\alpha(t)$ be the weight of the target cost function and $1-\alpha(t)$ be the weight of the initial cost function. In this case, the inverse traceability can be expressed in the form of $\frac{d\alpha(t)}{dt}v(\alpha)$, which is the same form as Eq. (7). It has also been verified that an optimal functional form of $\alpha(t)$ is obtained by minimizing the maximum value of the inverse traceability grover_optimal , which ensures a consistent traceability throughout the interpolation process. In CRS, we regard the probability-distributional change as the inverse traceability and minimize the maximum distance for determining the noise schedule.
4.3 Distance Metrics
Any distance metric is applicable to CRS. We introduce three distance metrics.
Distance metric based on FID: FID is a widely used distance metric to evaluate the sample fidelity in the field of image generation. The FID calculates the Fréchet distance fid_score under a normality assumption in the vision-relevant feature space embedded using the Inception-V3 model inception_v3 , which is trained on the ImageNet classification. When FID is used as the distance metric for the pixel-space diffusion models, CRS allocates more timesteps to the region where the sample fidelity significantly deteriorates in the diffusion process ($\alpha\simeq 1$). A weakness of FID-based distance metric is that it cannot accurately capture the probability-distributional change in the region with high-noise levels ($\alpha\simeq 0$), because the feature model is not trained on diffused images.
It is important to note that, in the latent-space diffusion models, the probability-distributional change must be measured in the latent space embedded using the autoencoder (the approximation in Eq. (3) is applied to the diffusion process defined in the latent space). We trained ResNet50 resnet on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model for computing FID.
We refer to $v(\alpha)$ computed using FID as $v_{\mathrm{FID}}(\alpha)$. Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain the noise schedule using Eq. (8). For the detailed computation procedure of $v_{\mathrm{FID}}(\alpha)$, refer to Appendix C.1.
Distance metric based on data prediction: We introduce a distance metric that uses a trained diffusion model as follows:
$v_{x}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{x}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].\$ (9)
Here, $\bm{x}_{\alpha}$ and $\bm{x}_{\alpha^{\prime}}$ are sampled from $q(\bm{x}_{\alpha},\bm{x}_{\alpha}^{\prime})=q(\bm{x}_{\alpha}^{\prime}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})$, and $\bm{x}_{\theta}(\bm{x},\alpha)$ is the data prediction, which is related to the noise prediction $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ as follows: $\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$. This distance metric aims at allocating more computational resources to the region where the data prediction changes rapidly. As shown in Appendix B, we can interpret this distance metric as a simplified version of the Kullback-Leibler divergence upper bound (KLUB) introduced in AYS. This distance metric can be used not only for determining the sampling schedule after training a diffusion model, but also for adaptively optimizing the training schedule during the training process. Pseudocodes for optimizing training and sampling schedules using $v_{x}(\alpha)$ are shown in Appendix C.2.
We can interpret that $v_{x}(\alpha)$ measures the probability-distributional change of diffused data. The data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ is given by the weighted sum of the data points $\bm{x}_{0}$ statistical_mechanics . The weight of each data point depends on the noise level $\alpha$. The weights of all data points are uniform when $\alpha=0$, while a small number of data points contribute to $\bm{x}_{\theta}(\bm{x},\alpha)$ when $\alpha\simeq 1$. The changes in these weights represent how the trajectories in the diffusion process gradually diverge as the noise level decreases. Therefore, we can interpret $v_{x}(\alpha)$ as measuring the probability-distributional change of diffused data.
Unlike the FID-based distance metric, $v_{x}(\alpha)$ can capture distributional changes even in regions with high-noise levels. However, $v_{x}(\alpha)$, which is related to the KLUB, is not necessarily optimal for improving sample fidelity at sampling. Therefore, as discussed in the next subsection, combining multiple distance metrics in a complementary manner enhances the performance of CRS.
Distance metric based on noise prediction: In the experiments presented in Sec. 5, we trained the noise prediction model. In this case, it is also considered effective to use noise predictions to measure the probability-distributional change as follows:
$v_{\varepsilon}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (10)
Hereafter, we refer to CRS using $v_{\mathrm{FID}}(\alpha)$, $v_{x}(\alpha)$, and $v_{\varepsilon}(\alpha)$ for optimizing noise schedules as CRS-$v_{\mathrm{FID}}$, CRS-$v_{x}$, and CRS-$v_{\varepsilon}$, respectively.
4.4 Combining Multiple Distance Metrics
As explained in the previous subsection, each distance metric has its own strengths an weaknesses. Therefore, it is considered effective to complementarily combining multiple distance metrics for optimizing noise schedules. We introduce a strategy for combining them as follows:
$v(\alpha)=\sum_{m=1}^{M}w_{m}\tilde{v}_{m}(\alpha),\ \ \mathrm{where}\ \ \tilde{v}_{m}(\alpha)=\frac{1}{C_{m}}v_{m}(\alpha)^{\xi_{m}}\ \ \mathrm{and}\ \ C_{m}=\int_{\alpha_{\min}}^{\alpha_{\max}}v_{m}(\alpha)^{\xi_{m}}.$ (11)
Here, $M$ is the number of distance metrics to be combined, $v_{m}(\alpha)$ is computed for each distance metric, $w_{m}$ is the weight for $v_{m}(\alpha)$ satisfying $\sum_{m=1}^{M}w_{m}=1$, and $\xi_{m}$ controls the dependence on $v_{m}(\alpha)$. Note that the value range of $v_{m}(\alpha)$ can significantly vary depending on the distance metric. Therefore, normalization of $v_{m}(\alpha)$ is required before combining them, as shown in Eq. (11).
Using this formulation, it is also possible to combine conventional noise schedules with CRS based noise schedules by converting $\alpha(t)$ to $v(\alpha)$. For example, when $\alpha(t)=\cos\left(\frac{\pi t}{2}\right)$, $v(\alpha)$ is obtained as follows:
$v(\alpha)\propto-\left(\frac{d\alpha(t)}{dt}\right)^{-1}=\frac{2}{\pi}\sin^{-1}\left(\frac{\pi t}{2}\right)=\frac{2}{\pi\sqrt{1-\alpha^{2}}}.$ (12)
This result suggests that every noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
5 Experiments
We experimentally demonstrated that CRS broadly improves the performance of diffusion models. Below, we provide a brief explanation of each component of the diffusion model.
Dataset: We use six image datasets: LSUN (Church, Bedroom, Horse) dataset_lsun , ImageNet dataset_imagenet , FFHQ dataset_ffhq , and CIFAR10 dataset_cifar10 . LSUN and FFHQ are used for unconditional image generation, and ImageNet and CIFAR10 for class-conditional image generation.
Noise Prediction Model: We use the U-Net architecture introduced in the ablated diffusion model (ADM) adm as the noise prediction model (see Table 8 in Appendix E for hyperparameter settings).
Training Schedule: We use linear ddpm_original , shifted cosine shifted_cosine , adaptive noise schedule proposed in VDM++ vdmpp , and CRS.
Loss Function: Except for the adaptive noise schedule of VDM++, we use the following simplified loss function introduced in DDPMs:
$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},t\sim\mathcal{U}(0,1),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x},\alpha(t))-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (13)
where $\mathcal{D}$ denotes all data points in the training dataset, $\alpha(t)$ is the training schedule, and $\bm{x}=\alpha(t)\bm{x}_{0}+\sqrt{1-\alpha(t)^{2}}\bm{\varepsilon}$.
When using the adaptive noise schedule of VDM++, the following loss function is applied:
$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\lambda\sim p(\lambda),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\frac{\omega(\lambda)}{p(\lambda)}\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (14)
where $\lambda=\log\left(\frac{\alpha^{2}}{\sigma^{2}}\right)$ is the log signal-to-noise ratio (log-SNR), $\omega(\lambda)$ is the weight for each noise level, and $p(\lambda)=-\frac{d\lambda}{dt}$ is the noise schedule for training. We use the EDM monotonic function as $\omega(\lambda)$, and $p(\lambda)$ is updated at each iteration to approximately satisfy the following condition:
$p(\lambda)\propto\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\omega(\lambda)\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right].$ (15)
Sampling Schedule: We use linear, shifted cosine, EDM sampling edm and CRS. In pixel-space diffusion models, we additionally evaluate CRS-$v_{x}+v_{\mathrm{FID}}$, which combines $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ for optimizing noise schedules. The reason for selecting the combination of $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ is that, as shown in Sec. 5.4, $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ have a complementary relationships. Recent studies edm ; vdmpp have shown that using different noise schedules for training and sampling can lead to improved performance. Therefore, in our experiments, we evaluate the performance of the diffusion models using arbitrary pairs of training and sampling schedules. Note that we can obtain different noise schedules for training and sampling using CRS, because it is capable of handling arbitrary distance metrics. Details of the conventional noise schedules used in our experiments can be found in Table 10 in Appendix E.
Sampling Methods: Images are mainly generated using two stochastic samplers and three deterministic samplers. For the stochastic samplers, we use Stochastic DDIM ($\eta=1$) ddim_original and SDE-DPM-Solver++(2M) sampler_dpmpp . For the deterministic samplers, we use DDIM ddim_original , DPM-Solver++(2M) sampler_dpmpp , and pseudo numerical methods for diffusion models (PNDMs) sampler_pndm . At small NFEs, we additionally use improved PNDM (iPNDM) sampler_ipndm and a unified predictor-corrector framework (UniPC) sampler_unipc .
Evaluation Metrics: We compute four metrics: FID fid_score , spatial FID (sFID) spatial_fid_score , and improved precision and recall recall_precision1 ; recall_precision2 . FID, sFID, and precision measure sample fidelity, while recall evaluates mode coverage.
5.1 Results on Pixel-Space Diffusion Models
We trained the continuous-time diffusion models on LSUH Horse 256 $\times$ 256 and Bedroom 256 $\times$ 256 for unconditional image generation. Detailed experimental results are listed in Appendix F.1.
LSUN Horse 256 $\times$ 256: The evaluation results are summarized in Table 2. The rows shaded in gray represent the dependence on the training schedule. Although sample fidelity and mode coverage are generally considered to be in a trade-off relationship, CRS-$v_{x}$ achieved the highest performance for all metrics. This result indicates that CRS-$v_{x}$ is not merely balancing trade-offs between evaluation metrics but fundamentally improves the quality of the generated samples.
The unshaded rows represent the dependence on sampling schedules. CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the highest performance for all metrics, indicating fundamental improvement in sample quality (see Appendix D for the hyperparameter setting of CRS-$v_{x}+v_{\mathrm{FID}}$). The FID score of 2.03 surpasses the current state-of-the-art FID score of 2.11 lsun_horse_sota .
LSUN Bedroom 256 $\times$ 256: We confirmed the advantage of CRS in the low-NFE regime by comparing it with the leading results reported in GITS noise_gits . We trained the continuous-time diffusion model and used CRS-$v_{x}$ for optimizing the training schedule. The results are listed in Table 2. The best FID scores at $\mathrm{NFE}=5$ and $\mathrm{NFE}=10$ were achieved when using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$ for optimizing the training and sampling schedules, respectively.
Table 1: Dependence on training and sampling schedules for LSUN Horse $256\times 256$ in pixel-space diffusion models at $\mathrm{NFE}=250$. SDE-DPM-Solver++(2M) was used as sampler. Bold values represent best training or sampling schedules for each metric.
Training Sampling Metrics
schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Linear Linear 2.90 6.82 0.66 0.56
Shifted cosine 2.95 6.43 0.67 0.56
VDM++ 3.82 7.40 0.66 0.54
CRS-$v_{\varepsilon}$ 2.91 6.56 0.66 0.56
CRS-$v_{x}$ 2.73 6.40 0.67 0.56
CRS-$v_{x}$ Linear 2.73 6.40 0.67 0.56
Shifted cosine 3.13 7.09 0.67 0.54
EDM 2.09 6.16 0.69 0.56
CRS-$v_{\varepsilon}$ 2.87 6.71 0.66 0.56
CRS-$v_{x}$ 5.46 8.34 0.62 0.52
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.03 6.06 0.69 0.56
Table 2: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion models at small NFEs. Bold values indicate best FID scores at each NFE. Results for rows other than "Ours" are taken from previously published study noise_gits .
Model Sampler Sampling schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$
EDM UniPC sampler_unipc EDM 23.34 5.75
AMED-Solver sampler_amed 5.38
iPNDM sampler_ipndm GITS noise_gits 15.85 5.28
Ours DPM-Solver++(2M) EDM 23.72 4.73
CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 4.88
UniPC EDM 82.52 4.94
CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 3.30
5.2 Results on Latent-Space Diffusion Models
We trained latent-space diffusion models using the pre-trained autoencoder VQ-4 used in latent diffusion models (LDMs) ldm . We used LSUN Church and Bedroom for unconditional image generation, and ImageNet for class-conditional image generation. The image resolution was set to $256\times 256$, which was reduced to $64\times 64$ in the latent space. We present the evaluation results on LSUN Church $256\times 256$. Detailed results are listed in Appendix F.2.
LSUN Church 256 $\times$ 256: The evaluation results are summarized in Table 4. The dependence on the training schedule is shown in the rows shaded in gray, and CRS-$v_{x}$ achieved the highest performance except for sFID. The unshaded rows represent the dependence on the sampling schedule, and CRS-$v_{x}$ further improves all evaluation metrics except for precision. These results indicate the superiority of CRS in latent-space diffusion models as well.
From the results in Table 4, CRS-$v_{x}$ appears to perform the best at both training and sampling; however, depending on the sampler and NFE, CRS-$v_{\varepsilon}$ is preferable at sampling.
5.3 Comparison with AYS Using EDM
Because $v_{x}(\alpha)$ is related to the KLUB introduced in AYS, we additionally compare CRS with AYS on FFHQ 64 $\times$ 64 using the pretrained model provided in EDMs edm . It is important to note that EDM adopts a variance-exploding process, while we have so far focused on the variance-preserving process. However, CRS can also be applied to the variance-exploding process by converting between $\alpha$ and the noise level in EDM $\sigma_{\mathrm{EDM}}$ via the log SNR, as shown below:
$\mathrm{logSNR}=\log\frac{\alpha^{2}}{1-\alpha^{2}}=\log\frac{1}{\sigma_{\mathrm{EDM}}^{2}}\Rightarrow\alpha=\frac{1}{\sqrt{1+\sigma_{\mathrm{EDM}}^{2}}}.$ (16)
The results are listed in Table 4. CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperformed AYS at all NFEs. Detailed results on pretrained EDM are listed in Appendix F.3.
Table 3: Dependence on training and sampling schedules for LSUN Church $256\times 256$ in latent-space diffusion models at $\mathrm{NFE}=30$. DPM-Solver++(2M) was used as sampler. Bold values represent best training or sampling schedules for each metric.
Training Sampling Metrics
schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Linear Linear 3.82 10.28 0.60 0.58
VDM++ 4.00 10.07 0.58 0.59
CRS-$v_{\mathrm{FID}}$ 3.78 10.51 0.60 0.58
CRS-$v_{\varepsilon}$ 3.87 10.30 0.60 0.59
CRS-$v_{x}$ 3.69 10.23 0.60 0.59
CRS-$v_{x}$ Linear 3.69 10.23 0.60 0.59
EDM 4.31 11.31 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.98 9.61 0.58 0.60
CRS-$v_{\varepsilon}$ 3.63 9.91 0.59 0.60
CRS-$v_{x}$ 3.59 9.51 0.59 0.60
Table 4: FID scores evaluated on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM. FID scores on AYS is taken from their original paper noise_ays .
Sampler Noise Schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) AYS noise_ays 5.43 3.29 2.62
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
5.4 Comparison of Noise Schedules
The CRS-based noise schedules in the pixel-space diffusion model are shown in Fig. 3. The hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$ was set to $w_{x}=w_{\mathrm{FID}}=0.5$ and $\xi_{x}=\xi_{\mathrm{FID}}=1.0$. Since $v_{\mathrm{FID}}(\alpha)$ takes large values in regions where sample fidelity degrades in the diffusion process, CRS-$v_{\mathrm{FID}}$ allocates many timesteps to the region where $\alpha\simeq 1$. In contrast, CRS-$v_{x}$ allocates more timesteps to regions with small $\alpha$. This is due to the fact that the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ can be expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$ using the noise prediction, where $\alpha$ appears in the denominator. Thus, CRS-$v_{\mathrm{FID}}$ and CRS-$v_{x}$ exhibits complementary characteristics, and combining them yields an effective sampling schedule. Since the noise schedule based on CRS-$v_{\varepsilon}$ can be approximately reproduced by tuning the hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$, we regard $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ as the fundamental compenents.
Figure 3 shows the CRS-based noise schedules in the latent-space diffusion model. The appropriate noise schedules differ significantly between the pixel-space and lantet-space diffusion models. The most notable difference is that, in the latent-space diffusion models, fewer timesteps are allocated to the region where $\alpha\simeq 1$. This can be attributed to the fact that, in latent-space diffusion models, the perceptual expression is largely generated by the autoencoder, and the autoencoder can absorb minor deviations in the generated latent variables. In contrast, the pixel-space diffusion models directly generate images, so careful control of the noise schedule is crucial at $\alpha\simeq 1$. We did not evaluate any combination of multiple distance metrics in latent-space diffusion models, because all the CRS-based noise schedules have a similar functional form.
Figure 2: Noise schedules on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model.
Figure 3: Noise schedules on LSUN Church 256 $\times$ 256 in latent-space diffusion model.
6 Limitations
This paper focuses on developing CRS and validating its effectiveness under fundamental experimental settings. However, to further strengthen the generality of CRS, additional investigations are needed from the following perspectives: 1) generality across various generation conditions, 2) scalability of the distance metrics with respect to dataset size and data dimensionality, 3) effectiveness in domains beyond image generation. While $v_{\mathrm{FID}}(\alpha)$ proved effective as the sampling schedule in pixel-space diffusion models, it is a distance metric specialized for image generation. Therefore, when applying CRS to other domains, it may be necessary to introduce new distance metrics that are better suited to the characteristics of the target domain.
7 Conclusion
We proposed a general approach called CRS to optimize noise schedules for training and sampling in diffusion models. Through extensive experiments, we confirmed that CRS broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and NFEs ranging from 5 to 250. CRS is fundamentally different from conventional approaches noise_ays ; noise_gits in that it can optimize the training schedule as well and use any distance metric for measuring probability-distributional change, allowing the introduction of domain specific distance metric and the use of different noise schedules for training and sampling. Although this study has the limitations discussed in Sec. 6, we believe that it successfully provides solid baselines. We hope that future research will further improve the generality of CRS, leading to its widespread adoption as a general-purpose approach.
Appendix A Probability Distribution of Diffused Data
We derive the probability distribution of diffused data used in the toy example (Fig. 1).
The probability distribution of diffused data is given by
[…]
$q(\bm{x}_{\alpha}|\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0},\sigma^{2}\bm{I}),$ (18)
where $q(\bm{x}_{0})$ is the probability distribution of a target dataset. We then approximate $q(\bm{x}_{0})$ with the empirical distribution as
$q(\bm{x}_{0})=\frac{1}{N}\sum_{n=1}^{N}\delta(\bm{x}_{0}-\bm{x}_{0}^{n}),$ (19)
where $N$ is the number of samples in the target dataset, $\bm{x}_{0}^{n}$ is the $n$-th sample, and $\delta$ denotes Dirac’s delta function. By substituting Eq. (19) into Eq. (17), we obtain
$q(\bm{x}_{\alpha})=\frac{1}{N}\sum_{n=1}^{N}\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0}^{n},\sigma^{2}\bm{I}).$ (20)
We used the following formula of Dirac’s delta function:
[…]
The resulting probability distribution is a Gaussian mixture distribution. We used Eq. (20) in the toy example.
Appendix B Derivation of Distance Metric Based on Data Prediction from KLUB
We show that $v_{x}(\alpha)$ can be derived as a simplified version of KLUB introduced in AYS.
B.1 KLUB in Variance-Preserving Process
The forward SDE of the continuous-time diffusion models [19] is defined as
$d\bm{x}=f(t)\bm{x}dt+g(t)d\bm{\omega},$ (22)
where $f(t)\bm{x}$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $\bm{\omega}$ is the standard Wiener process. The associated reverse SDE is given by
$d\bm{x}=\left[f(t)\bm{x}-g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\right]dt+g(t)d\bm{\omega}.$ (23)
For approximately computing the score $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, we train the following noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$:
$\bm{\varepsilon}_{\theta}(\bm{x},\alpha)\simeq-\sigma\nabla_{\bm{x}}\log p_{t}(\bm{x}).$ (24)
The noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ has the following relationship with the data prediction model $\bm{x}_{\theta}(\bm{x},\alpha)$:
$\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha).$ (25)
By substituting Eq. (24) into Eq. (23) and using Eq. (25), the reverse SDE can be rewritten as
$d\bm{x}$ $=\bm{f}(\bm{x},t)dt+g(t)d\bm{\omega},$ (26)
$\bm{f}(\bm{x},t)$ $=\left(f(t)+\frac{1}{\sigma(t)^{2}}g(t)^{2}\right)\bm{x}-\frac{\alpha(t)}{\sigma(t)^{2}}g(t)^{2}\bm{x}_{\theta}(\bm{x},\alpha(t)).$ (27)
Stochastic DDIM is a stochastic sampler designed to numerically solve the reverse SDE described above. In Stochastic DDIM, the solution for the first term of $\bm{f}(\bm{x},t)$ (linear term) is exactly computed, and discretization errors arise only from the second term of $\bm{f}(\bm{x},t)$ [22, 23]. According to a previous study [16], in the sub-interval $[t_{i-1},t_{i}]$, Stochastic DDIM solves the following discretized SDE:
$d\bm{x}$ $=\bm{f}^{\prime}(\bm{x},t)dt+g(t)d\bm{\omega},$ (28)
$\bm{f}^{\prime}(\bm{x},t)$ $=\left(f(t)+\frac{1}{\sigma(t)^{2}}g(t)^{2}\right)\bm{x}-\frac{\alpha(t)}{\sigma(t)^{2}}g(t)^{2}\bm{x}_{\theta}(\bm{x}_{t_{i}},\alpha(t_{i})).$ (29)
The KLUB between probability distributions of $\bm{x}_{t_{i-1}}$ obtained by exactly solving the reverse SDE and approximately solving the reverse SDE using Stochastic DDIM is introduced as
[…]
The diffusion coefficient $g(t)$ in the variance-preserving process is given by
$\frac{1}{2}g(t)^{2}=-\frac{\dot{\alpha}}{\alpha}.$ (32)
By substituting Eq. (32) into Eq. (31), we obtain
[…]
B.2 KLUB as Distance Metric of CRS
We transform the KLUB into a form that can be applied to CRS. Because the KLUB is used to evaluate the discretization error when solving the reverse SDE, it is reasonable to regard the KLUB as the traceability measure of the diffusion process.
First, we approximate the integral in the KLUB using Simpson’s rule as follows:
$\mathrm{KLUB}(t,t+\Delta t)$ $=\frac{\dot{\alpha(t)}\alpha(t)}{2\sigma(t)^{4}}\overline{\mathcal{D}_{x}^{2}}(t,t+\Delta t)\Delta t,$ (34)
$\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ $=\mathbb{E}_{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t^{\prime}}\sim q(\bm{x}_{t^{\prime}}|\bm{x}_{t})\end{subarray}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{t},\alpha(t))-\bm{x}(\bm{x}_{t^{\prime}},\alpha(t^{\prime}))\right\|_{2}^{2}\right].$ (35)
Next, to treat $\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ as a function of $\alpha$ rather than $t$, we introduce $\overline{D_{x}^{2}}(\alpha(t),\alpha(t^{\prime}))=\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$:
$\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (36)
Here, $q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ is given by
$q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ $=q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0}),$ (37)
$q(\bm{x}_{\alpha}|\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0},\sigma^{2}\bm{I}),$ (38)
$q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})$ $=\mathcal{N}(\bm{x}_{\alpha^{\prime}};\beta\bm{x}_{\alpha},\delta^{2}\bm{I}),$ (39)
where $\beta=\frac{\alpha^{\prime}}{\alpha}$, $\sigma=\sqrt{1-\alpha^{2}}$, and $\delta=\sqrt{1-\beta^{2}}$. Using $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$, we can rewrite $\mathrm{KLUB}(t,t+\Delta t)$ as
$\mathrm{KLUB}(t,t+\Delta t)$ $=\frac{\dot{\alpha}\alpha}{2\sigma^{4}}\overline{D_{x}^{2}}(\alpha,\alpha+\Delta\alpha)\Delta t,$ (40)
$=\frac{\dot{\alpha}\alpha}{2\sigma^{4}}\frac{\overline{D_{x}^{2}}(\alpha,\alpha+\Delta\alpha)-\overline{D_{x}^{2}}(\alpha,\alpha)}{\Delta\alpha}\frac{\Delta\alpha}{\Delta t}(\Delta t)^{2},$ (41)
$\simeq\left\{\frac{d\alpha}{dt}\cdot\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}\sqrt{\left.\frac{\partial\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\Delta t\right\}^{2}.$ (42)
We use $\overline{D_{x}^{2}}(\alpha,\alpha)=0$ in the second equality. Finally, by adopting $\sqrt{\mathrm{KLUB}(t,t^{\prime})}$ as the distance metric $D(t,t^{\prime})$ in CRS, we obtain the following equation:
$D(t,t+\Delta t)=\sqrt{\mathrm{KLUB}(t,t+\Delta t)}=-\frac{d\alpha}{dt}\cdot\underbrace{\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}\sqrt{\left.\frac{\partial\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}}_{v_{\mathrm{KLUB}}(\alpha)}\Delta t.$ (43)
which is the same form as Eq. (7). We can regard $v_{x}(\alpha)$ as a simplified version of $v_{\mathrm{KLUB}}(\alpha)$ in the sense that the weighting function of $v_{\mathrm{KLUB}}(\alpha)$, namely $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$, is discarded in $v_{x}(\alpha)$.
It is possible to optimize the noise schedule using $v_{\mathrm{KLUB}}(\alpha)$. However, when $v_{\mathrm{KLUB}}(\alpha)$ is directly used, the noise schedule is heavily impacted by the weighting function rather than $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$. The weighting function becomes extremely large around $\alpha\simeq 1(\sigma\simeq 0)$, which overwhelms the impact of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$. Recall that $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ can be interpreted as measuring probability-distributional changes of diffused data, as described in Sec. 4.3. Therefore, the simplified version we introduce is more aligned with our objective of optimizing the noise schedule on the basis of distributional changes.
Appendix C Implementation of CRS
We describe how to compute $v(\alpha)$ using each distance metric.
C.1 Distance Metric Based on FID
The forward process is simulated using the training dataset for computing $v_{\mathrm{FID}}(\alpha)$, as shown in Algorithm 1. One of the inputs $\{\alpha_{t}\}_{t=0,...,T}$ needs to be discretized to approximately represent the functional form of $v(\alpha)$ through linear interpolation. In our experiments, we set
$\alpha_{t}=1-\left(\frac{t}{T}\right)^{p},$ (44)
where $T=1000$, and set $p$ to $2$ and $1$ for the diffusion processes in pixel space and latent space, respectively. For the feature model $\bm{\phi}$, we used the Inception-V3 model trained on ImageNet classification for the pixel-space diffusion process. For the latent-space diffusion process, however, we need a feature model trained in the latent space embedded using the autoencoder. We trained ResNet50 on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model $\bm{\phi}$. Note that a trained diffusion model is not necessary for computing $v_{\mathrm{FID}}(\alpha)$. Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain noise schedules by solving Eq. (8).
The dependence of CRS-$v_{\mathrm{FID}}$ on hyperparameter $\xi$ is illustrated in Fig. 7. CRS-$v_{\mathrm{FID}}$ is computed on LSUN Horse 256 $\times$ 256 for the pixel-space diffusion process. As $\xi$ is increased, more timesteps are allocated to the region with low-noise level.
The computation time of $v_{\mathrm{FID}}(\alpha)$ is proportional to the number of images in the training dataset. For the pixel-space diffusion process, it takes about 35 hours for LSUN Horse 256 $\times$ 256 with 2M images using eight A100 GPUs. For the latent-space diffusion process, it takes about 12 hours for ImageNet 256 $\times$ 256 with 1.28M images using eight A100 GPUs. The computation of $v_{\mathrm{FID}}(\alpha)$ needs to be executed only once before training the diffusion model, and the computation time is sufficiently small compared with the training time of the diffusion model.
Algorithm 1 $v(\alpha)$ on basis of FID
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, number of timesteps for simulating the forward process $T$, discretized noise levels $\{\alpha_{t}\}_{t=0,...,T}$, feature model $\bm{\phi}$
[…]
1:  # Compute statistics of feature vectors
2:  for $n=0$ to $N-1$ do
3:     $\bm{x}^{(n)}_{0}\leftarrow\bm{x}^{(n)}$
4:     if diffusion process in latent space then
5:        $\bm{x}^{(n)}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}^{(n)}_{0})$
6:     end if
7:
8:     # Simulate the forward process
9:     for $t=1$ to $T$ do
10:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}}$
11:        $\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
12:        $x^{(n)}_{t}\leftarrow\beta_{t}\bm{x}^{(n)}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
13:     end for
14:  end for
15:  $\bm{\mu}_{t}\leftarrow mean\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
16:  $\bm{\Sigma}_{t}\leftarrow cov\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
17:
18:  # Compute $v(\alpha)$ using $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$
19:  for $t=0$ to $T-1$ do
20:     $v_{t}\leftarrow\frac{\mathrm{FID}(\bm{\mu}_{t},\bm{\Sigma}_{t},\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})}{\alpha_{t}-\alpha_{t+1}}$
21:  end for
22:  $v_{T}\leftarrow v_{T-1}$
Figure 4: Dependence of CRS-$v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$.
Figure 5: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $w_{x}$ and $w_{\mathrm{FID}}$ ($\xi_{x}=\xi_{\mathrm{FID}}=1)$.
Figure 6: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
Figure 7: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{x}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
C.2 Distance Metric Based on Data Prediction
Algorithm 2 presents the pseudocode for computing $v_{x}(\alpha)$ using a trained diffusion model. This pseudocode focuses on the optimization of sampling schedules. Similar to the computation of $v_{\mathrm{FID}}(\alpha)$, the forward process is simulated using the training dataset, and the change in the data prediction between adjacent timesteps is computed. A common configuration of hyperparameters $(T=10^{3},S=10^{4},\alpha_{s}=1,\alpha_{e}=10^{-4})$ was used in our experiments.
The computation time for obtaining $v_{x}(\alpha)$ is negligible compared with the training time of the diffusion model. It takes approximately 5.5 hours using eight H100 GPUs in pixel-space diffusion models. In latent-space diffusion models, it takes approximately 1.5 hours using eight A100 GPUs.
Algorithm 2 $v_{x}(\alpha)$ for optimizing sampling schedule.
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, trained diffusion model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$, number of timesteps for simulating the forward process $T$, number of samples used for calculating the mean $S$, maximum and minimum values of $\alpha$ $(\alpha_{s},\alpha_{e})$
0:  $LinearInterp1d(\{\alpha_{t}\}_{t=0,...,T},\{v_{t}\}_{t=0,...,T})$
1:  # Compute $\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right]$
2:  $\overline{D^{2}_{t}}\leftarrow 0,\ \forall{t}\in\{0,1,...,T\}$
3:  $\Delta\alpha\leftarrow\frac{1}{T}(\alpha_{s}-\alpha_{e})$
4:  for $n$ in $randperm(N)[:S]$ do
5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)}$
6:     if latent-space diffusion model then
7:        $\bm{x}_{0}\leftarrow autoencoder(\bm{x}_{0})$
8:     end if
9:     $\bm{y}_{0}\leftarrow\bm{x}_{0}$
10:
11:     # Simulate the forward process
12:     for $t=1$ to $T$ do
13:        $\alpha_{t}\leftarrow\alpha_{s}-\Delta\alpha t,\ \sigma_{t}\leftarrow\sqrt{1-\alpha_{t}^{2}}$
14:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}},\ \delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
15:        $\bm{x}_{t}\leftarrow\beta_{t}\bm{x}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
16:        $\bm{y}_{t}\leftarrow\frac{1}{\alpha_{t}}\left(\bm{x}_{t}-\sigma_{t}\bm{\varepsilon}_{\theta}(\bm{x}_{t},\alpha_{t})\right)$
17:        $\overline{D^{2}_{t}}\leftarrow\overline{D^{2}_{t}}+\frac{1}{S}\left\|\bm{y}_{t}-\bm{y}_{t-1}\right\|_{2}^{2}$
18:     end for
19:  end for
20:
21:  # Compute $v(\alpha)$ using $\overline{D^{2}_{t}}$
22:  for $t=1$ to $T$ do
23:     $v_{t}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{t}}}$
24:  end for
25:  $v_{0}\leftarrow v_{1}$
CRS-$v_{x}$ can be used for optimizing the training schedule as well. A pseudocode for training diffusion models with CRS is given in Algorithm 3. Since the data prediction is expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, $v_{x}(\alpha)$ tends to take extremely large values when $\alpha\simeq 0$. To avoid this, $v_{x}(\alpha)$ is evaluated within the range of $\alpha_{\mathrm{th}}\leq\alpha\leq\alpha_{\max}$, and we set $v_{x}(\alpha<\alpha_{\mathrm{th}})=v_{x}(\alpha_{\mathrm{th}})$. Similar to the implementation of the adaptive noise schedule in VDM++ [13], the range of $\alpha$ from $\alpha_{\mathrm{th}}$ to $\alpha_{\max}$ is divided into $B$ evenly spaced bins, and $\overline{D^{2}_{x}}(\alpha,\alpha+\Delta\alpha)$ is approximately evaluated using the exponential moving average (EMA). We used a common configuration of hyperparameters ($B=100,e=0.995,\alpha_{\max}=1,\alpha_{\min}=0,\alpha_{\mathrm{th}}=0.01,\Delta\alpha=10^{-3},\xi=1.0)$ in this study. The initial noise schedule is set to $\alpha(t)=\alpha_{\max}-(\alpha_{\max}-\alpha_{\min})\frac{t}{T}$. Although not explicitly included in the pseudocode, the update of $\overline{D^{2}_{b}}$ starts after 1000 iterations, and the noise schedule $\alpha(t)$ is updated every 100 iterations.
The training times of the diffusion models are listed in Table 5. When using CRS for optimizing training schedule, the training time increases by 20–30% compared with conventional noise schedules because CRS requires evaluating $v_{x}(\alpha)$ during training. To mitigate the increase in training time, one potential approach is to fix the training schedule after a certain number of epochs. In practice, CRS-$v_{x}$ showed little change beyond 10 epochs on LSUN Horse 256 $\times$ 256, suggesting the possibility of reducing training time while maintaining performance by fixing the training schedule after 10 epochs. Reducing the update frequency of $\overline{D^{2}_{x}}$ can be also an effective strategy. Developing an efficient implementation to reduce the training time remains as future work.
Algorithm 3 Adaptive noise schedule for training.
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N}$, initial noise schedule $\alpha(t)$, number of bins $B$, decay rate of EMA $e$, maximum and minimum values of $\alpha$ for training schedule $(\alpha_{\max},\alpha_{\min})$, infinitesimal variation of $\alpha$ for evaluating the derivative $\Delta\alpha$, minimum $\alpha$ for evaluating $v(\alpha)$ $\alpha_{\mathrm{th}}$, hyperparameter of CRS $\xi$
0:  noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$
1:  $\alpha_{b}\leftarrow\alpha_{\mathrm{th}}+(\alpha_{\max}-\alpha_{\mathrm{th}})\frac{b}{B},\ \forall{b}\in\{0,...,B\}$
2:  $\overline{D^{2}_{b}}\leftarrow 1e^{-6},\ \forall{b}\in\{0,...,B-1\}$
3:  for $n$ in $randperm(N)$ do
4:     # Compute loss for training diffusion models
5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)},\ t\sim\mathcal{U}(0,1),\ \bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
6:     $\alpha\leftarrow\alpha(t),\ \sigma\leftarrow\sqrt{1-\alpha^{2}}$
7:     $\bm{\varepsilon}_{\alpha}\leftarrow\bm{\varepsilon}_{\theta}(\alpha\bm{x}_{0}+\sigma\bm{\varepsilon},\alpha)$
8:     $loss\leftarrow\frac{1}{2}\left\|\bm{\varepsilon}_{\alpha}-\bm{\varepsilon}\right\|_{2}^{2}$
9:
10:     # Evaluate $\overline{D^{2}_{x}}$ using EMA (without gradient)
11:     if $\alpha\geq\alpha_{\mathrm{th}}$ then
12:        $\alpha^{\prime}\leftarrow\alpha-\Delta\alpha,\ \beta^{\prime}\leftarrow\frac{\alpha^{\prime}}{\alpha},\ \delta^{\prime}\leftarrow\sqrt{1-\beta^{\prime 2}}$
13:        $\bm{x}_{\alpha^{\prime}}\leftarrow\beta^{\prime}\bm{x}_{\alpha}+\delta^{\prime}\times\mathcal{N}(\bm{0},\bm{I})$
14:        $\bm{\varepsilon}_{\alpha^{\prime}}\leftarrow\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})$
15:        $\bm{y}_{\alpha}\leftarrow\frac{1}{\alpha}\left(\bm{x}_{\alpha}-\sigma\bm{\varepsilon}_{\alpha}\right),\ \bm{y}_{\alpha^{\prime}}\leftarrow\frac{1}{\alpha^{\prime}}\left(\bm{x}_{\alpha^{\prime}}-\sigma^{\prime}\bm{\varepsilon}_{\alpha^{\prime}}\right)$
16:        $b\leftarrow\min\left(\left\lfloor\frac{\alpha-\alpha_{\mathrm{th}}}{\alpha_{\max}-\alpha_{\mathrm{th}}}B\right\rfloor,B-1\right)$
17:        $\overline{D^{2}_{b}}\leftarrow e\overline{D^{2}_{b}}+(1-e)\left\|\bm{y}_{\alpha}-\bm{y}_{\alpha^{\prime}}\right\|_{2}^{2}$
18:        $v_{b}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{b}}}$
19:     end if
20:
21:     # Update training schedule $\alpha(t)$
22:     $\mathcal{X}\leftarrow\{\alpha_{\min},\alpha_{0},\alpha_{1},...,\alpha_{B}\}$
23:     $\mathcal{Y}\leftarrow\{v_{0},v_{0},v_{1},...,v_{B-1},v_{B-1}\}$
24:     $v(\alpha)\leftarrow LinearInterp1d(\mathcal{X},\mathcal{Y})$
25:     Update $\alpha(t)$ and model parameters $\theta$
26:  end for
Table 5: Training time of diffusion models.
Diffusion type Dataset GPUs Epochs Training schedule Training time
Latent space LSUN Church A100 $\times$ 8 1000
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
85 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
105 hours
LSUN Bedroom A100 $\times$ 8 100
CRS-$v_{\mathrm{FID}}$
190 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
245 hours
ImageNet A100 $\times$ 8 500
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
430 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
525 hours
Pixel space LSUN Horse H100 $\times$ 8 50
Linear
Shifted cosine
VDM++
170 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
215 hours
LSUN Bedroom H100 $\times$ 8 50
CRS-$v_{x}$
325 hours
Appendix D Hyperparameter Tuning for Combining Multiple Distance Metrics
CRS-$v_{x}+v_{\mathrm{FID}}$ has four hyperparameters; $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$. We tuned these hyperparameters in two steps. In the first step, we fixed $\xi_{x}=\xi_{\mathrm{FID}}=1$ and searched for appropriate values of $w_{x}$ and $w_{\mathrm{FID}}$, subject to the constraint $w_{x}+w_{\mathrm{FID}}=1$. In the second step, we fixed $w_{x}$ and $w_{\mathrm{FID}}$, then adjusted $\xi_{x}$ and $\xi_{\mathrm{FID}}$.
The detailed results on LSUN Horse 256 $\times$ 256 are listed in Table 6. In the first step, adjusting $w_{x}$ and $w_{\mathrm{FID}}$ led to significant changes in FID at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. In the second step, the improvement from tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ was small at $\mathrm{NFE}=250$. In contrast, for $\mathrm{NFE}=50$, the tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ resulted in a substantial improvement in FID, indicating that the importance of hyperparameter tuning increases as NFE decreases.
Table 6: Detailed results of hyperparameter tuning on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model. DPM-Solver++(2M) was used as sampler. To conserve computational resources, FID score was calculated using 10K samples.
NFE Tuning step $w_{\mathrm{FID}}$ $x_{x}$ $\xi_{\mathrm{FID}}$ $\xi_{x}$ FID (10K)
250 1st 0.9 0.1 1.0 1.0 3.78
0.7 0.3 3.70
0.5 0.5 3.70
0.3 0.7 3.73
0.1 0.9 3.76
2nd 0.7 0.3 0.5 0.5 3.77
1.0 4.30
1.2 3.75
1.4 3.94
1.0 0.5 3.76
1.0 3.70
1.2 3.69
1.4 3.78
1.2 0.5 3.82
1.0 3.69
1.2 3.78
1.4 3.77
1.4 0.5 3.83
1.0 3.78
1.2 3.77
1.4 3.77
50 1st 0.9 0.1 1.0 1.0 6.13
0.7 0.3 4.42
0.5 0.5 4.38
0.3 0.7 4.78
0.1 0.9 7.03
2nd 0.5 0.5 0.5 0.5 7.42
1.0 7.44
1.2 7.48
1.4 7.79
1.0 0.5 4.85
1.0 4.38
1.2 4.32
1.4 4.27
1.2 0.5 4.57
1.0 4.09
1.2 3.96
1.4 3.93
1.4 0.5 4.51
1.0 3.95
1.2 3.85
1.4 3.82
The dependence of noise schedules on the hyperparameters is shown in Figs. 7, 7, and 7. Increasing either $w_{\mathrm{FID}}$ or $\xi_{\mathrm{FID}}$ amplifies the impact of $v_{\mathrm{FID}}(\alpha)$, resulting in more timesteps being allocated to the region where $\alpha\simeq 1$ (see Figs. 7 and 7). On the other hand, increasing either $w_{x}$ or $\xi_{x}$ strengthens the impact of $v_{x}(\alpha)$, causing more timesteps to be allocated to the region where $\alpha\simeq 0$ (see Figs. 7 and 7).
The hyperparameter settings for CRS-$v_{x}+v_{\mathrm{FID}}$ are listed in Table 7. Noise schedules $\alpha(t)$ obtained using CRS-$v_{x}+v_{\mathrm{FID}}$ are illustrated in Figs. 13, 13, 13, 13, 13, and 13.
Table 7: Hyperparameter settings of CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse 256 $\times$ 256, LSUN Bedroom 256 $\times$ 256, FFHQ 64 $\times$ 64, and CIFAR10 32 $\times$ 32.
Dataset Sampler NFE $w_{\mathrm{FID}}$ $w_{x}$ $\xi_{\mathrm{FID}}$ $\xi_{x}$
LSUN Horse 256 $\times$ 256 SDE-DPM-Solver++(2M) 250 0.9 0.1 1.4 1.0
50 0.7 0.3 1.4 1.2
PNDM 250 0.9 0.1 1.0 0.5
50 0.5 0.5 1.2 0.5
DPM-Solver++(2M) 250 0.7 0.3 1.0 1.2
50 0.5 0.5 1.4 1.4
LSUN Bedroom 256 $\times$ 256 DPM-Solver++(2M) 10 0.5 0.5 1.0 1.0
5 0.5 0.5 1.0 1.2
UniPC 10 0.5 0.5 1.0 1.2
5 0.3 0.7 1.0 1.0
FFHQ 64 $\times$ 64 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.4
20 0.5 0.5 1.6 1.4
10 0.5 0.5 1.6 1.5
iPNDM 50 0.5 0.5 1.2 1.5
20 0.5 0.5 1.4 1.4
10 0.5 0.5 1.5 1.4
CIFAR10 32 $\times$ 32 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.0
20 0.5 0.5 1.8 2.4
10 0.5 0.5 1.6 3.0
iPNDM 50 0.7 0.3 1.6 1.2
20 0.5 0.5 1.6 1.6
10 0.5 0.5 1.6 3.2
Figure 8: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with SDE-DPM-Solver++(2M).
Figure 9: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with PNDM.
Figure 10: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with DPM-Solver++(2M).
Figure 11: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with DPM-Solver++(2M).
Figure 12: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with UniPC.
Figure 13: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating FFHQ 64 $\times$ 64 with DPM-Solver++(2M).
Appendix E Implementation Details
E.1 Unet Model
The hyperparameters of the U-Net model are listed in Table 8. For latent-space diffusion models, we used the same values as the ImageNet 64 $\times$ 64 in ADM [52]. For pixel-space diffusion models, we used the same values as LSUN 256 $\times$ 256 in ADM [52].
Table 8: Hyperparameters of UNet model.
Latent-space diffusion models Pixel-space diffusion models
Resolution 64 $\times$ 64 256 $\times$ 256
Number of parameters 296M 552M
Channels 192 256
Depth 3 2
Channels multiple 1,2,3,4 1,1,2,2,4,4
Heads channels 64 64
Attention resolution 32,16,8 32,16,8
BigGAN up/downsample True True
Dropout 0.1 0.1
E.2 Hyperparameters for Training
The hyperparameters for training the diffusion models are summarized in Table 9. We saved ten checkpoints for each model. The FID score with 10K samples was evaluated for all checkpoints, and the results are reported with the checkpoint with the best FID score.
Table 9: Hyperparameters for training diffusion models. Same configuration was used for training latent-space and pixel-space diffusion models.
Item Value
EMA decay 0.9999
Optimizer Adam
$\beta_{1}$ 0.9
$\beta_{2}$ 0.999
Learning rate 1e-4
Batch size 32 $\times$ 8
E.3 Noise Schedule for Training and Sampling
The conventional noise schedules used in this study are summarized in Table 10.
When using a CRS-based training schedule, a common hyperparameter setting of $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.0)$ is used (details are described in Appendix C.2). When using a CRS-based sampling schedule, the default setting is $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.01)$. An exception was made when using CRS-$v_{x}+v_{\mathrm{FID}}$, where the hyperparameters are tuned according to the procedure described in Appendix D. For the experiments using pretrained EDMs, we used $(\alpha_{\max},\alpha_{\min})=(0.999998,0.0125)$, which is equivalent to the boundary condition of EDM: $(\sigma_{\min},\sigma_{\max})=(0.002,80)$. The boundary condition of EDM is transformed to $\alpha_{\max}$ and $\alpha_{\min}$ using Eq. (16).
Table 10: Conventional noise schedules evaluated in this study. $T_{s}$ denotes number of timesteps for sampling.
Name Phase Noise schedule
Linear Training
$\alpha_{\mathrm{linear}}(t)=\mathrm{LinearInterp1d}\left(\left\{\frac{i}{T}\right\}_{i=0,...,T},\left\{\alpha_{i}\right\}_{i=0,...,T}\right),$
$\alpha_{0}=1,$
$\alpha_{1\leq i\leq T}=\sqrt{1-\tilde{\beta}_{i}}\alpha_{i-1},$
$\tilde{\beta}_{i}=\tilde{\beta}_{\min}+\frac{\tilde{\beta}_{\max}-\tilde{\beta}_{\min}}{T-1}(i-1),$
$T=1000,\tilde{\beta}_{\min}=10^{-4},\tilde{\beta}_{\max}=0.02.$
Sampling $\left\{\alpha_{\mathrm{linear}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\}$
Shifted cosine Training
$\alpha_{\mathrm{shifted}}(t)=\sqrt{\mathrm{sigmoid}(\lambda(t))},$
$\lambda(t)=-2\log\tan\left(\frac{\pi t}{2}\right)+2\log\left(\frac{64}{d}\right),$
where $d$ is resolution in diffusion process.
Sampling
$\left\{\tilde{\alpha}_{\mathrm{shifted}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\},$
$\tilde{\alpha}_{\mathrm{shifted}}(t)=\alpha_{\min}+(\alpha_{\max}-\alpha_{\min})\alpha_{\mathrm{shifted}}(t),$
$\alpha_{\min}=0.01,\alpha_{\max}=1.0.$
VDM++ Training
$p(\lambda)\propto\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\omega(\lambda)\left\|\bm{\varepsilon}_{\theta}(\bm{x},\tilde{\alpha}(\lambda))-\bm{\varepsilon}\right\|_{2}^{2}\right],$
[…]
$\tilde{\omega}(\lambda)=\mathcal{N}(\lambda;\lambda_{\mu},\lambda_{\sigma}^{2})(e^{-\lambda}+c^{2}),$
$\mu_{\lambda}=2.4,\sigma_{\lambda}=2.4,c=0.5.$
Sampling Only for training.
EDM Training Not evaluated in this study.
Sampling
$\alpha_{0}=1,\alpha_{t\geq 1}=\alpha_{\mathrm{EDM}}\left(\frac{t}{T_{s}}\right),$
$\alpha_{\mathrm{EDM}}(t)=\sqrt{\mathrm{sigmoid}(\lambda(t))},$
$\lambda(t)=-2\log\sigma(t),$
$\sigma(t)=\left\{\sigma_{\max}^{\frac{1}{\rho}}+\left(\sigma_{\min}^{\frac{1}{\rho}}-\sigma_{\max}^{\frac{1}{\rho}}\right)(1-t)\right\}^{\rho},$
$\sigma_{\max}=80,\sigma_{\min}=0.002,\rho=7.$
Appendix F Detailed Results
We present the detailed experimental results and confirm the superiority of CRS for optimizing both training and sampling schedules.
F.1 Results on Pixel-Space Diffusion Models
LSUN Horse 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=250$ are listed in Table 11. To verify the generality with respect to samplers, we used three different samplers: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M). As shown in the rows shaded in gray, multiple evaluation metrics, including FID score, are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule, regardless of the sampler.
Table 12 presents the dependence on the sampling schedules at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. When using SDE-DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ improved upon all four evaluation metrics at $\mathrm{NFE}=250$, achieving a new state-of-the-art FID score. When using PNDM and DPM-Solver++(2M), there was little difference between sampling schedules at $\mathrm{NFE}=250$. However, at $\mathrm{NFE}=50$, the performance became more sensitive to the sampling schedule, and CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the best FID score.
LSUN Bedroom 256 $\times$ 256: Since only the FID score is reported in the main part of this study, we include sFID, precision, and recall in Table 13 for reference.
F.2 Results on Latent-Space Diffusion Models
LSUN Church 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 14. To verify the generality with respect to samplers, we used five different samplers: Stochastic DDIM, SDE-DPM-Solver++(2M), DDIM, PNDM, and DPM-Solver++(2M). As shown in the rows shaded in gray, FID scores are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule across all samplers except for SDE-DPM-Solver++(2M).
The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 15 and 16, respectively. CRS-$v_{x}$ or CRS-$v_{\varepsilon}$ achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
LSUN Bedroom 256 $\times$ 256: The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 17 and 18, respectively. CRS-$v_{x}$ is used for optimizing the training schedule. CRS achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=30$ and $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
FID scores evaluated at small NFEs are listed in Table 19. Compared with the results reported in DPM-Solver-v3 [26], the better FID scores are achieved by using CRS for optimizing both the training and sampling schedules at $\mathrm{NFE}=10$ and $\mathrm{NFE}=20$. FID, sFID, precision, and recall are listed in Table 20. CRS based noise schedules consistently outperform the linear schedule.
ImageNet 256 $\times$ 256: We demonstrated the superiority of CRS in class-conditional image generation. The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 21. As shown in the rows shaded in gray, CRS-$v_{\mathrm{FID}}$ was the best for optimizing training schedule for ImageNet, while CRS-$v_{x}$, which consistently showed strong performance on other datasets, was the second-best. The reason CRS-$v_{\mathrm{FID}}$ performed better on ImageNet may be attributed to the fact that the feature model used to compute FID is trained on ImageNet classification.
The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Tables 22 and 23, respectively. We used CRS-$v_{x}$ for optimizing training schedule, which outperformed conventional noise schedules with significant margin on ImageNet and exhibited the best performance across a wide range of datasets. CRS achieved the best FID score in all cases except when using PNDM at $\mathrm{NFE}=20$ and DPM-Solver++(2M) at $\mathrm{NFE}=30$, and multiple evaluation metrics are simultaneously improved in most cases.
F.3 Results on Pretrained EDM
FFHQ 64 $\times$ 64: The evaluation results are listed in Table 24. When using DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ outperforms both EDM and AYS. Using iPNDM further improves FID scores, and CRS-$v_{x}+v_{\mathrm{FID}}$ consistently performs better than EDM.
CIFAR10 32 $\times$ 32: The evaluation results on class-conditional image generation are listed in Table 25. CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperforms EDM, and significant performance gains are observed at small NFEs.
Appendix G Generated Samples
Generated images are shown in the following figures:
  • LSUN Horse 256 $\times$ 256 in the pixel-space diffusion model: Figs. 14 and 15.
  • LSUN Bedroom 256 $\times$ 256 in the pixel-space diffusion model: Figs. 16 and 17.
  • LSUN Church 256 $\times$ 256 in the latent-space diffusion model: Fig. 18.
  • LSUN Bedroom 256 $\times$ 256 in the latent-space diffusion model: Fig. 19.
  • ImageNet 256 $\times$ 256 in the latent-space diffusion model: Fig. 20.
All models are trained using CRS-$v_{x}$ for optimizing training schedule.
Table 11: Dependence on training and sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models at $\mathrm{NFE}=250$. To verify generality with respect to samplers, we used three different samplers. Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric. Note that "cosine" refers to shifted cosine in this table.
Training Sampling Metrics
Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
SDE-DPM- Linear Linear 2.90 6.82 0.66 0.56
Solver++(2M) Cosine 2.95 6.43 0.67 0.56
VDM++ 3.82 7.40 0.66 0.54
CRS-$v_{\varepsilon}$ 2.91 6.56 0.66 0.56
CRS-$v_{x}$ 2.73 6.40 0.67 0.56
CRS-$v_{x}$ Linear 2.73 6.40 0.67 0.56
Cosine 3.13 7.09 0.67 0.54
EDM 2.09 6.16 0.69 0.56
CRS-$v_{\varepsilon}$ 2.87 6.71 0.66 0.56
CRS-$v_{x}$ 5.46 8.34 0.62 0.52
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.03 6.06 0.69 0.56
PNDM Linear Linear 4.06 6.23 0.57 0.61
Cosine 2.36 5.59 0.63 0.60
VDM++ 70.18 36.78 0.20 0.53
CRS-$v_{\mathrm{FID}}$ 8.20 7.92 0.51 0.60
CRS-$v_{x}$ 2.09 5.43 0.65 0.58
CRS-$v_{x}$ Linear 2.09 5.43 0.65 0.58
Cosine 2.19 5.59 0.67 0.57
EDM 2.35 5.85 0.67 0.56
CRS-$v_{\mathrm{\varepsilon}}$ 2.11 5.58 0.66 0.58
CRS-$v_{x}$ 2.50 5.94 0.64 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.12 5.50 0.67 0.57
DPM- Linear Linear 3.17 5.98 0.60 0.61
Solver++(2M) Cosine 2.31 5.63 0.63 0.60
VDM++ 47.63 29.77 0.26 0.56
CRS-$v_{\varepsilon}$ 7.12 7.45 0.54 0.60
CRS-$v_{x}$ 2.08 5.47 0.65 0.59
CRS-$v_{x}$ Linear 2.08 5.47 0.65 0.59
Cosine 2.18 5.72 0.66 0.58
EDM 2.35 5.85 0.67 0.56
CRS-$v_{x}$ 2.12 5.65 0.66 0.58
CRS-$v_{\varepsilon}$ 2.87 6.15 0.63 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.19 5.56 0.67 0.57
Table 12: Dependence on sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
SDE-DPM- 250 Linear 2.73 6.40 0.67 0.56
Solver++(2M) Shifted cosine 3.13 7.09 0.67 0.54
EDM 2.09 6.16 0.69 0.56
CRS-$v_{\varepsilon}$ 2.87 6.71 0.66 0.56
CRS-$v_{x}$ 5.46 8.34 0.62 0.52
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.03 6.06 0.69 0.56
50 Linear 7.54 9.73 0.59 0.48
Shifted cosine 11.37 12.46 0.53 0.43
EDM 2.75 6.91 0.68 0.55
CRS-$v_{\varepsilon}$ 6.08 8.75 0.61 0.50
CRS-$v_{x}$ 12.99 12.59 0.50 0.41
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.84 8.16 0.67 0.52
PNDM 250 Linear 2.09 5.43 0.65 0.58
Shifted cosine 2.19 5.59 0.67 0.57
EDM 2.35 5.85 0.67 0.56
CRS-$v_{\varepsilon}$ 2.11 5.58 0.66 0.58
CRS-$v_{x}$ 2.50 5.94 0.64 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.12 5.50 0.67 0.57
50 Linear 2.50 5.76 0.62 0.59
Shifted cosine 3.92 6.77 0.60 0.57
EDM 2.36 5.87 0.67 0.56
CRS-$v_{\varepsilon}$ 2.46 5.97 0.64 0.58
CRS-$v_{x}$ 6.03 7.32 0.57 0.56
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.04 5.60 0.66 0.58
DPM- 250 Linear 2.08 5.47 0.65 0.59
Solver++(2M) Shifted Cosine 2.18 5.72 0.66 0.58
EDM 2.35 5.85 0.67 0.56
CRS-$v_{\varepsilon}$ 2.12 5.65 0.66 0.58
CRS-$v_{x}$ 2.87 6.15 0.63 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.19 5.56 0.67 0.57
50 Linear 3.17 5.98 0.60 0.59
Shifted cosine 5.27 7.12 0.57 0.56
EDM 2.43 5.87 0.67 0.56
CRS-$v_{\varepsilon}$ 3.10 6.18 0.62 0.58
CRS-$v_{x}$ 7.44 7.83 0.55 0.55
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.26 5.90 0.66 0.57
Table 13: Evaluation results on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion model. CRS-$v_{x}$ was used for optimizing training schedule.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DPM-Solver++(2M) 10 EDM 4.73 7.75 0.53 0.51
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.88 8.67 0.49 0.52
5 EDM 23.72 19.24 0.20 0.28
CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 17.49 0.28 0.41
UniPC 10 EDM 4.94 8.20 0.53 0.40
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.30 6.60 0.55 0.53
5 EDM 82.52 33.35 0.06 0.08
CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 11.26 0.21 0.44
Table 14: Dependence on training and sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$. To verify generality with respect to samplers, we used five different samplers. Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric.
Training Sampling Metrics
Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM Linear Linear 10.73 18.31 0.57 0.33
VDM++ 11.05 17.33 0.56 0.35
CRS-$v_{\mathrm{FID}}$ 10.85 18.77 0.57 0.34
CRS-$v_{\varepsilon}$ 10.43 17.81 0.57 0.34
CRS-$v_{x}$ 9.82 16.39 0.58 0.34
CRS-$v_{x}$ Linear 9.82 16.39 0.58 0.34
EDM 21.73 29.33 0.43 0.22
CRS-$v_{\mathrm{FID}}$ 7.36 13.04 0.60 0.39
CRS-$v_{\varepsilon}$ 7.53 13.73 0.61 0.38
CRS-$v_{x}$ 6.77 12.57 0.63 0.40
SDE-DPM- Linear Linear 3.49 10.19 0.63 0.57
Solver++(2M) VDM++ 3.33 9.79 0.63 0.57
CRS-$v_{\mathrm{FID}}$ 3.50 10.69 0.63 0.57
CRS-$v_{\varepsilon}$ 3.57 10.26 0.64 0.56
CRS-$v_{x}$ 3.38 10.09 0.64 0.55
CRS-$v_{x}$ Linear 3.38 10.09 0.64 0.55
EDM 4.20 11.69 0.63 0.54
CRS-$v_{\mathrm{FID}}$ 3.61 9.51 0.61 0.57
CRS-$v_{\varepsilon}$ 3.19 9.56 0.62 0.59
CRS-$v_{x}$ 3.20 9.14 0.62 0.58
DDIM Linear Linear 6.18 13.16 0.59 0.51
VDM++ 6.50 12.67 0.56 0.54
CRS-$v_{\mathrm{FID}}$ 6.06 13.69 0.59 0.51
CRS-$v_{\varepsilon}$ 6.05 12.49 0.58 0.50
CRS-$v_{x}$ 5.67 12.33 0.59 0.51
CRS-$v_{x}$ Linear 5.67 12.33 0.59 0.51
EDM 9.15 16.86 0.56 0.44
CRS-$v_{\mathrm{FID}}$ 5.35 10.70 0.57 0.53
CRS-$v_{\varepsilon}$ 4.95 11.01 0.59 0.53
CRS-$v_{x}$ 4.81 10.79 0.59 0.55
PNDM Linear Linear 3.74 10.48 0.60 0.58
VDM++ 3.88 10.20 0.59 0.59
CRS-$v_{\mathrm{FID}}$ 3.59 10.61 0.61 0.59
CRS-$v_{\varepsilon}$ 3.71 10.36 0.60 0.59
CRS-$v_{x}$ 3.56 10.37 0.61 0.58
CRS-$v_{x}$ Linear 3.56 10.37 0.61 0.58
EDM 3.88 10.76 0.60 0.57
CRS-$v_{\mathrm{FID}}$ 3.59 9.69 0.59 0.61
CRS-$v_{\varepsilon}$ 3.48 9.79 0.59 0.59
CRS-$v_{x}$ 3.62 10.15 0.59 0.61
DPM- Linear Linear 3.82 10.28 0.60 0.58
Solver++(2M) VDM++ 4.00 10.07 0.58 0.59
CRS-$v_{\mathrm{FID}}$ 3.78 10.51 0.60 0.58
CRS-$v_{\varepsilon}$ 3.87 10.30 0.60 0.59
CRS-$v_{x}$ 3.69 10.23 0.60 0.59
CRS-$v_{x}$ Linear 3.69 10.23 0.60 0.59
EDM 4.31 11.31 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.98 9.61 0.58 0.60
CRS-$v_{\varepsilon}$ 3.63 9.91 0.59 0.60
CRS-$v_{x}$ 3.59 9.51 0.59 0.60
Table 15: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM 50 Linear 6.23 11.94 0.65 0.43
EDM 10.05 17.14 0.60 0.34
CRS-$v_{\mathrm{FID}}$ 5.17 10.49 0.65 0.44
CRS-$v_{\varepsilon}$ 5.26 10.94 0.65 0.44
CRS-$v_{x}$ 4.82 11.91 0.62 0.51
30 Linear 9.82 16.39 0.58 0.34
EDM 21.73 29.33 0.43 0.22
CRS-$v_{\mathrm{FID}}$ 7.36 13.04 0.60 0.39
CRS-$v_{\varepsilon}$ 7.53 13.73 0.61 0.38
CRS-$v_{x}$ 6.77 12.57 0.63 0.40
20 Linear 18.24 25.36 0.45 0.24
EDM 48.46 52.53 0.25 0.11
CRS-$v_{\mathrm{FID}}$ 12.42 18.38 0.51 0.29
CRS-$v_{\varepsilon}$ 12.53 19.27 0.52 0.29
CRS-$v_{x}$ 10.66 16.54 0.55 0.32
SDE-DPM- 50 Linear 3.42 10.58 0.63 0.57
Solver++(2M) EDM 3.57 10.93 0.64 0.57
CRS-$v_{\mathrm{FID}}$ 3.32 9.19 0.62 0.58
CRS-$v_{\varepsilon}$ 3.14 9.65 0.63 0.58
CRS-$v_{x}$ 3.09 9.34 0.63 0.59
30 Linear 3.38 10.09 0.64 0.55
EDM 4.20 11.69 0.63 0.54
CRS-$v_{\mathrm{FID}}$ 3.61 9.51 0.61 0.57
CRS-$v_{\varepsilon}$ 3.19 9.56 0.62 0.59
CRS-$v_{x}$ 3.20 9.14 0.62 0.58
20 Linear 3.55 10.37 0.63 0.56
EDM 5.92 13.79 0.62 0.52
CRS-$v_{\mathrm{FID}}$ 4.39 9.33 0.63 0.48
CRS-$v_{\varepsilon}$ 3.58 9.93 0.61 0.57
CRS-$v_{x}$ 3.57 9.55 0.62 0.56
Table 16: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DDIM 50 Linear 4.52 11.03 0.60 0.54
EDM 5.95 13.37 0.59 0.50
CRS-$v_{\mathrm{FID}}$ 4.28 10.07 0.59 0.56
CRS-$v_{\varepsilon}$ 4.17 10.66 0.60 0.56
CRS-$v_{x}$ 4.01 10.10 0.60 0.57
30 Linear 5.67 12.33 0.59 0.51
EDM 9.15 16.86 0.56 0.44
CRS-$v_{\mathrm{FID}}$ 5.35 10.70 0.57 0.53
CRS-$v_{\varepsilon}$ 4.95 11.01 0.59 0.53
CRS-$v_{x}$ 4.81 10.79 0.59 0.55
20 Linear 7.85 14.66 0.56 0.47
EDM 15.74 22.86 0.48 0.34
CRS-$v_{\mathrm{FID}}$ 7.47 12.74 0.54 0.48
CRS-$v_{\varepsilon}$ 6.49 12.55 0.56 0.49
CRS-$v_{x}$ 6.09 11.72 0.57 0.50
PNDM 50 Linear 3.52 10.42 0.61 0.58
EDM 3.71 10.59 0.60 0.57
CRS-$v_{\mathrm{FID}}$ 3.49 9.95 0.60 0.61
CRS-$v_{\varepsilon}$ 3.48 10.18 0.60 0.59
CRS-$v_{x}$ 3.48 9.92 0.60 0.60
30 Linear 3.56 10.37 0.61 0.58
EDM 3.88 10.76 0.60 0.57
CRS-$v_{\mathrm{FID}}$ 3.59 9.69 0.59 0.61
CRS-$v_{\varepsilon}$ 3.48 9.79 0.59 0.59
CRS-$v_{x}$ 3.62 10.15 0.59 0.61
20 Linear 3.74 10.24 0.60 0.59
EDM 4.57 11.49 0.59 0.56
CRS-$v_{\mathrm{FID}}$ 4.06 9.54 0.57 0.61
CRS-$v_{\varepsilon}$ 3.64 9.98 0.59 0.61
CRS-$v_{x}$ 3.86 9.54 0.62 0.52
DPM- 50 Linear 3.62 10.44 0.61 0.58
Solver++(2M) EDM 3.90 10.85 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.59 9.68 0.59 0.60
CRS-$v_{\varepsilon}$ 3.46 9.82 0.60 0.59
CRS-$v_{x}$ 3.43 9.45 0.60 0.59
30 Linear 3.69 10.23 0.60 0.59
EDM 4.31 11.31 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.98 9.61 0.58 0.60
CRS-$v_{\varepsilon}$ 3.63 9.91 0.59 0.60
CRS-$v_{x}$ 3.59 9.51 0.59 0.60
20 Linear 3.92 10.03 0.59 0.59
EDM 5.25 12.19 0.59 0.55
CRS-$v_{\mathrm{FID}}$ 4.95 9.92 0.56 0.58
CRS-$v_{\varepsilon}$ 3.89 9.73 0.58 0.60
CRS-$v_{x}$ 4.00 9.44 0.58 0.59
Table 17: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM 50 Linear 3.95 8.98 0.56 0.48
EDM 7.43 16.05 0.54 0.38
CRS-$v_{\mathrm{FID}}$ 3.09 7.17 0.57 0.50
CRS-$v_{\varepsilon}$ 3.25 7.81 0.57 0.49
CRS-$v_{x}$ 3.08 7.20 0.58 0.50
30 Linear 7.28 15.70 0.53 0.38
EDM 17.03 33.60 0.40 0.26
CRS-$v_{\mathrm{FID}}$ 5.17 10.62 0.57 0.43
CRS-$v_{\varepsilon}$ 5.58 12.33 0.56 0.42
CRS-$v_{x}$ 4.78 9.85 0.55 0.45
20 Linear 14.53 29.25 0.42 0.28
EDM 38.01 68.62 0.21 0.13
CRS-$v_{\mathrm{FID}}$ 9.18 16.67 0.50 0.35
CRS-$v_{\varepsilon}$ 9.70 19.40 0.49 0.34
CRS-$v_{x}$ 8.11 15.96 0.51 0.36
SDE-DPM- 50 Linear 2.47 6.44 0.57 0.54
Solver++(2M) EDM 2.48 6.32 0.57 0.54
CRS-$v_{\mathrm{FID}}$ 2.21 5.87 0.57 0.54
CRS-$v_{\varepsilon}$ 2.22 5.76 0.57 0.55
CRS-$v_{x}$ 2.23 5.67 0.57 0.54
30 Linear 2.45 6.70 0.56 0.55
EDM 2.90 7.41 0.57 0.52
CRS-$v_{\mathrm{FID}}$ 2.46 6.19 0.57 0.54
CRS-$v_{\varepsilon}$ 2.50 5.93 0.57 0.53
CRS-$v_{x}$ 2.48 5.73 0.56 0.54
20 Linear 2.78 6.79 0.56 0.53
EDM 4.00 9.57 0.55 0.50
CRS-$v_{\mathrm{FID}}$ 3.19 6.70 0.55 0.53
CRS-$v_{\varepsilon}$ 2.87 6.22 0.57 0.53
CRS-$v_{x}$ 3.07 6.34 0.56 0.52
Table 18: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DDIM 50 Linear 3.02 6.59 0.53 0.54
EDM 4.17 7.64 0.52 0.51
CRS-$v_{\mathrm{FID}}$ 2.66 6.03 0.54 0.55
CRS-$v_{\varepsilon}$ 2.71 6.03 0.54 0.55
CRS-$v_{x}$ 2.62 5.86 0.54 0.55
30 Linear 3.76 7.50 0.53 0.51
EDM 6.81 9.98 0.47 0.46
CRS-$v_{\mathrm{FID}}$ 3.23 6.61 0.53 0.53
CRS-$v_{\varepsilon}$ 3.22 6.65 0.53 0.53
CRS-$v_{x}$ 3.10 6.43 0.54 0.53
20 Linear 5.38 9.27 0.49 0.48
EDM 12.57 14.55 0.38 0.40
CRS-$v_{\mathrm{FID}}$ 4.39 7.86 0.51 0.49
CRS-$v_{\varepsilon}$ 4.28 7.97 0.51 0.49
CRS-$v_{x}$ 4.10 7.54 0.51 0.51
PNDM 50 Linear 2.51 6.02 0.54 0.56
EDM 2.64 6.16 0.53 0.57
CRS-$v_{\mathrm{FID}}$ 2.50 5.66 0.54 0.57
CRS-$v_{\varepsilon}$ 2.44 5.73 0.54 0.57
CRS-$v_{x}$ 2.45 5.61 0.54 0.56
30 Linear 2.57 6.03 0.54 0.56
EDM 2.93 6.45 0.52 0.57
CRS-$v_{\mathrm{FID}}$ 2.64 5.59 0.53 0.57
CRS-$v_{\varepsilon}$ 2.48 5.58 0.54 0.56
CRS-$v_{x}$ 2.65 5.53 0.54 0.57
20 Linear 2.77 6.13 0.53 0.57
EDM 3.76 7.25 0.52 0.54
CRS-$v_{\mathrm{FID}}$ 3.02 5.66 0.52 0.57
CRS-$v_{\varepsilon}$ 2.71 5.53 0.54 0.56
CRS-$v_{x}$ 3.63 6.23 0.61 0.57
DPM- 50 Linear 2.59 6.15 0.54 0.56
Solver++(2M) EDM 2.86 6.44 0.52 0.57
CRS-$v_{\mathrm{FID}}$ 2.46 5.63 0.54 0.56
CRS-$v_{\varepsilon}$ 2.46 5.63 0.54 0.56
CRS-$v_{x}$ 2.37 5.42 0.54 0.56
30 Linear 2.68 6.18 0.54 0.56
EDM 3.40 7.09 0.51 0.56
CRS-$v_{\mathrm{FID}}$ 2.60 5.63 0.54 0.56
CRS-$v_{\varepsilon}$ 2.60 5.53 0.54 0.56
CRS-$v_{x}$ 2.63 5.45 0.54 0.55
20 Linear 2.89 6.24 0.53 0.56
EDM 4.14 8.07 0.52 0.53
CRS-$v_{\mathrm{FID}}$ 3.03 5.81 0.52 0.56
CRS-$v_{\varepsilon}$ 2.88 5.59 0.53 0.56
CRS-$v_{x}$ 3.00 5.63 0.54 0.55
Table 19: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs. Bold values indicate best FID scores at each NFE. Results for rows other than "Ours" are taken from previously published work [26].
Sampling FID $\downarrow$
Model Sampler schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$ $\mathrm{NFE}=20$
LDM [57] DPM-Solver++ Linear 18.59 3.63 3.16
UniPC 12.24 3.56 3.07
DPM-Solver-v3 [26] 7.54 3.16 3.05
Ours iPNDM Linear 10.81 3.71 2.64
CRS-$v_{\varepsilon}$ 7.88 3.12 2.48
CRS-$v_{x}$ 8.07 3.45 2.50
Table 20: Detailed results evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
iPNDM 20 Linear 2.64 6.15 0.53 0.57
CRS-$v_{\varepsilon}$ 2.48 5.48 0.54 0.56
CRS-$v_{x}$ 2.50 5.41 0.55 0.56
10 Linear 3.71 6.88 0.52 0.54
CRS-$v_{\varepsilon}$ 3.12 5.69 0.52 0.56
CRS-$v_{x}$ 3.45 6.15 0.54 0.53
5 Linear 10.81 13.82 0.42 0.49
CRS-$v_{\varepsilon}$ 7.88 9.84 0.45 0.50
CRS-$v_{x}$ 8.07 10.43 0.45 0.48
Table 21: Dependence on training and sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$. To verify generality with respect to samplers, we used five different samplers. Rows shaded in gray represent dependence on training schedule, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric.
Training Sampling Metrics
Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM Linear Linear 27.31 21.39 0.46 0.55
VDM++ 29.00 17.67 0.46 0.57
CRS-$v_{\mathrm{FID}}$ 25.50 21.41 0.48 0.56
CRS-$v_{\varepsilon}$ 26.97 20.71 0.47 0.56
CRS-$v_{x}$ 25.83 19.92 0.47 0.57
CRS-$v_{x}$ Linear 25.83 19.92 0.47 0.57
EDM 49.72 50.08 0.31 0.51
CRS-$v_{\mathrm{FID}}$ 20.85 13.01 0.52 0.56
CRS-$v_{\varepsilon}$ 20.43 13.14 0.53 0.58
CRS-$v_{x}$ 17.88 10.14 0.55 0.59
SDE-DPM- Linear Linear 11.12 5.68 0.62 0.62
Solver++(2M) VDM++ 29.00 17.67 0.46 0.57
CRS-$v_{\mathrm{FID}}$ 10.22 5.81 0.62 0.64
CRS-$v_{\varepsilon}$ 11.18 5.84 0.62 0.63
CRS-$v_{x}$ 10.25 5.64 0.62 0.64
CRS-$v_{x}$ Linear 10.25 5.64 0.62 0.64
EDM 11.52 6.13 0.60 0.64
CRS-$v_{\mathrm{FID}}$ 10.49 5.39 0.63 0.63
CRS-$v_{\varepsilon}$ 10.24 5.22 0.63 0.63
CRS-$v_{x}$ 10.56 5.32 0.63 0.63
DDIM Linear Linear 17.33 6.79 0.55 0.62
VDM++ 19.53 5.97 0.54 0.64
CRS-$v_{\mathrm{FID}}$ 15.52 6.95 0.56 0.63
CRS-$v_{\varepsilon}$ 17.02 6.94 0.55 0.63
CRS-$v_{x}$ 16.28 6.27 0.56 0.64
CRS-$v_{x}$ Linear 16.28 6.27 0.56 0.64
EDM 22.17 9.22 0.49 0.62
CRS-$v_{\mathrm{FID}}$ 15.51 6.72 0.56 0.63
CRS-$v_{\varepsilon}$ 15.26 5.79 0.57 0.64
CRS-$v_{x}$ 15.28 5.67 0.57 0.64
PNDM Linear Linear 13.21 5.35 0.59 0.65
VDM++ 16.01 6.31 0.57 0.66
CRS-$v_{\mathrm{FID}}$ 11.42 5.31 0.59 0.66
CRS-$v_{\varepsilon}$ 12.69 5.45 0.59 0.65
CRS-$v_{x}$ 12.29 5.38 0.59 0.66
CRS-$v_{x}$ Linear 12.29 5.38 0.59 0.66
EDM 12.86 5.62 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 12.18 5.21 0.59 0.65
CRS-$v_{\varepsilon}$ 12.57 5.24 0.59 0.66
CRS-$v_{x}$ 12.65 5.38 0.59 0.66
DPM- Linear Linear 13.87 5.41 0.59 0.64
Solver++(2M) VDM++ 16.97 6.45 0.56 0.66
CRS-$v_{\mathrm{FID}}$ 12.09 5.40 0.59 0.65
CRS-$v_{\varepsilon}$ 13.41 5.54 0.58 0.65
CRS-$v_{x}$ 12.92 5.46 0.59 0.66
CRS-$v_{x}$ Linear 12.92 5.46 0.59 0.66
EDM 13.67 5.79 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 13.02 5.19 0.59 0.65
CRS-$v_{\varepsilon}$ 13.02 5.17 0.59 0.65
CRS-$v_{x}$ 13.33 5.13 0.59 0.65
Table 22: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM 50 Linear 15.61 8.97 0.57 0.60
EDM 25.66 19.17 0.46 0.57
CRS-$v_{\mathrm{FID}}$ 13.67 6.86 0.60 0.60
CRS-$v_{\varepsilon}$ 13.44 7.56 0.59 0.61
CRS-$v_{x}$ 13.05 6.45 0.60 0.61
30 Linear 25.83 19.92 0.47 0.57
EDM 49.72 50.08 0.31 0.51
CRS-$v_{\mathrm{FID}}$ 20.85 13.01 0.52 0.56
CRS-$v_{\varepsilon}$ 20.43 13.14 0.53 0.58
CRS-$v_{x}$ 17.88 10.14 0.55 0.59
20 Linear 44.15 42.96 0.34 0.53
EDM 78.05 99.25 0.20 0.44
CRS-$v_{\mathrm{FID}}$ 30.28 24.26 0.44 0.54
CRS-$v_{\varepsilon}$ 31.86 26.34 0.43 0.55
CRS-$v_{x}$ 27.20 20.09 0.47 0.56
SDE-DPM- 50 Linear 9.96 5.62 0.63 0.64
DPM-Solver++(2M) EDM 9.72 5.35 0.63 0.64
CRS-$v_{\mathrm{FID}}$ 10.26 5.43 0.63 0.63
CRS-$v_{\varepsilon}$ 9.68 5.33 0.63 0.64
CRS-$v_{x}$ 10.11 5.32 0.63 0.64
30 Linear 10.25 5.64 0.62 0.64
EDM 11.52 6.13 0.60 0.64
CRS-$v_{\mathrm{FID}}$ 10.49 5.39 0.63 0.63
CRS-$v_{\varepsilon}$ 10.24 5.22 0.63 0.63
CRS-$v_{x}$ 10.56 5.32 0.63 0.63
20 Linear 11.47 5.81 0.61 0.63
EDM 16.26 9.38 0.55 0.63
CRS-$v_{\mathrm{FID}}$ 11.10 5.50 0.63 0.62
CRS-$v_{\varepsilon}$ 10.83 5.36 0.62 0.63
CRS-$v_{x}$ 11.64 5.26 0.63 0.62
Table 23: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DDIM 50 Linear 13.91 5.37 0.58 0.65
EDM 16.25 6.24 0.55 0.65
CRS-$v_{\mathrm{FID}}$ 13.45 5.36 0.58 0.64
CRS-$v_{\varepsilon}$ 13.64 5.20 0.58 0.65
CRS-$v_{x}$ 13.75 5.12 0.59 0.65
30 Linear 16.28 6.27 0.56 0.64
EDM 22.17 9.22 0.49 0.62
CRS-$v_{\mathrm{FID}}$ 15.51 6.72 0.56 0.63
CRS-$v_{\varepsilon}$ 15.26 5.79 0.57 0.64
CRS-$v_{x}$ 15.28 5.67 0.57 0.64
20 Linear 20.49 8.82 0.51 0.62
EDM 33.08 15.78 0.40 0.60
CRS-$v_{\mathrm{FID}}$ 18.43 7.52 0.54 0.63
CRS-$v_{\varepsilon}$ 18.30 7.57 0.54 0.63
CRS-$v_{x}$ 18.16 7.05 0.55 0.62
PNDM 50 Linear 12.10 5.32 0.59 0.66
EDM 12.53 5.49 0.59 0.66
CRS-$v_{\mathrm{FID}}$ 11.86 5.20 0.59 0.65
CRS-$v_{\varepsilon}$ 12.35 5.28 0.59 0.66
CRS-$v_{x}$ 12.44 5.28 0.59 0.66
30 Linear 12.29 5.38 0.59 0.66
EDM 12.86 5.62 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 12.18 5.21 0.59 0.65
CRS-$v_{\varepsilon}$ 12.57 5.24 0.59 0.66
CRS-$v_{x}$ 12.65 5.38 0.59 0.66
20 Linear 12.83 5.52 0.59 0.66
EDM 13.99 5.99 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 12.97 5.38 0.58 0.66
CRS-$v_{\varepsilon}$ 12.92 5.30 0.59 0.65
CRS-$v_{x}$ 13.63 6.10 0.57 0.66
DPM- 50 Linear 12.42 5.41 0.59 0.66
Solver++(2M) EDM 12.90 5.57 0.59 0.66
CRS-$v_{\mathrm{FID}}$ 12.27 5.15 0.59 0.65
CRS-$v_{\varepsilon}$ 12.64 5.23 0.59 0.66
CRS-$v_{x}$ 12.75 5.21 0.59 0.65
30 Linear 12.92 5.46 0.59 0.66
EDM 13.67 5.79 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 13.02 5.19 0.59 0.65
CRS-$v_{\varepsilon}$ 13.02 5.17 0.59 0.65
CRS-$v_{x}$ 13.33 5.13 0.59 0.65
20 Linear 13.71 5.45 0.58 0.65
EDM 15.31 6.23 0.57 0.66
CRS-$v_{\mathrm{FID}}$ 14.10 5.25 0.59 0.65
CRS-$v_{\varepsilon}$ 13.70 5.13 0.59 0.65
CRS-$v_{x}$ 14.16 5.09 0.59 0.64
Table 24: Evaluation results on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM. Bold values indicate best sampling schedules for each combination of sampler and NFE. FID scores on AYS is taken from their original paper [16].
Sampling FID $\downarrow$
Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) EDM 5.57 3.02 2.58
AYS 5.43 3.29 2.62
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
iPNDM EDM 4.60 2.62 2.49
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.87 2.53 2.44
Table 25: Evaluation results for CIFAR10 32 $\times$ 32 on class-conditional image generation using the pretrained model provided in EDM. Bold values indicate best sampling schedules for each combination of sampler and NFE.
Sampling FID $\downarrow$
Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) EDM 4.12 2.20 1.94
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.03 2.07 1.89
iPNDM EDM 2.51 1.93 1.89
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.33 1.92 1.89
Figure 14: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=250$ ($\mathrm{FID}=2.03$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. SDE-DPM-Solver++(2M) was used as sampler.
Figure 15: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=50$ ($\mathrm{FID}=2.04$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. PNDM was used as sampler.
Figure 16: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=10$ ($\mathrm{FID}=3.30$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. UniPC was used as sampler.
Figure 17: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=5$ ($\mathrm{FID}=14.02$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. DPM-Solver++(2M) was used as sampler.
Figure 18: Samples of LSUN Church 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=3.89$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. DPM-Solver++(2M) was used as sampler.
Figure 19: Samples of LSUN Bedroom 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=2.71$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. PNDM was used as sampler.
Figure 20: Samples of ImageNet 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ with class condition ($\mathrm{FID}=10.83$). Six classes were randomly selected, and each row corresponds to one class. We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. SDE-DPM-Solver++(2M) was used as sampler.

### v4

@@
-Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models
+Constant Rate Scheduling: A General Framework for Optimizing Diffusion Noise Schedule via Distributional Change
 Abstract
-We propose a general approach to optimize noise schedules for training and sampling in diffusion models.
-Our approach optimizes the noise schedules to ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process.
-Any distance metric for measuring the probability-distributional change is applicable to our approach, and we introduce three distance metrics.
-We evaluated the effectiveness of our approach on unconditional and class-conditional image-generation tasks using the LSUN (Horse, Bedroom, Church), ImageNet, FFHQ, and CIFAR10 datasets.
-Through extensive experiments, we confirmed that our approach broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations ranging from 5 to 250.
-Notably, by using our approach for optimizing both training and sampling schedules, we achieved a state-of-the-art FID score of 2.03 without sacrificing mode coverage on LSUN Horse 256 $\times$ 256.
+We propose a general framework for optimizing noise schedules in diffusion models, applicable to both training and sampling.
+Our method enforces a constant rate of change in the probability distribution of diffused data throughout the diffusion process, where the rate of change is quantified using a user-defined discrepancy measure.
+We introduce three such measures, which can be flexibly selected or combined depending on the domain and model architecture.
+While our framework is inspired by theoretical insights, we do not aim to provide a complete theoretical justification of how distributional change affects sample quality.
+Instead, we focus on establishing a general-purpose scheduling framework and validating its empirical effectiveness.
+Through extensive experiments, we demonstrate that our approach consistently improves the performance of both pixel-space and latent-space diffusion models, across various datasets, samplers, and a wide range of number of function evaluations from 5 to 250.
+In particular, when applied to both training and sampling schedules, our method achieves a state-of-the-art FID score of 2.03 on LSUN Horse $\times256$, without compromising mode coverage.
 1 Introduction
-Diffusion models are a type of probabilistic generative model that have gained significant attention for their expressive ability to generate high-quality samples across various domains, such as image generation image_generation1 ; image_generation2 ; image_generation3 , image super-resolution image_super_resolution1 ; image_super_resolution2 , video synthesis video_generation1 ; video_generation2 , and audio generation audio_generation1 ; audio_generation2 .
-Diffusion models were originally proposed in the previous study nonequiliburium_thermodynamics , and denoising diffusion probabilistic models (DDPMs) ddpm_original demonstrate that diffusion models can achieve state-of-the-art performance in image generation.
-In diffusion models, two diffusion processes are defined using Markov chains, referred to as the forward process and reverse process.
-In the forward process, data are diffused by gradually adding Gaussian noise, until the probability distribution of diffused data becomes a Gaussian distribution.
-In the reverse process, data are generated from Gaussian noise by tracing the forward process in the reverse direction.
-Noise levels added to data at each step of the diffusion processes are hyperparameters and referred to as the noise schedule.
-We do not necessarily have to use a same noise schedule for training and sampling, and recent studies edm ; vdmpp have shown that using different noise schedules can lead to improved performance.
-Although diffusion models tend to perform well in terms of sample fidelity and mode coverage, they face the challenge of longer generation times due to the iterative function evaluations during sampling diffusion_gans .
-To improve the performance of diffusion models, we propose general approach for optimizing training and sampling schedules that ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process.
-We call our approach constant rate scheduling (CRS).
-As previously stated, the reverse process needs to accurately trace the forward process in the reverse direction to restore the data from pure noise.
-This motivates us to determine noise schedules that can enhance the traceability of the forward process.
-Figure 1 illustrates a toy example of diffused data distributions with three data points in one-dimensional data space (see Appendix A for details).
+Diffusion models are a class of probabilistic generative models that have attracted significant attention due to their ability to generate high-quality samples across a variety of domains, including image generation (Saharia et al., 2022b; Nichol et al., 2022b; Podell et al., 2024a), image super-resolution (Saharia et al., 2023; Doi et al., 2024), video synthesis (Ho et al., 2022b; a), and audio generation (Kong et al., 2021; Popov et al., 2021).
+Diffusion models were originally introduced in Sohl-Dickstein et al.
+(2015), and denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020) demonstrated that diffusion models can achieve state-of-the-art performance in image generation.
+A typical diffusion model consists of two Markovian processes: a forward process and a reverse process.
+The forward process gradually adds Gaussian noise to the input data, transforming its distribution toward a Gaussian.
+The reverse process aims to reconstruct data from Gaussian noise by tracing the forward process in the reverse direction.
+The noise levels applied at each step of these processes are treated as hyperparameters and are collectively referred to as the noise schedule.
+The noise schedule has a critical impact on both the quality and efficiency of training and sampling (Chen, 2023).
+Importantly, the schedules used during training and sampling do not have to be identical, and recent studies have shown that using different schedules for training and sampling can improve performance (Karras et al., 2022; Kingma and Gao, 2023; Hang et al., 2025).
+However, designing effective noise schedules remains a non-trivial and largely empirical process.
+In this work, we propose a general framework for optimizing both training and sampling schedules by enforcing a constant rate of change in the probability distribution of diffused data throughout the diffusion process.
+We refer to this framework as constant rate scheduling (CRS).
+Since the reverse process is expected to accurately trace the forward process, the key motivation is to determine noise schedules that can enhance the traceability of the forward process .
+Figure 1 presents a toy example of diffused data distributions with three data points in one-dimensional data space (see Appendix A for details).
 Here, $\alpha$ represents the decay rate of data at each step of the diffusion process.
-In this example, the diffusion process can be divided into three regions.
-The first is an area where the probability distributions barely change, and the noise level can be changed rapidly ($\alpha\lesssim 0.6$).
-The second is an area where three modes corresponding to the data points emerge, and careful adjustment of the noise level is necessary to enhance the mode coverage ($0.6\lesssim\alpha\lesssim 0.97$).
-The third is an area where the peaks of each mode become distinct, and the noise level should be changed slowly for sample fidelity ($\alpha\gtrsim 0.97$).
-As evident from this example, the diffusion process exhibits regions with rapid and slow changes in the probability distributions.
-We assume that the probability-distributional change of diffused data represents the traceability of the diffusion process and determines the noise schedule to ensure a constant rate of change in the probability distributions.
-CRS can be considered general from the following perspectives: 1) CRS can optimize both the training and sampling schedules, 2) any distance metric for measuring probability-distributional change is applicable, allowing the introduction of domain specific distance metrics and the use of different noise schedules for training and sampling, 3) any noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
-Our contributions are summarized as follows:
-  1. 1.
-    We propose a general approach to optimize noise schedules for both training and sampling.
-We can use any distance metric for measuring probability-distributional change and introduce three distance metrics.
-One is specialized for image generation, while the other two are expected to be applicable across a wide range of domains beyond image generation.
-  2. 2.
-    Through extensive experiments, we confirm that CRS broadly improves the performance of image generation in pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations (NFEs) from 5 to 250.
-CRS contributes to the performance improvement at both training and sampling, achieving a state-of-the-art Fréchet inception distance (FID) fid_score score of 2.03 on LSUN Horse 256 $\times$ 256 dataset.
-CRS is fundamentally different from conventional approaches, such as align your steps (AYS) noise_ays and geometry-inspired time scheduling (GITS) noise_gits , in that it can optimize the training schedule as well.
+In this example, the diffusion process can be divided into three regions: 1) where distributions barely change and noise can be decreased rapidly ($\alpha\lesssim 0.6$), 2) where modes corresponding to each data point emerge and require fine control of noise to avoid mode dropping ($0.6\lesssim\alpha\lesssim 0.97$), and 3) where mode peaks become sharp and noise must be reduced gradually for better fidelity ($\alpha\gtrsim 0.97$).
+This example illustrates that the distributional dynamics of the diffusion process are highly non-uniform.
+We hypothesize that the probability-distributional change represents the traceability of the diffusion process, and controlling the rate of distributional change can improve the quality and stability of the generative process.
+Recent advances in generative modeling have also emphasized the importance of well-behaved distributional evolution.
+For example, flow matching (Lipman et al., 2023) and rectified flow transformers (Esser et al., 2024) formulate generation as a deterministic flow between data and noise distributions, highlighting that stable distributional trajectories can lead to improved sample quality.
+Although these approaches differ from diffusion models in their formulation, they share the high-level motivation of designing generative processes with controlled and stable distributional dynamics.
 Figure 1: Toy example of diffused data distributions with three data points in one-dimensional data space.
@@
 Three modes become distinct when $\alpha\gtrsim 0.97$, requiring careful control of noise level for sample fidelity.
+CRS offers a unified and flexible framework for designing noise schedules, with the following key properties: 1) it supports both training and sampling schedules, 2) it allows the use of arbitrary discrepancy measures to quantify distributional change, enabling domain-specific tailoring and independent schedules for training and sampling, and 3) many existing schedules can be viewed as special cases of CRS under different discrepancy measures.
+Our contributions are summarized as follows:
+  1. 1.
+    We propose a general framework for optimizing noise schedules in diffusion models by enforcing a constant rate of distributional change throughout the diffusion process.
+Our framework supports both training and sampling schedules and allows the use of any user-specified discrepancy measure.
+  2. 2.
+    We introduce three practical discrepancy measures for quantifying distributional change, each capturing different aspects of the diffusion process.
+These measures can be flexibly selected or combined, depending on the domain or model type (e.g., pixel-space or latent-space diffusion models).
+  3. 3.
+    While theoretically motivated, this work does not aim to provide a full formal analysis of the link between distributional change and sample quality.
+Instead, we focus on empirically validating the effectiveness of CRS across diverse experimental settings.
+Through extensive experiments, we demonstrate that CRS consistently improves the performance of both pixel-space and latent-space diffusion models across various datasets, samplers, and numbers of function evaluations (NFEs), ranging from 5 to 250.
+The remainder of this paper is organized as follows: Sec.
+2 reviews related work.
+Sec.
+3 briefly introduces DDPMs.
+Sec.
+4 presents our proposed framework.
+Sec.
+5 provides the experimental results.
+Sec.
+6 discusses limitations, and Sec.
+7 concludes the paper.
 2 Related work
-Recent studies regarding fast sampling can be mainly categorized into the following three strategies.
-Samplers: This strategy devises update rules of diffused data in the reverse process.
-Denoising diffusion implicit models (DDIMs) ddim_original introduce a non-Markov chain to generalize DDPMs and propose a deterministic sampler.
-In the continuous-time diffusion models, the forward and reverse processes are formulated using a stochastic differential equation (SDE) and ordinary differential equation (ODE) score_matching3 .
-Numerous studies have been conducted to develop efficient solvers tailored for the reverse SDE and ODE in diffusion models sampler_adaptive_step ; sampler_pndm ; sampler_dpm_solver ; sampler_dpmpp ; sampler_ipndm ; sampler_unipc ; sampler_dpm-v3 ; sampler_amed .
-Noise schedules: Noise schedules significantly affect the performance of diffusion models noise_effect , and many noise schedules have been proposed ddpm_original ; improved_diffusion ; noise_effect ; shifted_cosine .
-However, the functional form in these studies, such as the linear ddim_original and cosine improved_diffusion functions, is determined somewhat arbitrarily.
-Recently, systematic optimization of noise schedules have begun to be actively studied noise_gits ; noise_jys .
-The most relevant approach to CRS is AYS noise_ays , which determines a noise schedule for sampling to minimize the discretization errors when solving the reverse SDE.
-However, noise schedules aimed at improving the training efficiency have been largely overlooked compared with the sampling efficiency, though several studies, such as elucidating diffusion models (EDMs) edm and variational diffusion models++ (VDM++) vdmpp , have investigated them.
-Distillations: The distillation-based acceleration is one of the most important direction for fast sampling distill_org1 ; distill_org2 ; distill_consistency ; distill_dmd ; distill_dmd2 ; distill_s4s .
-CRS is expected to provide better teacher models for distillation methods because it can optimize the training schedule as well.
+The performance of generative models is typically evaluated along three axes: sampling speed, fidelity, and mode coverage.
+Diffusion models are known to excel in terms of fidelity and mode coverage, but suffer from slow sampling speed due to the iterative nature of the reverse process (Xiao et al., 2022).
+To address this limitation, a wide range of acceleration strategies have been proposed.
+These approaches can be broadly categorized into five groups:
+Conditional generation: This strategy improves sampling speed by reducing sample diversity through conditioning (e.g., class labels or text prompts) (Dhariwal and Nichol, 2021; Ho and Salimans, 2021; Rombach et al., 2022; Saharia et al., 2022a; Nichol et al., 2022a; Podell et al., 2024b), which guides the model to generate specific outputs and avoid unnecessary exploration.
+Dimensionality reduction: Latent diffusion models (LDMs) (Rombach et al., 2022) employ pre-trained autoencoders to compress data representations, allowing the diffusion process to operate in a lower-dimensional latent space.
+This improves both training and sampling efficiency.
+Samplers: These methods improve the update rule of diffused data for the reverse process.
+Denoising diffusion implicit models (DDIMs) (Song et al., 2021a) introduce a non-Markovian process to generalize DDPMs and propose a deterministic sampler.
+In continuous-time diffusion models, the forward and reverse processes are formulated as stochastic differential equations (SDEs) and ordinary differential equations (ODEs) (Song et al., 2021b).
+Numerous studies have developed efficient solvers tailored to these formulations (Jolicoeur-Martineau et al., 2021; Liu et al., 2022; Lu et al., 2022a; b; Zhang and Chen, 2023; Zhao et al., 2023; Zheng et al., 2023a; Zhou et al., 2024).
+Noise schedules: The choice of noise schedule significantly affects training and sampling efficiency.
+Early schedules, such as linear (Ho et al., 2020) and cosine (Nichol and Dhariwal, 2021), were chosen heuristically.
+More recently, principled approaches for optimizing sampling schedules have been proposed (Chen et al., 2024; Sabour et al., 2024; Williams et al., 2024; Park et al., 2025; Tong et al., 2025).
+Among them, Align Your Steps (AYS) (Sabour et al., 2024) determines a noise schedule to minimize the discretization errors measured using Kullback-Leibler divergence.
+One of the discrepancy measures in our approach can be interpreted as a simplified variant of Kullback-Leibler divergence upper bound (KLUB) introduced in AYS, and thus CRS generalizes AYS and related methods as special cases.
+In contrast, elucidating diffusion models (EDMs) (Karras et al., 2022) and variational diffusion models++ (VDM++) (Kingma and Gao, 2023) have proposed optimization of the training schedule.
+These works also highlight that using different schedules for training and sampling can improve performance.
+However, most prior work has focused predominantly on sampling schedules, and optimization of training schedules remains relatively underexplored.
+CRS is designed to optimize both schedules within a unified framework, supporting the use of arbitrary and even composite discrepancy measures.
+This flexibility allows practitioners to adapt scheduling strategies based on empirical behavior and task-specific requirements.
+Distillation: Distillation-based approaches are another major direction for fast sampling (Luhman and Luhman, 2021; Salimans and Ho, 2022; Song et al., 2023; Yin et al., 2024b; a; Frankel et al., 2025).
+These methods compress multi-step diffusion into fewer steps.
+CRS is orthogonal to this line of work and can provide better teacher models for distillation, as it can improve training schedules.
 3 Background
-We briefly introduce DDPMs.
-In DDPMs ddpm_original , the forward process, which starts from a data point $\bm{x}_{0}\sim q(\bm{x})$, is defined as
-$q(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})\right)q(\bm{x}_{0})\ \ \mathrm{and}\ \ q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\beta_{t}\bm{x}_{t-1},\delta_{t}^{2}\bm{I}),$ (1)
-where $\beta_{t}$ is the pre-defined decay rate of the data, $\delta_{t}$ represents the noise strength added to $\bm{x}_{t-1}$, and $T$ is the number of timesteps in the forward process.
-It is straightforward to verify that
-$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I}),$ (2)
-where $\alpha_{t}=\beta_{t}\alpha_{t-1}$ and $\sigma_{t}^{2}=\delta_{t}^{2}+\beta_{t}^{2}\sigma_{t-1}^{2}$.
-We focus on the variance-preserving process, where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ and $\beta_{t}^{2}+\delta_{t}^{2}=1$ are satisfied.
-The reverse process starts from Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$.
-In the limit of $T\to\infty$, the probability distribution of the reverse process is well approximated by the Gaussian distribution ddpm_approx :
-$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\nu_{t}^{2}\bm{I}).$ (3)
-Using this approximation, the noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},t)$:
-$\bm{\mu}_{\theta}(\bm{x},t)=\frac{1}{\beta_{t}}\left(\bm{x}_{t}-\frac{\delta_{t}^{2}}{\sigma_{t}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)\right),$ (4)
-is optimized using a simplified loss function derived from the variational lower bound.
+We briefly introduce DDPMs (Ho et al., 2020).
+In the forward process, Gaussian noise is gradually added to the data through a sequence of timesteps, transforming the data distribution into an isotropic Gaussian.
+This process is typically modeled as a Markov chain that starts from a data sample $\bm{x}_{0}\sim q(\bm{x}_{0})$:
+$q(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})\right)q(\bm{x}_{0}),\quad\mathrm{where}\quad q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\beta_{t}\bm{x}_{t-1},\delta_{t}^{2}\bm{I}).$ (1)
+Here, $T$ denotes the number of timesteps, $\beta_{t}$ is a decay factor controlling the signal strength, $\delta_{t}$ represents the added noise level, and $q(\bm{x}_{0})$ is the data distribution.
+It is straightforward to show that $q(\bm{x}_{t}|\bm{x}_{0})$ takes the following form:
+$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$ (2)
+where the coefficients satisfy the recursive relations: $\alpha_{t}=\beta_{t}\alpha_{t-1}$ and $\sigma_{t}^{2}=\delta_{t}^{2}+\beta_{t}^{2}\sigma_{t-1}^{2}$.
+This formulation allows sampling of $\bm{x}_{t}$ from $\bm{x}_{0}$ in a single step.
+In this work, we parameterize the forward process using $(\alpha_{t},\sigma_{t})$ instead of $(\beta_{t},\delta_{t})$, and adopt a variance-preserving formulation in which both $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ and $\beta_{t}^{2}+\delta_{t}^{2}=1$ hold.
+The sequence $\{\alpha_{t}\}_{t=0}^{T}$ defines the noise schedule, which controls how rapidly the signal component of the data decays during diffusion.
+Our goal is to optimize this noise schedule to enable more effective and efficient training and sampling in diffusion models.
+In the reverse process, data samples are generated from Gaussian noise by tracing the forward process in the reverse direction. This process is also modeled as a Markov chain and is defined as:
+$p_{\theta}(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})\right)p(\bm{x}_{T}),\quad\mathrm{where}\quad p(\bm{x}_{T})=\mathcal{N}(\bm{x}_{T};\bm{0},\bm{I}).$ (3)
+Here, $\theta$ denotes the learnable parameters of the diffusion model. In the limit of $T\to\infty$, the reverse transitions can be well approximated by the Gaussian distributions (Feller, 1949):
+$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}\bigl(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\nu_{t}^{2}\bm{I}\bigr).$ (4)
+DDPMs employ a noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)$, which predicts the noise added at timestep $t$. The predicted mean $\bm{\mu}_{\theta}(\bm{x},t)$ is given by:
+$\bm{\mu}_{\theta}(\bm{x},t)=\frac{1}{\beta_{t}}\left(\bm{x}_{t}-\frac{\delta_{t}^{2}}{\sigma_{t}}\bm{\varepsilon}_{\theta}(\bm{x},t)\right).$ (5)
+The noise prediction model is trained by maximizing a simplified version of the variational lower bound, which yields a practical and stable learning objective.
 4 General Approach for Optimizing Noise Schedule
-We explain our motivation and present CRS in more detail.
+We explain our motivation and present CRS.
 4.1 Motivation
 As stated in Sec.
-3, when $T$ is sufficiently large, it is a good approximation to assume the probability distribution of the reverse process $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ as a Gaussian distribution, as expressed in Eq.
+3, when $T$ is sufficiently large, it is a good approximation to assume that the reverse transitions $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ follow a Gaussian distribution, as expressed in Eq.
 (4).
-In other words, the distributional change between adjacent timesteps must be small enough to be well approximated by a Gaussian distribution.
-This observation motivates us to minimize the maximum distributional change in the diffusion process for holding the approximation in Eq.
-(3) throughout the diffusion process.
-We can understand the importance of the probability-distributional change in the diffusion process from the perspective of noise conditional score networks (NCSNs) score_matching1 ; score_matching2 ; score_matching3 , which are also the origin of diffusion models.
-According to the manifold hypothesis, natural data typically lie on a low-dimensional manifold embedded within a high-dimensional data space.
-Based on this property, there exist two main challenges in generating samples using only the score of the data distribution $\nabla_{\bm{x}}\log q(\bm{x})$ starting from a random initialization $\bm{x}_{\mathrm{rnd}}$ score_matching1 .
-First, $\bm{x}_{\mathrm{rnd}}$ is likely to be located in regions with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, where it is difficult to learn the score accurately.
-Second, even if the score can be learned correctly, the score tends to be small in the region with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, leading to slow mixing.
-To address these challenges, NCSNs interpolate between the data distribution and a Gaussian distribution, the support of which is the whole space.
-If the distributional change between adjacent timesteps is sufficiently small during the interpolation process, a sample generated at a certain timestep is likely to fall into a region where the probability density of the adjacent timestep is high.
-Therefore, by minimizing the maximum distributional change between adjacent timesteps, it becomes possible to generate samples using the reliable score at all timesteps of the interpolation process.
+This implies that the distributional change between adjacent timesteps must be small enough to be well approximated by a Gaussian distribution.
+In diffusion models, sample quality depends on how faithfully the reverse process can trace the forward process.
+Large distributional changes can lead to poor approximation, resulting in artifacts or mode dropping.
+This observation motivates us to minimize the maximum distributional change during the diffusion process, in order to maintain the validity of Eq.
+(4) throughout the trajectory.
+The importance of controlling distributional change can also be understood from the perspective of noise conditional score networks (NCSNs) (Song and Ermon, 2019; 2020; Song et al., 2021b), which are the conceptual predecessors of diffusion models.
+According to the manifold hypothesis, natural data typically lie on a low-dimensional manifold embedded in high-dimensional space.
+This leads to two main challenges when generating samples using only the score function $\nabla_{\bm{x}}\log q(\bm{x})$ from a randomly initialized point $\bm{x}_{\mathrm{rnd}}$ (Song and Ermon, 2019): first, $\bm{x}_{\mathrm{rnd}}$ is likely to fall in regions where $q(\bm{x}_{\mathrm{rnd}})\ll 1$, making it difficult to estimate accurate gradients; second, even if the score is correctly estimated, it tends to be small in such regions, leading to slow mixing.
+To address these challenges, NCSNs interpolate between the data distribution and a Gaussian distribution, whose support spans the entire space.
+If the distributional change between adjacent steps is kept small during this interpolation, each reverse step is likely to remain within regions of high probability density.
+This improves the reliability of local updates and enhances overall sample quality.
+These theoretical motivations form the basis for our proposed framework: to design noise schedules that ensure a constant rate of change in the distribution of the diffused data across timesteps.
 4.2 Constant Rate Scheduling
-As described above, our criterion to optimize noise schedules is to minimize the maximum distance between consecutive probability distributions in the diffusion process.
-Using the distance $D(t,t+\Delta t)$ between the probability distributions of diffused data $p(\bm{x}_{t})$ and $p(\bm{x}_{t+\Delta t})$, our criterion is expressed as
+As described above, our goal is to optimize noise schedules by minimizing the maximum distributional change between adjacent timesteps in the diffusion process.
+Let $D(t,t+\Delta t)$ denote a discrepancy measure between the distributions of diffused data $q(\bm{x}_{t})$ and $q(\bm{x}_{t+\Delta t})$.
+Our criterion can be formulated as:
 $\underset{\alpha(t)}{\operatorname{argmin}}\left(\max_{t}D(t,t+\Delta t)\right).$ (6)
-This optimization problem has a straightforward solution as follows:
+The optimal schedule $\alpha(t)$ must satisfy the following condition:
 $D(t,t+\Delta t)=\frac{D(t,t+\Delta t)-D(t,t)}{\Delta t}\Delta t\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=\mathrm{const.},$ (7)
-where we use $D(t,t)=0$.
-To explicitly show the dependence on the noise schedule, we rewrite $D(t,t^{\prime})$ as $\tilde{D}(\alpha(t),\alpha(t^{\prime}))\equiv\tilde{D}(\alpha,\alpha^{\prime})$.
-Here, $\alpha(t)$ is the noise-schedule function to be optimized.
-This results in
-$D(t,t+\Delta t)\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=-v(\alpha)\frac{d\alpha(t)}{dt}\Delta t=\mathrm{const.},$ (7)
-where $v(\alpha)=\left.\frac{\partial\tilde{D}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}$ represents the rate of change in the probability distributions of diffused data.
-Therefore, the noise-schedule function is determined to satisfy $-\frac{d\alpha(t)}{dt}\propto v(\alpha)^{-1}$.
-The proportionality coefficient can be calculated using the boundary conditions: $\alpha(0)=\alpha_{\max}$ and $\alpha(1)=\alpha_{\min}$.
-Finally, we obtain the noise-schedule function as follows:
+where we used $D(t,t)=0$.
+To explicitly reflect the dependence on the noise schedule, we rewrite $D(t,t^{\prime})$ as $\tilde{D}(\alpha(t),\alpha(t^{\prime}))\equiv\tilde{D}(\alpha,\alpha^{\prime})$, where $\alpha(t)$ is the noise schedule to be optimized.
+This leads to:
+$D(t,t+\Delta t)\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=v(\alpha)\frac{d\alpha(t)}{dt}\Delta t=\mathrm{const.},$ (8)
+where $v(\alpha)$ represents the rate of change in the probability distribution of the diffused data.
+Therefore, the noise schedule must satisfy $\frac{d\alpha(t)}{dt}\propto v(\alpha)^{-1}$.
+The proportionality constant is computed from the boundary conditions $\alpha(0)=\alpha_{\max}$ and $\alpha(1)=\alpha_{\min}$, yielding:
+$\frac{d\alpha(t)}{dt}=Cv(\alpha)^{-\xi},\quad\mathrm{where}\quad C=\int_{\alpha_{\min}}^{\alpha_{\max}}v(\alpha)^{\xi}d\alpha.$ (9)
+Here, we introduce a hyperparameter $\xi>0$ to control the dependence of $\alpha(t)$ on $v(\alpha)$.
+Larger values of $\xi$ allocate more timesteps to regions where $v(\alpha)$ is large, i.e., where the distribution changes rapidly.
+The noise schedule optimization using CRS consists of the following steps:
+  1. 1.
+    Choose a discrepancy measure to quantify distributional change.
+  2. 2.
+    Compute $v(\alpha)$ using the selected discrepancy measure.
+  3. 3.
+    Solve Eq. (9) to obtain $\alpha(t)$.
+  4. 4.
+    Apply the resulting schedule as follows:
+    • For training, use $\alpha(t)$ directly.
+    • For sampling, discretize the schedule uniformly in time: $\{\alpha(t/T)|t=0,1,...,T\}$ where $T$ is the number of sampling steps.
+We propose three discrepancy measures for computing $v(\alpha)$, which are detailed in Section 4.3.
+The derivation of Eqs.
+(7) and (8) relies on a first-order Taylor expansion with respect to $\Delta t$, which is accurate when the NFE is sufficiently large.
+However, when optimizing the sampling schedule at small NFEs, $\Delta t$ becomes large and the approximation error may no longer be negligible.
+Empirically, we observe that CRS with $\xi=1$ remains effective for moderate NFEs, while the importance of tuning $\xi$ increases as the NFE decreases (see Table 27 in Appendix F.2).
+We attribute this increased sensitivity to the growing discretization error of the first-order expansion.
+For small NFEs, the optimal value of $\xi$ tends to be greater than $1$, resulting in more timesteps being allocated to regions where $v(\alpha)$ is large.
+Since discretization error is expected to be more pronounced in regions with larger $v(\alpha)$, it is intuitively reasonable that values of $\xi>1$ become optimal when the NFE is small.
+Although introducing $\xi$ partially compensates for the discretization error, it does not fully resolve the mismatch between the continuous-time formulation and the discrete-time implementation.
+A fully principled solution would require directly optimizing the sampling schedule in the discrete-time domain, without relying on the continuous-time assumption underlying Eq.
+(7).
+We consider this an important direction for future work.
+Similar to NCSNs inspired by simulated annealing (Kirkpatrick et al., 1983), our formulation draws additional motivation from the adiabatic theorem (Morita and Nishimori, 2008) in quantum annealing (Kadowaki and Nishimori, 1998).
+Quantum annealing solves combinatorial optimization problems by interpolating between a target cost function and an initial one that has a trivial optimum.
+The optimization starts from the known optimum of the initial function and aims to trace the instantaneous optimums along the interpolation path.
+It is crucial to accurately trace the optimum of the instantaneous cost function, and the adiabatic theorem provides the traceability of the optimum at each step.
+Let $\alpha(t)$ be the weight of the target cost function and $1-\alpha(t)$ be the weight of the initial cost function.
+In this setting, the inverse traceability is expressed in the form of $\frac{d\alpha(t)}{dt}v(\alpha)$, which takes the same form as Eq.
+(8).
+Prior work has shown that minimizing the maximum value of this inverse traceability yields the optimal $\alpha(t)$ (Roland and Cerf, 2002), that ensures consistent traceability throughout the interpolation.
+In CRS, we regard the rate of probability-distributional change as the inverse traceability and minimize its maximum value to derive an optimal noise schedule.
+4.3 Discrepancy Measures
+Any discrepancy measure is applicable to CRS. We introduce three discrepancy measures.
+4.3.1 Discrepancy measure based on FID
+FID is a widely used metric for evaluating sample fidelity in the field of image generation.
+It computes the Fréchet distance (Heusel et al., 2017) between the feature distributions of real and generated images under a Gaussian assumption, where the features are extracted using the Inception-V3 network (Szegedy et al., 2016) trained on ImageNet classification.
+When used as a discrepancy measure in our framework, $v(\alpha)$ tends to take large values in the region where sample fidelity significantly degrades (i.e., $\alpha\simeq 1$).
+As a result, CRS allocates more timesteps to this region, thereby improving sample fidelity.
+However, a limitation of the FID-based measure is that it fails to accurately capture distributional changes in high-noise regions ($\alpha\simeq 0$), since the feature extractor is not trained on heavily diffused images and may not provide meaningful representations in that regime.
+Therefore, while the FID-based measure is effective for capturing fidelity-related changes in low-noise regions, it may benefit from being complemented by other discrepancy measures that can better capture high-noise dynamics.
+Once the forward process is defined, $v(\alpha)$ can be computed by simulating the forward diffusion using the training dataset (see Algorithm 1 in Appendix C.1).
+Since this procedure is independent of model parameters, it only needs to be executed once prior to training the diffusion model.
+The computation time of $v(\alpha)$ is proportional to the number of training images, which may be prohibitively expensive for large-scale training datasets.
+To address the scalability issue for large-scale datasets, we introduce a practical alternative in Sec.
+5.7.
+We refer to $v(\alpha)$ computed using FID as $v_{\mathrm{FID}}(\alpha)$.
+4.3.2 Discrepancy measure based on data prediction
+We introduce a discrepancy measure that leverages a trained diffusion model as follows:
+$v_{x}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}},\quad\mathrm{where}\quad\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (10)
+Here, $q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ is the joint distribution of diffused data at different noise levels, and $\bm{x}_{\theta}(\bm{x},\alpha)$ denotes the data prediction, which is related to the noise prediction $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ via $\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$.
+This measure is designed to allocate more computational resources to regions where the data prediction changes rapidly with respect to $\alpha$.
+As shown in Appendix B, $v_{x}(\alpha)$ can be interpreted as a simplified variant of KLUB, introduced in AYS.
+Importantly, $v_{x}(\alpha)$ can be used not only to determine the sampling schedule after training, but also to adaptively optimize the training schedule during learning.
+We further interpret $v_{x}(\alpha)$ as measuring the probability-distributional change of the diffused data.
+According to prior work (Ambrogioni, 2024), the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ can be viewed as a weighted average over training data points $\bm{x}_{0}$, where the weighting depends on the noise level $\alpha$.
+When $\alpha=0$, the weights are approximately uniform across the dataset.
+As $\alpha\rightarrow 1$, only a few data points contribute significantly, indicating sharper and more specific representations.
+The change in these implicit weights reflects how sample trajectories diverge as noise decreases, and thus $v_{x}(\alpha)$ provides a proxy for the rate of distributional change.
+The pseudocode for optimizing training and sampling schedules is provided in Appendix C.2.
+For sampling schedule optimization, the computation time of $v_{x}(\alpha)$ is negligible compared to the overall training time.
+In contrast, training schedule optimization requires updating $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ during training as the model parameters evolve, which increases the total training time by approximately 20–30%.
+However, our experiments show that the schedule typically stabilizes early in training.
+Therefore, fixing the schedule after a few epochs or reducing the update frequency of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ may mitigate the computational cost without compromising performance.
+Developing a more efficient implementation remains an important direction for future work.
+4.3.3 Discrepancy measure based on noise prediction
+In the experiments presented in Sec.
+5, we trained the noise prediction model.
+In this setting, it is also effective to use the noise predictions themselves to quantify distributional change, leading to the following definition:
+$v_{\varepsilon}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}},\quad\mathrm{where}\quad\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (11)
+We refer to CRS using $v_{\mathrm{FID}}(\alpha)$, $v_{x}(\alpha)$, and $v_{\varepsilon}(\alpha)$ as $v_{\mathrm{FID}}, CRS$-$v_{x}$, and CRS-$v_{\varepsilon}$, respectively, depending on the choice of discrepancy measure used to compute $v(\alpha)$.
+4.4 Combining Multiple Discrepancy Measures
+Each discrepancy measure has its own strengths and weaknesses.
+For example, CRS-$v_{\mathrm{FID}}$ effectively captures distributional changes relevant to sample fidelity, but fails to represent changes accurately in high-noise regions ($\alpha\simeq 0$), as the feature extractor is not trained on diffused images.
+On the other hand, CRS-$v_{x}$ captures distributional change even in high-noise regions, but it is not necessarily optimal for improving sample fidelity.
+Therefore, it is considered effective to complementarily combining multiple discrepancy measures for optimizing noise schedules.
+We introduce the following strategy for combining multiple discrepancy measures:
+$v(\alpha)=\sum_{m=1}^{M}w_{m}\tilde{v}_{m}(\alpha),\ \ \mathrm{where}\ \ \tilde{v}_{m}(\alpha)=\frac{1}{C_{m}}v_{m}(\alpha)^{\xi_{m}}\ \ \mathrm{and}\ \ C_{m}=\int_{\alpha_{\min}}^{\alpha_{\max}}v_{m}(\alpha)^{\xi_{m}}.$ (12)
+where $M$ is the number of discrepancy measures, $v_{m}(\alpha)$ is the rate of change computed using the $m$-th measure, $w_{m}$ is its corresponding weight (satisfying $\sum_{m=1}^{M}w_{m}=1$), and $\xi_{m}$ controls the emphasis on $v_{m}(\alpha)$.
+Since the value ranges of different discrepancy measures can vary significantly, we normalize each $v_{m}(\alpha)$ as in Eq.
+(12) before combining them.
+This formulation also enables conventional noise schedules to be interpreted and incorporated within the CRS framework.
+By converting a given noise schedule $\alpha(t)$ into its corresponding form of $v(\alpha)$, we can treat it as an implicit discrepancy measure.
+For example, if $\alpha(t)=\cos\left(\frac{\pi t}{2}\right)$, then:
+$v(\alpha)\propto-\left(\frac{d\alpha(t)}{dt}\right)^{-1}\propto\sin^{-1}\left(\frac{\pi t}{2}\right)=\frac{1}{\sqrt{1-\alpha^{2}}}.$ (13)
+This observation implies that any noise schedule can be viewed as the optimization result of CRS with an associated (possibly implicit) discrepancy measure, thus providing a unified perspective on the design of noise schedules.
+While combining multiple discrepancy measures increases the flexibility of CRS, it also introduces additional hyperparameters.
+In our experiments, we found that tuning these hyperparameters is required only when optimizing sampling schedules in pixel-space diffusion models.
+In all other cases—training schedules for both pixel-space and latent-space models, and sampling schedules for latent-space models—CRS-$v_{x}$ with $\xi=1$ consistently performed well and can serve as a default setting.
+In the next section, we show that CRS-$v_{x}+v_{\mathrm{FID}}$, which combines $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$, is effective for optimizing the sampling schedule of pixel-space diffusion models.
+CRS-$v_{x}+v_{\mathrm{FID}}$ has four hyperparameters: $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$.
+The procedure for tuning these hyperparameters is described in Appendix D, and the hyperparameter settings used in all experiments are summarized in Table 12.
+For configurations with $\mathrm{NFE}\leq 50$, Table 12 shows that $w_{x}=w_{\mathrm{FID}}=0.5$ is optimal in most cases.
+These observations suggest that, in practical use, it is generally sufficient to tune only $\xi_{x}$ and $\xi_{\mathrm{FID}}$.
+5 Experiments
+We experimentally demonstrate that CRS broadly improves the performance of diffusion models.
+5.1 Experimental Settings
+Below, we provide a brief overview of the components and settings used in our experiments.
+Dataset: We use six image datasets: LSUN (Church, Bedroom, Horse) (Yu et al., 2016), ImageNet (Deng et al., 2009), FFHQ (Karras et al., 2019), and CIFAR10 (Krizhevsky, 2009).
+LSUN, FFHQ, and CIFAR10 are used for unconditional image generation, while ImageNet is used for class-conditional image generation.
+Noise Prediction Model: We adopt the U-Net architecture introduced in the ablated diffusion model (ADM) (Dhariwal and Nichol, 2021) as the noise prediction model.
+Hyperparameter settings are listed in Table 13 (Appendix E).
+Training Schedule: We compare four training noise schedules: linear (Ho et al., 2020), shifted cosine (Hoogeboom et al., 2023), the adaptive schedule proposed in VDM++ (Kingma and Gao, 2023), and our proposed CRS.
+The CRS hyperparameter $\xi$ is fixed to $1$ in all experiments.
+Loss Function: Except for the adaptive schedule in VDM++, we use the simplified loss function from DDPMs:
+$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},t\sim\mathcal{U}(0,1),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x},\alpha(t))-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (14)
+where $\bm{x}=\alpha(t)\bm{x}_{0}+\sqrt{1-\alpha(t)^{2}}\bm{\varepsilon}$, $\mathcal{D}$ denotes the training dataset, and $\alpha(t)$ is the training schedule.
+When using the adaptive schedule from VDM++, the following loss function is used:
+$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\lambda\sim p(\lambda),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\frac{\omega(\lambda)}{p(\lambda)}\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right],\\$ (15)
+where $\lambda=\log\left(\frac{\alpha^{2}}{\sigma^{2}}\right)$ is the log signal-to-noise ratio (log-SNR), $\omega(\lambda)$ is the weighting function, and $p(\lambda)=-\frac{d\lambda}{dt}$ denotes the noise schedule.
+We adopt the EDM monotonic function as $\omega(\lambda)$, and update $p(\lambda)$ at each iteration to approximately satisfy:
+$p(\lambda)\propto\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\omega(\lambda)\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right].$ (16)
 […]
-Here, we introduce a hyperparameter $\xi>0$ to control the dependence of $\alpha(t)$ on $v(\alpha)$. As $\xi$ is increased, more timesteps are allocated to the region where $v(\alpha)$ is large.
-Similar to NCSN (score_matching1, ) inspired by simulated annealing (sa_original, ), our formulation draws inspiration from the adiabatic theorem (adiabatic_theorem, ) in quantum annealing (qa_original, ).
-Quantum annealing is an algorithm for solving combinatorial optimization problems, where the target cost function to be optimized is interpolated with an initial cost function with a trivial optimal solution.
-The initial state of the optimization is set to the optimal solution of the initial cost function.
-By tracing the optimal solutions of the instantaneous cost functions during the interpolation, the optimal solution of the target cost function is then obtained.
-In quantum annealing, accurately tracing the optimal solution of an instantaneous cost function is crucial, and the adiabatic theorem derives the traceability of the optimal solution at each step.
-Let $\alpha(t)$ be the weight of the target cost function and $1-\alpha(t)$ be the weight of the initial cost function.
-In this case, the inverse traceability can be expressed in the form of $\frac{d\alpha(t)}{dt}v(\alpha)$, which is the same form as Eq.
+Sampling Methods: To assess the generality across samplers, we employ two stochastic samplers and three deterministic samplers.
+The stochastic samplers include Stochastic DDIM ($\eta=1$) (Song et al., 2021a) and SDE-DPM-Solver++(2M) (Lu et al., 2022b).
+The deterministic samplers include DDIM (Song et al., 2021a), DPM-Solver++(2M) (Lu et al., 2022b), and pseudo numerical methods (PNDM) (Liu et al., 2022).
+At low NFE regimes, we also evaluate improved PNDM (iPNDM) (Zhang and Chen, 2023) and the unified predictor-corrector framework (UniPC) (Zhao et al., 2023).
+Evaluation Metrics: We use four evaluation metrics: FID (Heusel et al., 2017), spatial FID (sFID) (Nash et al., 2021), and improved precision and recall (Sajjadi et al., 2018; Kynkäänniemi et al., 2019).
+FID, sFID, and precision measure sample fidelity, while recall quantifies mode coverage.
+5.2 Results on Pixel-Space Diffusion Models
+We trained continuous-time diffusion models on LSUN Horse $\times256$ and LSUN Bedroom $\times256$ for unconditional image generation. Detailed experimental results are provided in Appendix F.1.
+LSUN Horse $\times256$: The evaluation results are summarized in Table 2.
+Rows shaded in gray represent results obtained under different training schedules.
+Although sample fidelity and mode coverage are generally considered to be in a trade-off relationship, CRS-$v_{x}$ achieved the highest scores across all evaluation metrics.
+This indicates that CRS-$v_{x}$ does not merely balance fidelity and diversity but leads to an overall improvement in sample quality.
+The unshaded rows represent the dependence on sampling schedules.
+CRS-$v_{x}+v_{\mathrm{FID}}$ outperformed all baselines across all metrics, indicating a fundamental improvement in generation quality.
+Notably, our model achieved an FID score of 2.03, surpassing the previous state-of-the-art score of 2.11 reported in Kumari et al.
 (2022).
-It has also been verified that an optimal functional form of $\alpha(t)$ is obtained by minimizing the maximum value of the inverse traceability grover_optimal , which ensures a consistent traceability throughout the interpolation process.
-In CRS, we regard the probability-distributional change as the inverse traceability and minimize the maximum distance for determining the noise schedule.
-4.3 Distance Metrics
-Any distance metric is applicable to CRS. We introduce three distance metrics.
-Distance metric based on FID: FID is a widely used distance metric to evaluate the sample fidelity in the field of image generation.
-The FID calculates the Fréchet distance fid_score under a normality assumption in the vision-relevant feature space embedded using the Inception-V3 model inception_v3 , which is trained on the ImageNet classification.
-When FID is used as the distance metric for the pixel-space diffusion models, CRS allocates more timesteps to the region where the sample fidelity significantly deteriorates in the diffusion process ($\alpha\simeq 1$).
-A weakness of FID-based distance metric is that it cannot accurately capture the probability-distributional change in the region with high-noise levels ($\alpha\simeq 0$), because the feature model is not trained on diffused images.
-It is important to note that, in the latent-space diffusion models, the probability-distributional change must be measured in the latent space embedded using the autoencoder (the approximation in Eq.
-(3) is applied to the diffusion process defined in the latent space).
-We trained ResNet50 resnet on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model for computing FID.
-We refer to $v(\alpha)$ computed using FID as $v_{\mathrm{FID}}(\alpha)$.
-Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain the noise schedule using Eq.
-(8).
-For the detailed computation procedure of $v_{\mathrm{FID}}(\alpha)$, refer to Appendix C.1.
-Distance metric based on data prediction: We introduce a distance metric that uses a trained diffusion model as follows:
-$v_{x}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{x}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].\$ (9)
-Here, $\bm{x}_{\alpha}$ and $\bm{x}_{\alpha^{\prime}}$ are sampled from $q(\bm{x}_{\alpha},\bm{x}_{\alpha}^{\prime})=q(\bm{x}_{\alpha}^{\prime}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})$, and $\bm{x}_{\theta}(\bm{x},\alpha)$ is the data prediction, which is related to the noise prediction $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ as follows: $\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$.
-This distance metric aims at allocating more computational resources to the region where the data prediction changes rapidly.
-As shown in Appendix B, we can interpret this distance metric as a simplified version of the Kullback-Leibler divergence upper bound (KLUB) introduced in AYS.
-This distance metric can be used not only for determining the sampling schedule after training a diffusion model, but also for adaptively optimizing the training schedule during the training process.
-Pseudocodes for optimizing training and sampling schedules using $v_{x}(\alpha)$ are shown in Appendix C.2.
-We can interpret that $v_{x}(\alpha)$ measures the probability-distributional change of diffused data.
-The data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ is given by the weighted sum of the data points $\bm{x}_{0}$ statistical_mechanics .
-The weight of each data point depends on the noise level $\alpha$.
-The weights of all data points are uniform when $\alpha=0$, while a small number of data points contribute to $\bm{x}_{\theta}(\bm{x},\alpha)$ when $\alpha\simeq 1$.
-The changes in these weights represent how the trajectories in the diffusion process gradually diverge as the noise level decreases.
-Therefore, we can interpret $v_{x}(\alpha)$ as measuring the probability-distributional change of diffused data.
-Unlike the FID-based distance metric, $v_{x}(\alpha)$ can capture distributional changes even in regions with high-noise levels.
-However, $v_{x}(\alpha)$, which is related to the KLUB, is not necessarily optimal for improving sample fidelity at sampling.
-Therefore, as discussed in the next subsection, combining multiple distance metrics in a complementary manner enhances the performance of CRS.
-Distance metric based on noise prediction: In the experiments presented in Sec.
-5, we trained the noise prediction model.
-In this case, it is also considered effective to use noise predictions to measure the probability-distributional change as follows:
-$v_{\varepsilon}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (10)
-Hereafter, we refer to CRS using $v_{\mathrm{FID}}(\alpha)$, $v_{x}(\alpha)$, and $v_{\varepsilon}(\alpha)$ for optimizing noise schedules as CRS-$v_{\mathrm{FID}}$, CRS-$v_{x}$, and CRS-$v_{\varepsilon}$, respectively.
-4.4 Combining Multiple Distance Metrics
-As explained in the previous subsection, each distance metric has its own strengths an weaknesses.
-Therefore, it is considered effective to complementarily combining multiple distance metrics for optimizing noise schedules.
-We introduce a strategy for combining them as follows:
-$v(\alpha)=\sum_{m=1}^{M}w_{m}\tilde{v}_{m}(\alpha),\ \ \mathrm{where}\ \ \tilde{v}_{m}(\alpha)=\frac{1}{C_{m}}v_{m}(\alpha)^{\xi_{m}}\ \ \mathrm{and}\ \ C_{m}=\int_{\alpha_{\min}}^{\alpha_{\max}}v_{m}(\alpha)^{\xi_{m}}.$ (11)
-Here, $M$ is the number of distance metrics to be combined, $v_{m}(\alpha)$ is computed for each distance metric, $w_{m}$ is the weight for $v_{m}(\alpha)$ satisfying $\sum_{m=1}^{M}w_{m}=1$, and $\xi_{m}$ controls the dependence on $v_{m}(\alpha)$.
-Note that the value range of $v_{m}(\alpha)$ can significantly vary depending on the distance metric.
-Therefore, normalization of $v_{m}(\alpha)$ is required before combining them, as shown in Eq.
-(11).
-Using this formulation, it is also possible to combine conventional noise schedules with CRS based noise schedules by converting $\alpha(t)$ to $v(\alpha)$.
-For example, when $\alpha(t)=\cos\left(\frac{\pi t}{2}\right)$, $v(\alpha)$ is obtained as follows:
-$v(\alpha)\propto-\left(\frac{d\alpha(t)}{dt}\right)^{-1}=\frac{2}{\pi}\sin^{-1}\left(\frac{\pi t}{2}\right)=\frac{2}{\pi\sqrt{1-\alpha^{2}}}.$ (12)
-This result suggests that every noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
-5 Experiments
-We experimentally demonstrated that CRS broadly improves the performance of diffusion models. Below, we provide a brief explanation of each component of the diffusion model.
-Dataset: We use six image datasets: LSUN (Church, Bedroom, Horse) dataset_lsun , ImageNet dataset_imagenet , FFHQ dataset_ffhq , and CIFAR10 dataset_cifar10 .
-LSUN and FFHQ are used for unconditional image generation, and ImageNet and CIFAR10 for class-conditional image generation.
-Noise Prediction Model: We use the U-Net architecture introduced in the ablated diffusion model (ADM) adm as the noise prediction model (see Table 8 in Appendix E for hyperparameter settings).
-Training Schedule: We use linear ddpm_original , shifted cosine shifted_cosine , adaptive noise schedule proposed in VDM++ vdmpp , and CRS.
-Loss Function: Except for the adaptive noise schedule of VDM++, we use the following simplified loss function introduced in DDPMs:
-$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},t\sim\mathcal{U}(0,1),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x},\alpha(t))-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (13)
-where $\mathcal{D}$ denotes all data points in the training dataset, $\alpha(t)$ is the training schedule, and $\bm{x}=\alpha(t)\bm{x}_{0}+\sqrt{1-\alpha(t)^{2}}\bm{\varepsilon}$.
-When using the adaptive noise schedule of VDM++, the following loss function is applied:
-$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\lambda\sim p(\lambda),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\frac{\omega(\lambda)}{p(\lambda)}\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (14)
-where $\lambda=\log\left(\frac{\alpha^{2}}{\sigma^{2}}\right)$ is the log signal-to-noise ratio (log-SNR), $\omega(\lambda)$ is the weight for each noise level, and $p(\lambda)=-\frac{d\lambda}{dt}$ is the noise schedule for training.
-We use the EDM monotonic function as $\omega(\lambda)$, and $p(\lambda)$ is updated at each iteration to approximately satisfy the following condition:
-$p(\lambda)\propto\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\omega(\lambda)\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right].$ (15)
-Sampling Schedule: We use linear, shifted cosine, EDM sampling edm and CRS.
-In pixel-space diffusion models, we additionally evaluate CRS-$v_{x}+v_{\mathrm{FID}}$, which combines $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ for optimizing noise schedules.
-The reason for selecting the combination of $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ is that, as shown in Sec.
-5.4, $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ have a complementary relationships.
-Recent studies edm ; vdmpp have shown that using different noise schedules for training and sampling can lead to improved performance.
-Therefore, in our experiments, we evaluate the performance of the diffusion models using arbitrary pairs of training and sampling schedules.
-Note that we can obtain different noise schedules for training and sampling using CRS, because it is capable of handling arbitrary distance metrics.
-Details of the conventional noise schedules used in our experiments can be found in Table 10 in Appendix E.
-Sampling Methods: Images are mainly generated using two stochastic samplers and three deterministic samplers.
-For the stochastic samplers, we use Stochastic DDIM ($\eta=1$) ddim_original and SDE-DPM-Solver++(2M) sampler_dpmpp .
-For the deterministic samplers, we use DDIM ddim_original , DPM-Solver++(2M) sampler_dpmpp , and pseudo numerical methods for diffusion models (PNDMs) sampler_pndm .
-At small NFEs, we additionally use improved PNDM (iPNDM) sampler_ipndm and a unified predictor-corrector framework (UniPC) sampler_unipc .
-Evaluation Metrics: We compute four metrics: FID fid_score , spatial FID (sFID) spatial_fid_score , and improved precision and recall recall_precision1 ; recall_precision2 .
-FID, sFID, and precision measure sample fidelity, while recall evaluates mode coverage.
-5.1 Results on Pixel-Space Diffusion Models
-We trained the continuous-time diffusion models on LSUH Horse 256 $\times$ 256 and Bedroom 256 $\times$ 256 for unconditional image generation.
-Detailed experimental results are listed in Appendix F.1.
-LSUN Horse 256 $\times$ 256: The evaluation results are summarized in Table 2.
-The rows shaded in gray represent the dependence on the training schedule.
-Although sample fidelity and mode coverage are generally considered to be in a trade-off relationship, CRS-$v_{x}$ achieved the highest performance for all metrics.
-This result indicates that CRS-$v_{x}$ is not merely balancing trade-offs between evaluation metrics but fundamentally improves the quality of the generated samples.
-The unshaded rows represent the dependence on sampling schedules.
-CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the highest performance for all metrics, indicating fundamental improvement in sample quality (see Appendix D for the hyperparameter setting of CRS-$v_{x}+v_{\mathrm{FID}}$).
-The FID score of 2.03 surpasses the current state-of-the-art FID score of 2.11 lsun_horse_sota .
-LSUN Bedroom 256 $\times$ 256: We confirmed the advantage of CRS in the low-NFE regime by comparing it with the leading results reported in GITS noise_gits .
-We trained the continuous-time diffusion model and used CRS-$v_{x}$ for optimizing the training schedule.
-The results are listed in Table 2.
-The best FID scores at $\mathrm{NFE}=5$ and $\mathrm{NFE}=10$ were achieved when using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$ for optimizing the training and sampling schedules, respectively.
-Table 1: Dependence on training and sampling schedules for LSUN Horse $256\times 256$ in pixel-space diffusion models at $\mathrm{NFE}=250$.
-SDE-DPM-Solver++(2M) was used as sampler.
-Bold values represent best training or sampling schedules for each metric.
+LSUN Bedroom $\times256$: We further validated the effectiveness of CRS in low-NFE regimes by comparing it with strong baselines reported in GITS (Chen et al., 2024).
+We trained continuous-time diffusion models with CRS-$v_{x}$ used for optimizing the training schedule.
+Table 2 lists the results.
+At $\mathrm{NFE}=5$ and $\mathrm{NFE}=10$, the best FID scores were achieved by using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$ for optimizing training and sampling schedules, respectively.
+As shown in Table 20 in Appendix F.1, we also tuned the $\rho$ parameter in the EDM sampling schedule. Even with the optimal $\rho$ setting, CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperformed EDM.
+Table 1: Performance of pixel-space diffusion models on LSUN Horse $\times256$ with $\mathrm{NFE}=250$, using SDE-DPM-Solver++(2M) as the sampler.
+Gray-shaded rows indicate variations in the training schedule, while unshaded rows correspond to different sampling schedules.
+Bold values highlight the best performance for each metric under varying training or sampling schedules.
 Training Sampling Metrics
@@
 CRS-$v_{x}+v_{\mathrm{FID}}$ 2.03 6.06 0.69 0.56
-Table 2: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion models at small NFEs.
-Bold values indicate best FID scores at each NFE.
-Results for rows other than "Ours" are taken from previously published study noise_gits .
+Table 2: FID scores for LSUN Bedroom $\times256$ in pixel-space diffusion models under low NFE conditions.
+Our results are compared with prior work from Chen et al.
+(2024).
+Bold entries indicate the best FID achieved at each NFE setting.
 Model Sampler Sampling schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$
-EDM UniPC sampler_unipc EDM 23.34 5.75
-AMED-Solver sampler_amed 5.38
-iPNDM sampler_ipndm GITS noise_gits 15.85 5.28
+EDM UniPC (Zhao et al., 2023) EDM 23.34 5.75
+AMED-Solver (Zhou et al., 2024) 5.38
+iPNDM (Zhang and Chen, 2023) GITS (Chen et al., 2024) 15.85 5.28
 Ours DPM-Solver++(2M) EDM 23.72 4.73
@@
 5.3 Results on Latent-Space Diffusion Models
-We trained latent-space diffusion models using the pre-trained autoencoder VQ-4 used in latent diffusion models (LDMs) ldm .
-We used LSUN Church and Bedroom for unconditional image generation, and ImageNet for class-conditional image generation.
-The image resolution was set to $256\times 256$, which was reduced to $64\times 64$ in the latent space.
-We present the evaluation results on LSUN Church $256\times 256$.
-Detailed results are listed in Appendix F.2.
-LSUN Church 256 $\times$ 256: The evaluation results are summarized in Table 4.
-The dependence on the training schedule is shown in the rows shaded in gray, and CRS-$v_{x}$ achieved the highest performance except for sFID.
-The unshaded rows represent the dependence on the sampling schedule, and CRS-$v_{x}$ further improves all evaluation metrics except for precision.
-These results indicate the superiority of CRS in latent-space diffusion models as well.
-From the results in Table 4, CRS-$v_{x}$ appears to perform the best at both training and sampling; however, depending on the sampler and NFE, CRS-$v_{\varepsilon}$ is preferable at sampling.
-5.3 Comparison with AYS Using EDM
-Because $v_{x}(\alpha)$ is related to the KLUB introduced in AYS, we additionally compare CRS with AYS on FFHQ 64 $\times$ 64 using the pretrained model provided in EDMs edm .
-It is important to note that EDM adopts a variance-exploding process, while we have so far focused on the variance-preserving process.
-However, CRS can also be applied to the variance-exploding process by converting between $\alpha$ and the noise level in EDM $\sigma_{\mathrm{EDM}}$ via the log SNR, as shown below:
-$\mathrm{logSNR}=\log\frac{\alpha^{2}}{1-\alpha^{2}}=\log\frac{1}{\sigma_{\mathrm{EDM}}^{2}}\Rightarrow\alpha=\frac{1}{\sqrt{1+\sigma_{\mathrm{EDM}}^{2}}}.$ (16)
-The results are listed in Table 4. CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperformed AYS at all NFEs. Detailed results on pretrained EDM are listed in Appendix F.3.
-Table 3: Dependence on training and sampling schedules for LSUN Church $256\times 256$ in latent-space diffusion models at $\mathrm{NFE}=30$.
-DPM-Solver++(2M) was used as sampler.
-Bold values represent best training or sampling schedules for each metric.
+We trained latent-space diffusion models using the pre-trained autoencoder from LDMs (Rombach et al., 2022).
+In our experiments, we adopted the VQ-regularized variant for all datasets, which is the configuration primarily used in the publicly available pretrained models on GitHub (https://github.com/CompVis/latent-diffusion).
+Although a downsampling factor of $f=8$ is also commonly used, we set $f=4$ (VQ-f4), as it is expected to yield better FID scores according to Table 10 in the LDMs paper.
+LSUN Church and LSUN Bedroom were used for unconditional image generation, and ImageNet was used for class-conditional image generation.
+All images were resized to $\times256$ and encoded into $\times64$ latent representations.
+Here, we present the evaluation results on LSUN Church $\times256$.
+Detailed results are provided in Appendix F.2.
+LSUN Church $\times256$: Table 4 summarizes the results.
+Rows shaded in gray show the effect of different training schedules, where CRS-$v_{x}$ achieved the best performance on all metrics except sFID.
+Unshaded rows show the dependence on the sampling schedule, and again, CRS-$v_{x}$ achieved the best performance on all metrics except precision.
+These results indicate that CRS is also effective in latent-space diffusion models.
+Although CRS-$v_{x}$ performs best in most settings, we observe that depending on the sampler and NFEs, CRS-$v_{\varepsilon}$ can be preferable for sampling.
+Table 3: Performance of latent-space diffusion models on LSUN Church $\times256$ with $\mathrm{NFE}=30$, using DPM-Solver++(2M) as the sampler.
+Gray-shaded rows indicate variations in the training schedule, while unshaded rows correspond to different sampling schedules.
+Bold values highlight the best performance for each metric under varying training or sampling schedules.
 Training Sampling Metrics
@@
 CRS-$v_{x}$ 3.59 9.51 0.59 0.60
-Table 4: FID scores evaluated on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM. FID scores on AYS is taken from their original paper noise_ays .
-Sampler Noise Schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
-DPM-Solver++(2M) AYS noise_ays 5.43 3.29 2.62
-CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
+Table 4: FID scores on CIFAR10 $\times32$ and FFHQ $\times64$ evaluated using the pretrained model provided by EDM (Karras et al., 2022).
+Bold numbers indicate the best sampling schedule for each combination of dataset, sampler, and NFE.
+For reference, the FID scores of AYS (Sabour et al., 2024) and LD3 (Tong et al., 2025) are taken directly from their original papers.
+Sampling CIFAR10 $\times32$ FFHQ $\times64$
+Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=10$ $\mathrm{NFE}=20$
+DPM-Solver++(2M) AYS 2.98 2.10 5.43 3.29
+GITS 4.03 2.32 5.51 3.81
+LD3 3.38 2.36 3.98 2.89
+CRS-$v_{x}+v_{\mathrm{FID}}$ 3.92 2.27 5.35 3.10
+UniPC(3M) GITS 3.50 1.99 4.52 2.71
+LD3 2.84 3.27
+CRS-$v_{x}+v_{\mathrm{FID}}$ 2.52 1.99 5.23 2.64
+iPNDM(3M) GITS 2.73 2.08 3.96 3.09
+LD3 2.38 3.25
+CRS-$v_{x}+v_{\mathrm{FID}}$ 2.87 2.11 4.07 2.75
+iPNDM(4M) GITS 2.49 2.02 3.62 3.00
+CRS-$v_{x}+v_{\mathrm{FID}}$ 2.46 2.02 3.45 2.53
+5.4 Comparison with Recent Sampling Schedules
+Using the pretrained models released with EDM (Karras et al., 2022), we compare the performance of CRS as a sampling schedule against AYS (Sabour et al., 2024), GITS (Chen et al., 2024), and learning to discretize denoising diffusion ODEs (LD3) (Tong et al., 2025).
+All evaluations in this subsection are conducted using the official repository provided in the GITS paper, ensuring full consistency with their experimental protocol.
+It is important to note that EDM adopts a variance-exploding formulation, whereas all of our previous experiments have been conducted under the variance-preserving setting.
+However, CRS can still be applied to variance-exploding processes by converting between $\alpha$ and the EDM noise level $\sigma_{\mathrm{EDM}}$ via the log-SNR, as shown below:
+$\log\mathrm{SNR}=\log\left(\frac{\alpha^{2}}{1-\alpha^{2}}\right)=\log\left(\frac{1}{\sigma_{\mathrm{EDM}}^{2}}\right)\quad\Rightarrow\quad\alpha=\frac{1}{\sqrt{1+\sigma_{\mathrm{EDM}}^{2}}}.$ (17)
+We also note that, whereas all earlier experiments in this paper used the noise-prediction parameterization, EDM adopts the data-prediction parameterization with preconditioning.
+The evaluation results are presented in Table 4.
+The optimal sampling schedule varies across datasets, samplers, and NFE configurations.
+On average, LD3 achieves the highest performance, while CRS consistently outperforms GITS and can therefore be regarded as the second-best sampling schedule overall.
+Although CRS underperforms LD3 as a sampling schedule, a key advantage of CRS is that it also provides a unified framework for optimizing the training schedule, which LD3 does not address.
+Further improving the performance of CRS will likely require identifying more effective discrepancy measures, and exploring such alternatives is an important direction for future work.
+For completeness and fairness, we note that CRS-$v_{x}+v_{\mathrm{FID}}$ requires both the computation of $v_{\mathrm{FID}}(\alpha)$ and hyperparameter tuning, and the time needed to optimize the sampling schedule is substantially larger than that of GITS or LD3.
+However, sampling-schedule optimization is performed only once after training the diffusion model.
+Therefore, as long as the tuning cost remains small relative to the total training time, we believe that performance under the best hyperparameter configuration is the most meaningful criterion for comparison.
 5.5 Comparison of Noise Schedules
-The CRS-based noise schedules in the pixel-space diffusion model are shown in Fig.
-3.
-The hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$ was set to $w_{x}=w_{\mathrm{FID}}=0.5$ and $\xi_{x}=\xi_{\mathrm{FID}}=1.0$.
-Since $v_{\mathrm{FID}}(\alpha)$ takes large values in regions where sample fidelity degrades in the diffusion process, CRS-$v_{\mathrm{FID}}$ allocates many timesteps to the region where $\alpha\simeq 1$.
-In contrast, CRS-$v_{x}$ allocates more timesteps to regions with small $\alpha$.
-This is due to the fact that the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ can be expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$ using the noise prediction, where $\alpha$ appears in the denominator.
-Thus, CRS-$v_{\mathrm{FID}}$ and CRS-$v_{x}$ exhibits complementary characteristics, and combining them yields an effective sampling schedule.
-Since the noise schedule based on CRS-$v_{\varepsilon}$ can be approximately reproduced by tuning the hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$, we regard $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ as the fundamental compenents.
-Figure 3 shows the CRS-based noise schedules in the latent-space diffusion model.
-The appropriate noise schedules differ significantly between the pixel-space and lantet-space diffusion models.
-The most notable difference is that, in the latent-space diffusion models, fewer timesteps are allocated to the region where $\alpha\simeq 1$.
-This can be attributed to the fact that, in latent-space diffusion models, the perceptual expression is largely generated by the autoencoder, and the autoencoder can absorb minor deviations in the generated latent variables.
-In contrast, the pixel-space diffusion models directly generate images, so careful control of the noise schedule is crucial at $\alpha\simeq 1$.
-We did not evaluate any combination of multiple distance metrics in latent-space diffusion models, because all the CRS-based noise schedules have a similar functional form.
-Figure 2: Noise schedules on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model.
-Figure 3: Noise schedules on LSUN Church 256 $\times$ 256 in latent-space diffusion model.
+Figure 3 shows examples of CRS-based noise schedules in pixel-space diffusion models, computed using different discrepancy measures.
+For CRS-$v_{x}+v_{\mathrm{FID}}$, the hyperparameters were set to $w_{x}=w_{\mathrm{FID}}=0.5$ and $\xi_{x}=\xi_{\mathrm{FID}}=1.0$.
+Since $v_{\mathrm{FID}}(\alpha)$ tends to be large in regions where sample fidelity degrades (i.e., near $\alpha\simeq 1$), CRS-$v_{\mathrm{FID}}$ allocates more timesteps to these regions.
+In contrast, CRS-$v_{x}$ allocates more timesteps to regions with smaller $\alpha$.
+This is due to the formulation of the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, where $\alpha$ appears in the denominator, making the data prediction more sensitive to small changes in $\alpha$ when $\alpha$ is small.
+Thus, CRS-$v_{\mathrm{FID}}$ and CRS-$v_{x}$ exhibit complementary behaviors, and combining both yields an effective and balanced sampling schedule.
+Furthermore, we observed that the noise schedule derived from CRS-$v_{\varepsilon}$ can be approximately reproduced by appropriately tuning the hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$.
+Therefore, in practice, $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ serve as the two principal components for constructing robust sampling schedules.
+Figure 3 illustrates CRS-based noise schedules for latent-space diffusion models.
+The optimized schedules differ notably between pixel-space and latent-space models.
+In latent-space diffusion, fewer timesteps are allocated to the region where $\alpha\simeq 1$.
+This difference can be attributed to the nature of latent-space diffusion models: perceptual expression is largely reconstructed by the autoencoder, and the autoencoder is robust to minor deviations in the generated latent variables.
+Consequently, precise control near $\alpha\simeq 1$ is less critical.
+In contrast, pixel-space models directly generate the final image, so the schedule must allocate sufficient timesteps near $\alpha\simeq 1$ to preserve high-fidelity details.
+Because all discrepancy measures result in similar noise schedules for latent-space models, we did not evaluate combinations of multiple discrepancy measures.
+Figure 2: Noise schedules computed using different discrepancy measures on LSUN Horse $\times256$ in the pixel-space diffusion model.
+Figure 3: Noise schedules computed using different discrepancy measures on LSUN Church $\times256$ in the latent-space diffusion model.
+5.6 Computational Overhead of CRS
+CRS requires computing $v(\alpha)$. Below, we report the computational overhead in scenarios where the cost is expected to be most significant.
+Cost of computing $v_{\mathrm{FID}}(\alpha)$: Although $v_{\mathrm{FID}}(\alpha)$ needs to be computed only once before training the diffusion model, its computation requires processing the entire training set to compute feature statistics, and therefore the cost scales linearly with the number of images.
+Table 6 reports both the cost relative to a single training epoch and the cost relative to the total training time.
+Because the computation of $v_{\mathrm{FID}}(\alpha)$ and the duration of one training epoch both scale proportionally with dataset size, the cost relative to one training epoch remains nearly constant across datasets for both pixel-space and latent-space diffusion models.
+The relative cost appears larger for latent-space diffusion models simply because one latent-space epoch is much faster than one pixel-space epoch.
+When compared to the total training time, the overhead of computing $v_{\mathrm{FID}}(\alpha)$ remains below 15% even in the worst case.
+Overhead of computing $v_{x}(\alpha)$ during training: Table 6 compares the training time when using the linear versus CRS-$v_{x}$ as the training schedule.
+Because CRS-$v_{x}$ requires evaluating $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ during training, the overall training time increases by approximately 20–30%.
+Nevertheless, CRS-$v_{x}$ typically accelerates the convergence of FID scores during training (see Figs.
+13 and 13 in Appendix F for details), which can largely compensate for the per-epoch overhead.
+Table 5: Wall-clock computational cost of computing $v_{\mathrm{FID}}(\alpha)$.
+We report both the cost relative to one training epoch and the cost relative to the total training time for each diffusion setting.
+Computational cost
+Diffusion type Dataset Relative to one training epoch Relative to total training time
+Pixel space LSUN Horse $\sim$5 epochs $\sim$10%
+LSUN Bedroom $\sim$5 epochs $\sim$10%
+Latent space LSUN Church $\sim$14 epochs $\sim$1.4%
+LSUN Bedroom $\sim$14 epochs $\sim$14%
+ImageNet $\sim$14 epochs $\sim$2.8%
+Table 6: Comparison of wall-clock training time when using linear and CRS-$v_{x}$ as the training schedule.
+CRS-$v_{x}$ requires evaluating $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ during training, resulting in a 20–-30% increase in overall training time.
+Diffusion type Dataset Linear CRS-$v_{x}$ Increase
+Pixel space LSUN Horse 170 hours 215 hours +26.5%
+Latent space LSUN Church 85 hours 105 hours +23.5%
+LSUN Bedroom 190 hours 245 hours +28.9%
+ImageNet 430 hours 525 hours +22.1%
+5.7 A Practical Alternative to FID-Based Scheduling
+In pixel-space diffusion models, CRS-$v_{x}+v_{\mathrm{FID}}$ consistently achieves the best performance as a sampling schedule.
+However, for large-scale training datasets, computing $v_{\mathrm{FID}}(\alpha)$ becomes prohibitively expensive.
+As shown in the subsection 5.5, CRS-$v_{\mathrm{FID}}$ allocates a large number of timesteps to regions near $\alpha\simeq 1$, resulting in a schedule shape similar to that of the cosine schedule.
+Motivated by this observation, we evaluate a lightweight alternative: CRS-$v_{x}+v_{\mathrm{cos}}$, which combines CRS-$v_{x}$ with the cosine schedule (see Eq.
+13 for $v(\alpha)$ of the cosine schedule).
+Table 7 compares the FID scores of CRS-$v_{x}+v_{\mathrm{FID}}$ and CRS-$v_{x}+v_{\mathrm{cos}}$ on LSUN Bedroom $\times256$.
+In all settings, CRS-$v_{x}+v_{\mathrm{cos}}$ achieves FID scores comparable to CRS-$v_{x}+v_{\mathrm{FID}}$, demonstrating that CRS-$v_{x}+v_{\mathrm{cos}}$ is a practical and effective alternative when $v_{\mathrm{FID}}(\alpha)$ is costly to compute.
+Table 7: Evaluation of CRS-$v_{x}+v_{\mathrm{cos}}$ as a practical alternative to CRS-$v_{x}+v_{\mathrm{FID}}$ for sampling in pixel-space diffusion models on LSUN Bedroom $\times256$.
+Despite not requiring the computation of $v_{\mathrm{FID}}(\alpha)$, CRS-$v_{x}+v_{\mathrm{cos}}$ achieves similar FID scores.
+Sampler Sampling Schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$
+DPM-Solver++(2M) CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 4.88
+CRS-$v_{x}+v_{\mathrm{cos}}$ 14.10 4.46
+UniPC CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 3.30
+CRS-$v_{x}+v_{\mathrm{cos}}$ 20.77 3.30
 6 Limitations
-This paper focuses on developing CRS and validating its effectiveness under fundamental experimental settings.
-However, to further strengthen the generality of CRS, additional investigations are needed from the following perspectives: 1) generality across various generation conditions, 2) scalability of the distance metrics with respect to dataset size and data dimensionality, 3) effectiveness in domains beyond image generation.
-While $v_{\mathrm{FID}}(\alpha)$ proved effective as the sampling schedule in pixel-space diffusion models, it is a distance metric specialized for image generation.
-Therefore, when applying CRS to other domains, it may be necessary to introduce new distance metrics that are better suited to the characteristics of the target domain.
+This paper focuses on developing CRS and empirically validating its effectiveness under fundamental experimental settings.
+While the results are promising, there are several directions for future work to further enhance the generality, applicability, and performance of CRS.
+Below, we summarize six key limitations and avenues for future work.
+(1) Theoretical analysis.
+The goal of this work is to provide a theoretically motivated and empirically effective framework capable of incorporating a wide range of discrepancy measures.
+However, we do not provide a theoretical guarantee that minimizing the maximum discrepancy necessarily leads to improved sample quality.
+Establishing a formal connection between distributional change and sample quality remains an important direction for future research.
+(2) Evaluation on more complex tasks.
+Our experiments focus on fundamental settings—unconditional and class-conditional image generation.
+We did not evaluate CRS in more complex practical scenarios such as text-conditional generation or classifier-free guidance (CFG).
+CFG is known to improve FID at the cost of reduced recall, and an adaptive noise schedule may help mitigate this trade-off.
+Evaluating CRS together with CFG or other conditioning mechanisms could further reveal the benefits of schedule optimization.
+(3) Applicability beyond image generation.
+Although our experiments are limited to image generation, CRS is derived from general diffusion-model principles and is therefore expected to generalize to other domains.
+In this work, $v_{\mathrm{FID}}(\alpha)$ proved effective as a sampling discrepancy measure for pixel-space diffusion models; however, it is specific to image generation.
+Applying CRS to non-image modalities will require identifying discrepancy measures that reflect the characteristics of each target domain.
+(4) Improving discrepancy measures.
+As a sampling schedule optimizer, CRS currently performs slightly worse than LD3, suggesting that further improvements may be achieved by designing better discrepancy measures.
+A promising direction is to build on advances in schedule-optimization theory such as AYS.
+As shown in Appendix B, CRS-$v_{x}$ can be interpreted as employing a simplified variant of the KLUB introduced in AYS.
+Thus, future theoretical progress in schedule optimization may naturally lead to new discrepancy measures compatible with CRS.
+With respect to image-specific metrics, although our experiments primarily relied on FID, many feature-based perceptual metrics have been proposed for image generation.
+For example, LPIPS (Zhang et al., 2018) is a strong candidate for capturing visually meaningful differences.
+Since FID and LPIPS depend on feature extractors, a key challenge is designing robust feature representations for noised data, which are essential for accurately assessing distributional change.
+Self-supervised learning on noised samples is a promising direction, as it requires no labels and can be made noise-level aware (e.g., by predicting or conditioning on $\alpha$).
+Such learned embeddings may lead to more expressive and domain-adaptive discrepancy measures.
+(5) Generality across prediction parameterizations and architectures.
+Our experiments primarily evaluate CRS under the noise-prediction parameterization.
+However, the data- and velocity-prediction formulations are also widely used in recent diffusion models, and the generality of CRS under these parameterizations has not been thoroughly assessed.
+Furthermore, modern diffusion architectures such as Diffusion Transformers (Peebles and Xie, 2023) were not evaluated.
+Because CRS is grounded in the dynamics of distributional evolution rather than in model architecture, we expect architectural dependence to be limited, but empirical validation remains an important direction for future work.
+(6) Scalability to dataset size and image resolution.
+Although CRS-$v_{x}+v_{\mathrm{FID}}$ is effective for optimizing sampling schedules in pixel-space diffusion models, it requires both the computation of $v_{\mathrm{FID}}(\alpha)$ and hyperparameter tuning.
+In particular, the computation of $v_{\mathrm{FID}}(\alpha)$ scales linearly with the number of training images, making this variant impractical for large-scale datasets.
+Although Section 5.7 demonstrates that CRS-$v_{x}+v_{\mathrm{cos}}$ can serve as a practical alternative, fundamentally improving scalability will require developing new discrepancy measures that are both computationally efficient and robust to noise-level variation.
+Furthermore, all of our experiments were conducted at a resolution of $\times256$.
+Validating whether CRS remains effective for higher-resolution image generation, and analyzing how the optimal noise schedule depends on image size, are compelling directions for future research.
 7 Conclusion
-We proposed a general approach called CRS to optimize noise schedules for training and sampling in diffusion models.
-Through extensive experiments, we confirmed that CRS broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and NFEs ranging from 5 to 250.
-CRS is fundamentally different from conventional approaches noise_ays ; noise_gits in that it can optimize the training schedule as well and use any distance metric for measuring probability-distributional change, allowing the introduction of domain specific distance metric and the use of different noise schedules for training and sampling.
-Although this study has the limitations discussed in Sec.
-6, we believe that it successfully provides solid baselines.
-We hope that future research will further improve the generality of CRS, leading to its widespread adoption as a general-purpose approach.
+We proposed a general framework called CRS for optimizing noise schedules in diffusion models.
+CRS enforces a constant rate of distributional change throughout the diffusion process and provides a unified perspective for designing both training and sampling schedules.
+Our framework supports arbitrary user-specified discrepancy measures, and we introduced three practical options that capture different aspects of the distributional dynamics.
+Through extensive experiments, we demonstrated that CRS consistently improves the performance of both pixel-space and latent-space diffusion models across various datasets, samplers, and NFEs ranging from 5 to 250.
+In particular, for sampling in pixel-space diffusion models, CRS-$v_{x}+v_{\mathrm{FID}}$ with tuned hyperparameters achieved the best results.
+In all other settings, CRS-$v_{x}$ with $\xi=1$ performed consistently well and is recommended as a default configuration due to its simplicity and robustness.
+Although this study has limitations, as discussed in Sec.
+6—including the lack of a complete theoretical analysis and the need for further validation in other domains—we believe that CRS provides a solid and extensible foundation for future research on schedule optimization in diffusion models.
+We hope that this work encourages broader exploration of discrepancy-based scheduling, and that future research will further generalize CRS and extend its applicability beyond image generation.
+Acknowledgments
+We thank the anonymous reviewers for their constructive and insightful comments, which helped improve the quality of this paper.
+We are also grateful to the team responsible for operating and maintaining the AI Cloud Platform, which provides the internal GPU infrastructure used in this work.
+Without their support, this research would not have been possible.
 Appendix A Probability Distribution of Diffused Data
-We derive the probability distribution of diffused data used in the toy example (Fig. 1).
-The probability distribution of diffused data is given by
-[…]
+We derive the probability distribution of the diffused data used in the toy example (Fig. 1).
+The distribution of the diffused data is given by:
+$q(\bm{x}_{\alpha})$ $=\int q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})d\bm{x}_{0},$ (18)
 $q(\bm{x}_{\alpha}|\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0},\sigma^{2}\bm{I}),$ (19)
-where $q(\bm{x}_{0})$ is the probability distribution of a target dataset. We then approximate $q(\bm{x}_{0})$ with the empirical distribution as
+where $q(\bm{x}_{0})$ is the probability distribution of the target dataset.
+We approximate $q(\bm{x}_{0})$ by the empirical distribution:
 $q(\bm{x}_{0})=\frac{1}{N}\sum_{n=1}^{N}\delta(\bm{x}_{0}-\bm{x}_{0}^{n}),$ (20)
-where $N$ is the number of samples in the target dataset, $\bm{x}_{0}^{n}$ is the $n$-th sample, and $\delta$ denotes Dirac’s delta function. By substituting Eq. (19) into Eq. (17), we obtain
+where $N$ is the number of data samples, $\bm{x}_{0}^{n}$ is the $n$-th sample, and $\delta$ denotes the Dirac delta function. Substituting Eq. (20) into Eq. (18) yields:
 $q(\bm{x}_{\alpha})=\frac{1}{N}\sum_{n=1}^{N}\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0}^{n},\sigma^{2}\bm{I}).$ (21)
-We used the following formula of Dirac’s delta function:
-[…]
+This derivation uses the following formula of Dirac’s delta function:
+$\int f(\bm{x})\delta(\bm{x}-\bm{y})d\bm{x}=f(\bm{y}).$ (22)
 The resulting probability distribution is a Gaussian mixture distribution. We used Eq. (21) in the toy example.
-Appendix B Derivation of Distance Metric Based on Data Prediction from KLUB
-We show that $v_{x}(\alpha)$ can be derived as a simplified version of KLUB introduced in AYS.
+Appendix B Derivation of Discrepancy Measure Based on Data Prediction from KLUB
+We show that $v_{x}(\alpha)$ can be derived as a simplified variant of the KLUB introduced in AYS.
 B.1 KLUB in Variance-Preserving Process
-The forward SDE of the continuous-time diffusion models [19] is defined as
+The forward SDE in continuous-time diffusion models (Song et al., 2021b) is defined as:
 $d\bm{x}=f(t)\bm{x}dt+g(t)d\bm{\omega},$ (23)
-where $f(t)\bm{x}$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $\bm{\omega}$ is the standard Wiener process. The associated reverse SDE is given by
+where $f(t)\bm{x}$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $\bm{\omega}$ is the standard Wiener process. The corresponding reverse SDE is given by:
 $d\bm{x}=\left[f(t)\bm{x}-g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\right]dt+g(t)d\bm{\omega}.$ (24)
-For approximately computing the score $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, we train the following noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$:
+To approximate the score $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, we train a noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$:
 $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)\simeq-\sigma\nabla_{\bm{x}}\log p_{t}(\bm{x}).$ (25)
-The noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ has the following relationship with the data prediction model $\bm{x}_{\theta}(\bm{x},\alpha)$:
+The data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ is related to the noise prediction via:
 $\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha).$ (26)
@@
 $\bm{f}(\bm{x},t)$ $=\left(f(t)+\frac{1}{\sigma(t)^{2}}g(t)^{2}\right)\bm{x}-\frac{\alpha(t)}{\sigma(t)^{2}}g(t)^{2}\bm{x}_{\theta}(\bm{x},\alpha(t)).$ (28)
-Stochastic DDIM is a stochastic sampler designed to numerically solve the reverse SDE described above.
-In Stochastic DDIM, the solution for the first term of $\bm{f}(\bm{x},t)$ (linear term) is exactly computed, and discretization errors arise only from the second term of $\bm{f}(\bm{x},t)$ [22, 23].
-According to a previous study [16], in the sub-interval $[t_{i-1},t_{i}]$, Stochastic DDIM solves the following discretized SDE:
+Stochastic DDIM numerically solves this reverse SDE.
+It exactly integrates the linear term and approximates the remaining term, introducing discretization error (Lu et al., 2022a; b).
+According to Sabour et al.
+(2024), in the interval $t_{i-1},t_{i}]$, Stochastic DDIM solves the discretized SDE:
 $d\bm{x}$ $=\bm{f}^{\prime}(\bm{x},t)dt+g(t)d\bm{\omega},$ (29)
 $\bm{f}^{\prime}(\bm{x},t)$ $=\left(f(t)+\frac{1}{\sigma(t)^{2}}g(t)^{2}\right)\bm{x}-\frac{\alpha(t)}{\sigma(t)^{2}}g(t)^{2}\bm{x}_{\theta}(\bm{x}_{t_{i}},\alpha(t_{i})).$ (30)
-The KLUB between probability distributions of $\bm{x}_{t_{i-1}}$ obtained by exactly solving the reverse SDE and approximately solving the reverse SDE using Stochastic DDIM is introduced as
-[…]
-[…]
-The diffusion coefficient $g(t)$ in the variance-preserving process is given by
+Furthermore, the KLUB between probability distributions of $\bm{x}_{t_{i-1}}$ obtained by solving the true reverse SDE and the descretized one is introduced as
+$\mathrm{KLUB}(t_{i-1},t_{t})$ $=\frac{1}{2}\int_{t_{i-1}}^{t_{i}}\mathbb{E}_{{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t_{i}}\sim q(\bm{x}_{t_{i}}|\bm{x}_{t})\end{subarray}}}\left[\frac{\left\|\bm{f}(\bm{x}_{t},t)-\bm{f}^{\prime}(\bm{x}_{t},t)\right\|_{2}^{2}}{g(t)^{2}}\right]dt,$ (31)
+$=\frac{1}{2}\int_{t_{i-1}}^{t_{i}}\mathbb{E}_{{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t_{i}}\sim q(\bm{x}_{t_{i}}|\bm{x}_{t})\end{subarray}}}\left[\frac{\alpha(t)^{2}}{\sigma(t)^{4}}g(t)^{2}\left\|\bm{x}_{\theta}(\bm{x}_{t},\alpha(t))-\bm{x}_{\theta}(\bm{x}_{t_{i}},\alpha(t_{i}))\right\|_{2}^{2}\right]dt.$ (32)
+In the variance-preserving setting, the diffusion coefficient satisfies:
 $\frac{1}{2}g(t)^{2}=-\frac{\dot{\alpha}}{\alpha}.$ (33)
-By substituting Eq. (32) into Eq. (31), we obtain
-[…]
-B.2 KLUB as Distance Metric of CRS
-We transform the KLUB into a form that can be applied to CRS.
-Because the KLUB is used to evaluate the discretization error when solving the reverse SDE, it is reasonable to regard the KLUB as the traceability measure of the diffusion process.
-First, we approximate the integral in the KLUB using Simpson’s rule as follows:
+Substituting Eq. (33) into Eq. (32) gives:
+$\mathrm{KLUB}(t_{i-1},t_{i})=-\int_{t_{i-1}}^{t_{i}}dt\dot{\alpha}(t)\frac{\alpha(t)}{\sigma(t)^{4}}\mathbb{E}_{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t_{i}}\sim q(\bm{x}_{t_{i}}|\bm{x}_{t})\end{subarray}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{t},\alpha(t))-\bm{x}_{\theta}(\bm{x}_{t_{i}},\alpha(t_{i}))\right\|_{2}^{2}\right].$ (34)
+B.2 KLUB as Discrepancy Measure of CRS
+To apply KLUB within the CRS framework, we reinterpret it as a discrepancy measure over adjacent steps in the diffusion process.
+Because the KLUB is introduced to evaluate the discretization error when solving the reverse SDE, it is reasonable to regard the KLUB as the traceability measure of the diffusion process.
+We approximate the integral in the KLUB using Simpson’s rule as follows:
 $\mathrm{KLUB}(t,t+\Delta t)$ $=\frac{\dot{\alpha(t)}\alpha(t)}{2\sigma(t)^{4}}\overline{\mathcal{D}_{x}^{2}}(t,t+\Delta t)\Delta t,$ (35)
 $\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ $=\mathbb{E}_{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t^{\prime}}\sim q(\bm{x}_{t^{\prime}}|\bm{x}_{t})\end{subarray}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{t},\alpha(t))-\bm{x}(\bm{x}_{t^{\prime}},\alpha(t^{\prime}))\right\|_{2}^{2}\right].$ (36)
-Next, to treat $\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ as a function of $\alpha$ rather than $t$, we introduce $\overline{D_{x}^{2}}(\alpha(t),\alpha(t^{\prime}))=\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$:
-$\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (36)
+Then, to treat $\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ as a function of $\alpha$ rather than $t$, we introduce $\overline{D_{x}^{2}}(\alpha(t),\alpha(t^{\prime}))=\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$:
+$\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right],$ (37)
 Here, $q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ is given by
-$q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ $=q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0}),$ (37)
+$q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ $=\int q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})d\bm{x}_{0},$ (38)
 $q(\bm{x}_{\alpha}|\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0},\sigma^{2}\bm{I}),$ (39)
@@
 We use $\overline{D_{x}^{2}}(\alpha,\alpha)=0$ in the second equality.
-Finally, by adopting $\sqrt{\mathrm{KLUB}(t,t^{\prime})}$ as the distance metric $D(t,t^{\prime})$ in CRS, we obtain the following equation:
+Finally, by adopting $\sqrt{\mathrm{KLUB}(t,t^{\prime})}$ as the discrepancy measure $D(t,t^{\prime})$ in CRS, we obtain the following equation:
 $D(t,t+\Delta t)=\sqrt{\mathrm{KLUB}(t,t+\Delta t)}=-\frac{d\alpha}{dt}\cdot\underbrace{\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}\sqrt{\left.\frac{\partial\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}}_{v_{\mathrm{KLUB}}(\alpha)}\Delta t.$ (44)
-which is the same form as Eq.
+which corresponds to the general form of CRS in Eq.
 (8).
-We can regard $v_{x}(\alpha)$ as a simplified version of $v_{\mathrm{KLUB}}(\alpha)$ in the sense that the weighting function of $v_{\mathrm{KLUB}}(\alpha)$, namely $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$, is discarded in $v_{x}(\alpha)$.
-It is possible to optimize the noise schedule using $v_{\mathrm{KLUB}}(\alpha)$.
-However, when $v_{\mathrm{KLUB}}(\alpha)$ is directly used, the noise schedule is heavily impacted by the weighting function rather than $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$.
-The weighting function becomes extremely large around $\alpha\simeq 1(\sigma\simeq 0)$, which overwhelms the impact of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$.
-Recall that $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ can be interpreted as measuring probability-distributional changes of diffused data, as described in Sec.
-4.3.
-Therefore, the simplified version we introduce is more aligned with our objective of optimizing the noise schedule on the basis of distributional changes.
+We can interpret $v_{x}(\alpha)$ as a simplified variant of $v_{\mathrm{KLUB}}(\alpha)$, in which the multiplicative weighting term $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$ is omitted.
+Although it is possible to optimize the noise schedule using $v_{\mathrm{KLUB}}(\alpha)$ directly, this can lead to undesirable behavior, as the weighting term $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$ grows very large when $\alpha\simeq 1$ (i.e., $\sigma\simeq 0$), which can overwhelm the influence of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$.
+Since $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ can be interpreted as capturing the distributional change between diffused data distributions (as discussed in Section 4.3), the simplified variant $v_{x}(\alpha)$ is more aligned with the objective of CRS.
 Appendix C Implementation of CRS
-We describe how to compute $v(\alpha)$ using each distance metric.
-C.1 Distance Metric Based on FID
-The forward process is simulated using the training dataset for computing $v_{\mathrm{FID}}(\alpha)$, as shown in Algorithm 1.
-One of the inputs $\{\alpha_{t}\}_{t=0,...,T}$ needs to be discretized to approximately represent the functional form of $v(\alpha)$ through linear interpolation.
-In our experiments, we set
+We describe how to compute $v(\alpha)$ using each discrepancy measure.
+C.1 Discrepancy Measure Based on FID
+Algorithm 1 presents the pseudocode for computing $v_{\mathrm{FID}}(\alpha)$, and hyperparameters are summarized in Tabel 8.
+Once the forward process is defined, $v(\alpha)$ can be computed by simulating the forward diffusion using the training dataset.
+Since this procedure is independent of model parameters, it only needs to be executed once prior to training the diffusion model.
+It is important to note that, in latent-space diffusion models, the probability-distributional change must be measured in the latent space embedded by the autoencoder, as the Gaussian approximation in Eq.
+(4) applies to the diffusion process defined in the latent space.
+Therefore, for latent-space models, we trained a ResNet-50 (He et al., 2016) classifier on ImageNet in the autoencoder’s latent space and use it as the feature extractor $\bm{\phi}$.
+To compute $v(\alpha)$, we discretize the range of $\alpha$ using a one-dimensional grid $\{\alpha_{t}\}_{t=0,\dots,T}$ and approximate the continuous function via linear interpolation.
+In our experiments, we define the discretized grid of $\alpha$ values as:
 $\alpha_{t}=1-\left(\frac{t}{T}\right)^{p},$ (45)
-where $T=1000$, and set $p$ to $2$ and $1$ for the diffusion processes in pixel space and latent space, respectively.
-For the feature model $\bm{\phi}$, we used the Inception-V3 model trained on ImageNet classification for the pixel-space diffusion process.
-For the latent-space diffusion process, however, we need a feature model trained in the latent space embedded using the autoencoder.
-We trained ResNet50 on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model $\bm{\phi}$.
-Note that a trained diffusion model is not necessary for computing $v_{\mathrm{FID}}(\alpha)$.
-Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain noise schedules by solving Eq.
-(8).
-The dependence of CRS-$v_{\mathrm{FID}}$ on hyperparameter $\xi$ is illustrated in Fig.
-7.
-CRS-$v_{\mathrm{FID}}$ is computed on LSUN Horse 256 $\times$ 256 for the pixel-space diffusion process.
-As $\xi$ is increased, more timesteps are allocated to the region with low-noise level.
-The computation time of $v_{\mathrm{FID}}(\alpha)$ is proportional to the number of images in the training dataset.
-For the pixel-space diffusion process, it takes about 35 hours for LSUN Horse 256 $\times$ 256 with 2M images using eight A100 GPUs.
-For the latent-space diffusion process, it takes about 12 hours for ImageNet 256 $\times$ 256 with 1.28M images using eight A100 GPUs.
-The computation of $v_{\mathrm{FID}}(\alpha)$ needs to be executed only once before training the diffusion model, and the computation time is sufficiently small compared with the training time of the diffusion model.
+where $T=1000$.
+For pixel-space diffusion models, where $v(\alpha)$ exhibits sharp variation near $\alpha\simeq 1$, we set $p=2$ to allocate more evaluation points in that region.
+For latent-space models, where $v(\alpha)$ is more uniformly distributed, we use $p=1$.
+The computational cost of $v(\alpha)$ is proportional to the number of training images.
+For the pixel-space diffusion process on LSUN Horse $\times256$ with 2M images, the computation takes approximately 35 hours using eight A100 GPUs.
+For the latent-space process on ImageNet $\times256$ with 1.28M images, it takes about 12 hours using eight A100 GPUs.
+To address the scalability issue for large-scale datasets, we introduce a practical alternative in Sec.
+5.7.
 Algorithm 1 $v(\alpha)$ on basis of FID
-0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, number of timesteps for simulating the forward process $T$, discretized noise levels $\{\alpha_{t}\}_{t=0,...,T}$, feature model $\bm{\phi}$
+1:training dataset $\{\bm{x}^{(n)}\}_{n=0,\dots,N-1}$, timesteps $T$, noise levels $\{\alpha_{t}\}_{t=0,\dots,T}$, feature model $\bm{\phi}$
 […]
-1:  # Compute statistics of feature vectors
-2:  for $n=0$ to $N-1$ do
-3:     $\bm{x}^{(n)}_{0}\leftarrow\bm{x}^{(n)}$
-4:     if diffusion process in latent space then
-5:        $\bm{x}^{(n)}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}^{(n)}_{0})$
-6:     end if
-7:
-8:     # Simulate the forward process
-9:     for $t=1$ to $T$ do
-10:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}}$
-11:        $\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
-12:        $x^{(n)}_{t}\leftarrow\beta_{t}\bm{x}^{(n)}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
-13:     end for
-14:  end for
-15:  $\bm{\mu}_{t}\leftarrow mean\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
-16:  $\bm{\Sigma}_{t}\leftarrow cov\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
+3:# Compute feature vectors of diffused data
+4:for n=0n=0toN1N-1 do
+5:  $\bm{x}^{(n)}_{0}\leftarrow\bm{x}^{(n)}$
+6:  if is_latent_space then
+7:   $\bm{x}^{(n)}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}^{(n)}_{0})$
+8:  end if
+9:
+10:  # Simulate the forward process
+11:  for t=1t=1toTT do
+12:   $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}}$
+13:   $\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
+14:   $\bm{x}^{(n)}_{t}\leftarrow\beta_{t}\bm{x}^{(n)}_{t-1}+\delta_{t}\,\mathcal{N}(\bm{0},\bm{I})$
+15:  end for
+16:end for
 17:
-18:  # Compute $v(\alpha)$ using $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$
-19:  for $t=0$ to $T-1$ do
-20:     $v_{t}\leftarrow\frac{\mathrm{FID}(\bm{\mu}_{t},\bm{\Sigma}_{t},\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})}{\alpha_{t}-\alpha_{t+1}}$
-21:  end for
-22:  $v_{T}\leftarrow v_{T-1}$
-Figure 4: Dependence of CRS-$v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$.
-Figure 5: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $w_{x}$ and $w_{\mathrm{FID}}$ ($\xi_{x}=\xi_{\mathrm{FID}}=1)$.
-Figure 6: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
-Figure 7: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{x}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
-C.2 Distance Metric Based on Data Prediction
+18:# Compute statistics of feature vectors
+[…]
+[…]
+21:
+22:# Compute $v(\alpha)$ based on FID
+23:for t=0t=0toT1T-1 do
+24:  $v_{t}\leftarrow\dfrac{\mathrm{FID}(\bm{\mu}_{t},\bm{\Sigma}_{t},\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})}{\alpha_{t}-\alpha_{t+1}}$
+25:end for
+[…]
+Table 8: Hyperparameters for computing $v_{\mathrm{FID}}(\alpha)$.
+Parameter Description Used Value
+$T$
+Number of timesteps
+for simulating the forward process
+$10^{3}$
+$\{\alpha_{t}\}_{t=0,...,T}$
+Grid of $\alpha$-values
+for evaluating $v(\alpha)$
+$\alpha_{t}=1-\left(\frac{t}{T}\right)^{p},$
+where $p=\begin{cases}2,&\mathrm{pixel\ space}\\ 1,&\mathrm{latent\ space}\end{cases}$
+$\phi$
+Feature extractor
+for computing FID
+Pixel space: Inception-V3 model used
+in the EDM repository for FID evaluation
+Latent space: ResNet-50 trained by us
+on ImageNet classification in the latent space
+C.2 Discrepancy Measure Based on Data Prediction
 Algorithm 2 presents the pseudocode for computing $v_{x}(\alpha)$ using a trained diffusion model.
-This pseudocode focuses on the optimization of sampling schedules.
-Similar to the computation of $v_{\mathrm{FID}}(\alpha)$, the forward process is simulated using the training dataset, and the change in the data prediction between adjacent timesteps is computed.
-A common configuration of hyperparameters $(T=10^{3},S=10^{4},\alpha_{s}=1,\alpha_{e}=10^{-4})$ was used in our experiments.
+As in the case of $v_{\mathrm{FID}}(\alpha)$, this involves simulating the forward process on the training data and measuring changes in the predicted data across adjacent timesteps.
+Hyperparameters are summarized in Table 9.
+We use the same configuration for all our experiments, regardless of the diffusion type or dataset.
 The computation time for obtaining $v_{x}(\alpha)$ is negligible compared with the training time of the diffusion model.
@@
 Algorithm 2 $v_{x}(\alpha)$ for optimizing sampling schedule.
-0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, trained diffusion model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$, number of timesteps for simulating the forward process $T$, number of samples used for calculating the mean $S$, maximum and minimum values of $\alpha$ $(\alpha_{s},\alpha_{e})$
-0:  $LinearInterp1d(\{\alpha_{t}\}_{t=0,...,T},\{v_{t}\}_{t=0,...,T})$
-1:  # Compute $\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right]$
-2:  $\overline{D^{2}_{t}}\leftarrow 0,\ \forall{t}\in\{0,1,...,T\}$
-3:  $\Delta\alpha\leftarrow\frac{1}{T}(\alpha_{s}-\alpha_{e})$
-4:  for $n$ in $randperm(N)[:S]$ do
-5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)}$
-6:     if latent-space diffusion model then
-7:        $\bm{x}_{0}\leftarrow autoencoder(\bm{x}_{0})$
-8:     end if
-9:     $\bm{y}_{0}\leftarrow\bm{x}_{0}$
-10:
-11:     # Simulate the forward process
-12:     for $t=1$ to $T$ do
-13:        $\alpha_{t}\leftarrow\alpha_{s}-\Delta\alpha t,\ \sigma_{t}\leftarrow\sqrt{1-\alpha_{t}^{2}}$
-14:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}},\ \delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
-15:        $\bm{x}_{t}\leftarrow\beta_{t}\bm{x}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
-16:        $\bm{y}_{t}\leftarrow\frac{1}{\alpha_{t}}\left(\bm{x}_{t}-\sigma_{t}\bm{\varepsilon}_{\theta}(\bm{x}_{t},\alpha_{t})\right)$
-17:        $\overline{D^{2}_{t}}\leftarrow\overline{D^{2}_{t}}+\frac{1}{S}\left\|\bm{y}_{t}-\bm{y}_{t-1}\right\|_{2}^{2}$
-18:     end for
-19:  end for
-20:
-21:  # Compute $v(\alpha)$ using $\overline{D^{2}_{t}}$
-22:  for $t=1$ to $T$ do
-23:     $v_{t}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{t}}}$
-24:  end for
-25:  $v_{0}\leftarrow v_{1}$
-CRS-$v_{x}$ can be used for optimizing the training schedule as well.
-A pseudocode for training diffusion models with CRS is given in Algorithm 3.
-Since the data prediction is expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, $v_{x}(\alpha)$ tends to take extremely large values when $\alpha\simeq 0$.
-To avoid this, $v_{x}(\alpha)$ is evaluated within the range of $\alpha_{\mathrm{th}}\leq\alpha\leq\alpha_{\max}$, and we set $v_{x}(\alpha<\alpha_{\mathrm{th}})=v_{x}(\alpha_{\mathrm{th}})$.
-Similar to the implementation of the adaptive noise schedule in VDM++ [13], the range of $\alpha$ from $\alpha_{\mathrm{th}}$ to $\alpha_{\max}$ is divided into $B$ evenly spaced bins, and $\overline{D^{2}_{x}}(\alpha,\alpha+\Delta\alpha)$ is approximately evaluated using the exponential moving average (EMA).
-We used a common configuration of hyperparameters ($B=100,e=0.995,\alpha_{\max}=1,\alpha_{\min}=0,\alpha_{\mathrm{th}}=0.01,\Delta\alpha=10^{-3},\xi=1.0)$ in this study.
-The initial noise schedule is set to $\alpha(t)=\alpha_{\max}-(\alpha_{\max}-\alpha_{\min})\frac{t}{T}$.
+1:training dataset $\{\bm{x}^{(n)}\}_{n=0,\dots,N-1}$, trained diffusion model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$, number of timesteps $T$, number of samples for mean $S$, range of $\alpha$ $(\alpha_{s},\alpha_{e})$
+[…]
+3:# Evaluate $\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\!\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right]$
+[…]
+[…]
+6:for $n$ in $\mathrm{randperm}(N)[:S]$ do
+7:  $\bm{x}_{0}\leftarrow\bm{x}^{(n)}$
+8:  if model operates in latent space then
+9:   $\bm{x}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}_{0})$
+10:  end if
+11:  $\bm{y}_{0}\leftarrow\bm{x}_{0}$
+12:
+13:  # Simulate the forward process
+14:  for t=1t=1toTT do
+15:   $\alpha_{t}\leftarrow\alpha_{s}-\Delta\alpha\,t,\quad\sigma_{t}\leftarrow\sqrt{1-\alpha_{t}^{2}}$
+16:   $\beta_{t}\leftarrow\dfrac{\alpha_{t}}{\alpha_{t-1}},\quad\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
+17:   $\bm{x}_{t}\leftarrow\beta_{t}\bm{x}_{t-1}+\delta_{t}\,\mathcal{N}(\bm{0},\bm{I})$
+18:   $\bm{y}_{t}\leftarrow\dfrac{1}{\alpha_{t}}\!\left(\bm{x}_{t}-\sigma_{t}\,\bm{\varepsilon}_{\theta}(\bm{x}_{t},\alpha_{t})\right)$
+19:   $\overline{D^{2}_{t}}\leftarrow\overline{D^{2}_{t}}+\dfrac{1}{S}\left\|\bm{y}_{t}-\bm{y}_{t-1}\right\|_{2}^{2}$
+20:  end for
+21:end for
+22:
+23:# Compute $v(\alpha)$ using $\overline{D^{2}_{t}}$
+24:for t=1t=1toTT do
+25:  $v_{t}\leftarrow\sqrt{\dfrac{1}{\Delta\alpha}\,\overline{D^{2}_{t}}}$
+26:end for
+[…]
+Table 9: Hyperparameters used to compute $v_{x}(\alpha)$ for optimizing sampling schedule.
+Parameter Description Used Value
+$T$ Number of timesteps for simulating the forward process $10^{3}$
+$S$ Number of samples used to estimate $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ $10^{4}$
+$\alpha_{s}$ Maximum $\alpha$ used to evaluate $v(\alpha)$ 1.0
+$\alpha_{e}$ Minimum $\alpha$ used to evaluate $v(\alpha)$ 0.0
+Algorithm 3 shows how $v_{x}(\alpha)$ can be used to adaptively optimize the training schedule.
+Since the data prediction is expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, $v_{x}(\alpha)$ can become numerically unstable as $\alpha\to 0$.
+To prevent this, we restrict evaluation to the range $\alpha_{\mathrm{th}}\leq\alpha\leq\alpha_{\max}$ and fix $v_{x}(\alpha)=v_{x}(\alpha_{\mathrm{th}})$ for $\alpha<\alpha_{\mathrm{th}}$.
+Similar to VDM++ (Kingma and Gao, 2023), we divide the range of $\alpha$ into $B$ uniform bins and estimate $\overline{D^{2}_{x}}$ for each bin using an exponential moving average (EMA).
+Hyperparameters are summarized in Table 10.
+We use the same configuration for all our experiments, regardless of the diffusion type or dataset.
 Although not explicitly included in the pseudocode, the update of $\overline{D^{2}_{b}}$ starts after 1000 iterations, and the noise schedule $\alpha(t)$ is updated every 100 iterations.
-The training times of the diffusion models are listed in Table 5.
-When using CRS for optimizing training schedule, the training time increases by 20–30% compared with conventional noise schedules because CRS requires evaluating $v_{x}(\alpha)$ during training.
-To mitigate the increase in training time, one potential approach is to fix the training schedule after a certain number of epochs.
-In practice, CRS-$v_{x}$ showed little change beyond 10 epochs on LSUN Horse 256 $\times$ 256, suggesting the possibility of reducing training time while maintaining performance by fixing the training schedule after 10 epochs.
-Reducing the update frequency of $\overline{D^{2}_{x}}$ can be also an effective strategy.
-Developing an efficient implementation to reduce the training time remains as future work.
-Algorithm 3 Adaptive noise schedule for training.
-0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N}$, initial noise schedule $\alpha(t)$, number of bins $B$, decay rate of EMA $e$, maximum and minimum values of $\alpha$ for training schedule $(\alpha_{\max},\alpha_{\min})$, infinitesimal variation of $\alpha$ for evaluating the derivative $\Delta\alpha$, minimum $\alpha$ for evaluating $v(\alpha)$ $\alpha_{\mathrm{th}}$, hyperparameter of CRS $\xi$
-0:  noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$
-1:  $\alpha_{b}\leftarrow\alpha_{\mathrm{th}}+(\alpha_{\max}-\alpha_{\mathrm{th}})\frac{b}{B},\ \forall{b}\in\{0,...,B\}$
-2:  $\overline{D^{2}_{b}}\leftarrow 1e^{-6},\ \forall{b}\in\{0,...,B-1\}$
-3:  for $n$ in $randperm(N)$ do
-4:     # Compute loss for training diffusion models
-5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)},\ t\sim\mathcal{U}(0,1),\ \bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
-6:     $\alpha\leftarrow\alpha(t),\ \sigma\leftarrow\sqrt{1-\alpha^{2}}$
-7:     $\bm{\varepsilon}_{\alpha}\leftarrow\bm{\varepsilon}_{\theta}(\alpha\bm{x}_{0}+\sigma\bm{\varepsilon},\alpha)$
-8:     $loss\leftarrow\frac{1}{2}\left\|\bm{\varepsilon}_{\alpha}-\bm{\varepsilon}\right\|_{2}^{2}$
-9:
-10:     # Evaluate $\overline{D^{2}_{x}}$ using EMA (without gradient)
-11:     if $\alpha\geq\alpha_{\mathrm{th}}$ then
-12:        $\alpha^{\prime}\leftarrow\alpha-\Delta\alpha,\ \beta^{\prime}\leftarrow\frac{\alpha^{\prime}}{\alpha},\ \delta^{\prime}\leftarrow\sqrt{1-\beta^{\prime 2}}$
-13:        $\bm{x}_{\alpha^{\prime}}\leftarrow\beta^{\prime}\bm{x}_{\alpha}+\delta^{\prime}\times\mathcal{N}(\bm{0},\bm{I})$
-14:        $\bm{\varepsilon}_{\alpha^{\prime}}\leftarrow\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})$
-15:        $\bm{y}_{\alpha}\leftarrow\frac{1}{\alpha}\left(\bm{x}_{\alpha}-\sigma\bm{\varepsilon}_{\alpha}\right),\ \bm{y}_{\alpha^{\prime}}\leftarrow\frac{1}{\alpha^{\prime}}\left(\bm{x}_{\alpha^{\prime}}-\sigma^{\prime}\bm{\varepsilon}_{\alpha^{\prime}}\right)$
-16:        $b\leftarrow\min\left(\left\lfloor\frac{\alpha-\alpha_{\mathrm{th}}}{\alpha_{\max}-\alpha_{\mathrm{th}}}B\right\rfloor,B-1\right)$
-17:        $\overline{D^{2}_{b}}\leftarrow e\overline{D^{2}_{b}}+(1-e)\left\|\bm{y}_{\alpha}-\bm{y}_{\alpha^{\prime}}\right\|_{2}^{2}$
-18:        $v_{b}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{b}}}$
-19:     end if
-20:
-21:     # Update training schedule $\alpha(t)$
-22:     $\mathcal{X}\leftarrow\{\alpha_{\min},\alpha_{0},\alpha_{1},...,\alpha_{B}\}$
-23:     $\mathcal{Y}\leftarrow\{v_{0},v_{0},v_{1},...,v_{B-1},v_{B-1}\}$
-24:     $v(\alpha)\leftarrow LinearInterp1d(\mathcal{X},\mathcal{Y})$
-25:     Update $\alpha(t)$ and model parameters $\theta$
-26:  end for
-Table 5: Training time of diffusion models.
-Diffusion type Dataset GPUs Epochs Training schedule Training time
-Latent space LSUN Church A100 $\times$ 8 1000
-Linear
-VDM++
-CRS-$v_{\mathrm{FID}}$
-85 hours
-CRS-$v_{\varepsilon}$
-CRS-$v_{x}$
-105 hours
-LSUN Bedroom A100 $\times$ 8 100
-CRS-$v_{\mathrm{FID}}$
-190 hours
-CRS-$v_{\varepsilon}$
-CRS-$v_{x}$
-245 hours
-ImageNet A100 $\times$ 8 500
-Linear
-VDM++
-CRS-$v_{\mathrm{FID}}$
-430 hours
-CRS-$v_{\varepsilon}$
-CRS-$v_{x}$
-525 hours
-Pixel space LSUN Horse H100 $\times$ 8 50
-Linear
-Shifted cosine
-VDM++
-170 hours
-CRS-$v_{\varepsilon}$
-CRS-$v_{x}$
-215 hours
-LSUN Bedroom H100 $\times$ 8 50
-CRS-$v_{x}$
-325 hours
-Appendix D Hyperparameter Tuning for Combining Multiple Distance Metrics
-CRS-$v_{x}+v_{\mathrm{FID}}$ has four hyperparameters; $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$.
+While this adaptive scheme slightly increases training time (by 20–30% compared to fixed schedules; see Table 15 in Appendix E), we find that the schedule stabilizes early: for example, on LSUN Horse $\times256$, $v_{x}(\alpha)$ changes very little after 10 epochs.
+Thus, fixing the schedule after a few epochs or reducing the frequency of $\overline{D^{2}_{x}}$ updates may mitigate the cost without degrading performance.
+Developing a more efficient implementation remains an important direction for future work.
+Algorithm 3 Adaptive optimization of training schedule using $v_{x}(\alpha)$.
+1:training dataset $\{\bm{x}^{(n)}\}_{n=0,\dots,N}$, initial noise schedule $\alpha(t)$, number of bins $B$, decay rate of EMA $e$, range of $\alpha$ $(\alpha_{\max},\alpha_{\min})$, derivative step $\Delta\alpha$, minimum value for evaluating $v(\alpha)$ $\alpha_{\mathrm{th}}$, hyperparameter of CRS $\xi$
+2:noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$
+[…]
+[…]
+5:for $n$ in $\mathrm{randperm}(N)$ do
+6:  # Compute loss for training diffusion models
+7:  $\bm{x}_{0}\leftarrow\bm{x}^{(n)},\quad t\sim\mathcal{U}(0,1),\quad\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
+8:  $\alpha\leftarrow\alpha(t),\quad\sigma\leftarrow\sqrt{1-\alpha^{2}}$
+9:  $\bm{\varepsilon}_{\alpha}\leftarrow\bm{\varepsilon}_{\theta}(\alpha\bm{x}_{0}+\sigma\bm{\varepsilon},\,\alpha)$
+10:  $loss\leftarrow\dfrac{1}{2}\left\|\bm{\varepsilon}_{\alpha}-\bm{\varepsilon}\right\|_{2}^{2}$
+11:
+12:  # Evaluate $\overline{D^{2}_{b}}$ using EMA (without gradient)
+13:  if $\alpha\geq\alpha_{\mathrm{th}}$ then
+14:   $\alpha^{\prime}\leftarrow\alpha-\Delta\alpha,\quad\sigma^{\prime}\leftarrow\sqrt{1-\alpha^{\prime 2}},\quad\beta^{\prime}\leftarrow\dfrac{\alpha^{\prime}}{\alpha},\quad\delta^{\prime}\leftarrow\sqrt{1-{\beta^{\prime}}^{2}}$
+15:   $\bm{x}_{\alpha^{\prime}}\leftarrow\beta^{\prime}\bm{x}_{\alpha}+\delta^{\prime}\mathcal{N}(\bm{0},\bm{I})$
+16:   $\bm{\varepsilon}_{\alpha^{\prime}}\leftarrow\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\,\alpha^{\prime})$
+17:   $\bm{y}_{\alpha}\leftarrow\dfrac{1}{\alpha}\!\left(\bm{x}_{\alpha}-\sigma\bm{\varepsilon}_{\alpha}\right),\quad\bm{y}_{\alpha^{\prime}}\leftarrow\dfrac{1}{\alpha^{\prime}}\!\left(\bm{x}_{\alpha^{\prime}}-\sigma^{\prime}\bm{\varepsilon}_{\alpha^{\prime}}\right)$
+18:   $b\leftarrow\min\!\left(\left\lfloor\dfrac{\alpha-\alpha_{\mathrm{th}}}{\alpha_{\max}-\alpha_{\mathrm{th}}}\,B\right\rfloor,\,B-1\right)$
+19:   $\overline{D^{2}_{b}}\leftarrow e\,\overline{D^{2}_{b}}+(1-e)\left\|\bm{y}_{\alpha}-\bm{y}_{\alpha^{\prime}}\right\|_{2}^{2}$
+20:   $v_{b}\leftarrow\sqrt{\dfrac{1}{\Delta\alpha}\,\overline{D^{2}_{b}}}$
+21:  end if
+22:
+23:  # Update training schedule $\alpha(t)$
+24:  $\mathcal{X}\leftarrow\{\alpha_{\min},\alpha_{0},\alpha_{1},\dots,\alpha_{B}\}$
+25:  $\mathcal{Y}\leftarrow\{v_{0},v_{0},v_{1},\dots,v_{B-1},v_{B-1}\}$
+26:  $v(\alpha)\leftarrow\mathrm{LinearInterp1d}(\mathcal{X},\mathcal{Y})$
+27:  Update noise schedule $\alpha(t)$ by solving Eq. (9)
+28:
+29:  # Update model parameters $\theta$
+30:  loss.backward()
+31:  optimizer.step()
+32:end for
+Table 10: Hyperparameters used for optimizing the training schedule with CRS-$v_{x}$.
+Parameter Description Used Value
+$\alpha_{\min}$ Minimum value of $\alpha$ for the training schedule 0.0
+$\alpha_{\max}$ Maximum value of $\alpha$ for the training schedule 1.0
+$\Delta\alpha$ Small step used for approximating the derivative $10^{-3}$
+$\alpha_{\mathrm{th}}$ Threshold of $\alpha$ for clipping $v(\alpha)$ 0.01
+$\xi$ A parameter controlling the dependence of $\alpha(t)$ on $v(\alpha)$ 1
+$B$ Number of bins used to divide the range of $\alpha$ 100
+$e$ EMA decay rate for approximately evaluating $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ 0.995
+$\alpha(t)$ Initial noise schedule 1t1-t
+Appendix D Hyperparameter Tuning for Combining Multiple Discrepancy Measures
+CRS-$v_{x}+v_{\mathrm{FID}}$ achieves the best performance for sampling in pixel-space diffusion models.
+CRS-$v_{x}+v_{\mathrm{FID}}$ involves four hyperparameters: $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$.
+The effects of these hyperparameters on the resulting noise schedules are illustrated in Figs.
+6, 6, and 6.
+Increasing either $w_{\mathrm{FID}}$ or $\xi_{\mathrm{FID}}$ amplifies the influence of $v_{\mathrm{FID}}(\alpha)$, resulting in more timesteps being allocated to the region where $\alpha\simeq 1$ (see Figs.
+6 and 6).
+Conversely, increasing either $w_{x}$ or $\xi_{x}$ strengthens the contribution of $v_{x}(\alpha)$, leading to more timesteps being assigned to the region where $\alpha\simeq 0$ (see Figs.
+6 and 6).
+Figure 4: Effect of the weights $w_{x}$ and $w_{\mathrm{FID}}$ on the noise schedule generated by CRS-$v_{x}+v_{\mathrm{FID}}$, with fixed exponents $\xi_{x}=\xi_{\mathrm{FID}}=1$.
+Figure 5: Effect of the exponent $\xi_{\mathrm{FID}}$ on the noise schedule generated by CRS-$v_{x}+v_{\mathrm{FID}}$, with fixed weights $w_{x}=w_{\mathrm{FID}}=0.5$.
+Figure 6: Effect of the exponent $\xi_{x}$ on the noise schedule generated by CRS-$v_{x}+v_{\mathrm{FID}}$, with fixed weights $w_{x}=w_{\mathrm{FID}}=0.5$.
 We tuned these hyperparameters in two steps.
-In the first step, we fixed $\xi_{x}=\xi_{\mathrm{FID}}=1$ and searched for appropriate values of $w_{x}$ and $w_{\mathrm{FID}}$, subject to the constraint $w_{x}+w_{\mathrm{FID}}=1$.
-In the second step, we fixed $w_{x}$ and $w_{\mathrm{FID}}$, then adjusted $\xi_{x}$ and $\xi_{\mathrm{FID}}$.
-The detailed results on LSUN Horse 256 $\times$ 256 are listed in Table 6.
-In the first step, adjusting $w_{x}$ and $w_{\mathrm{FID}}$ led to significant changes in FID at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$.
-In the second step, the improvement from tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ was small at $\mathrm{NFE}=250$.
-In contrast, for $\mathrm{NFE}=50$, the tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ resulted in a substantial improvement in FID, indicating that the importance of hyperparameter tuning increases as NFE decreases.
-Table 6: Detailed results of hyperparameter tuning on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model.
+First, we fixed $\xi_{x}=\xi_{\mathrm{FID}}=1$ and searched for appropriate values of $w_{x}$ and $w_{\mathrm{FID}}$ under the constraint $w_{x}+w_{\mathrm{FID}}=1$.
+Then, we fixed the weights and adjusted $\xi_{x}$ and $\xi_{\mathrm{FID}}$.
+Although further iterations between tuning weights and exponents are possible, we found that just one iteration per step was typically sufficient to outperform conventional noise schedules.
+The detailed tuning results on LSUN Horse $\times256$ are presented in Table 11.
+In the first step, adjusting $w_{x}$ and $w_{\mathrm{FID}}$ led to substantial changes in FID at both $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$.
+In the second step, tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ had little impact at $\mathrm{NFE}=250$, but resulted in a significant FID improvement at $\mathrm{NFE}=50$.
+This suggests that hyperparameter sensitivity increases as the NFEs decrease.
+Table 11: Detailed results of hyperparameter tuning for CRS-$v_{x}+v_{\mathrm{FID}}$ on LSUN Horse $\times256$ in pixel-space diffusion model.
 DPM-Solver++(2M) was used as sampler.
-To conserve computational resources, FID score was calculated using 10K samples.
+To reduce computational cost, FID scores were calculated using 10K samples.
 NFE Tuning step $w_{\mathrm{FID}}$ $x_{x}$ $\xi_{\mathrm{FID}}$ $\xi_{x}$ FID (10K)
@@
 1.4 3.82
-The dependence of noise schedules on the hyperparameters is shown in Figs.
-7, 7, and 7.
-Increasing either $w_{\mathrm{FID}}$ or $\xi_{\mathrm{FID}}$ amplifies the impact of $v_{\mathrm{FID}}(\alpha)$, resulting in more timesteps being allocated to the region where $\alpha\simeq 1$ (see Figs.
-7 and 7).
-On the other hand, increasing either $w_{x}$ or $\xi_{x}$ strengthens the impact of $v_{x}(\alpha)$, causing more timesteps to be allocated to the region where $\alpha\simeq 0$ (see Figs.
-7 and 7).
-The hyperparameter settings for CRS-$v_{x}+v_{\mathrm{FID}}$ are listed in Table 7.
-Noise schedules $\alpha(t)$ obtained using CRS-$v_{x}+v_{\mathrm{FID}}$ are illustrated in Figs.
-13, 13, 13, 13, 13, and 13.
-Table 7: Hyperparameter settings of CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse 256 $\times$ 256, LSUN Bedroom 256 $\times$ 256, FFHQ 64 $\times$ 64, and CIFAR10 32 $\times$ 32.
+The final hyperparameter settings used in our experiments are summarized in Table 12.
+The resulting noise schedules $\alpha(t)$ for CRS-$v_{x}+v_{\mathrm{FID}}$ are visualized in Figs.
+11, 11, 11, 11, and 11.
+Table 12: Hyperparameter settings used for CRS-$v_{x}+v_{\mathrm{FID}}$ in our experiments on LSUN Horse $\times256$, LSUN Bedroom $\times256$, FFHQ $\times64$, and CIFAR10 $\times32$.
 Dataset Sampler NFE $w_{\mathrm{FID}}$ $w_{x}$ $\xi_{\mathrm{FID}}$ $\xi_{x}$
-LSUN Horse 256 $\times$ 256 SDE-DPM-Solver++(2M) 250 0.9 0.1 1.4 1.0
+LSUN Horse $\times256$ SDE-DPM-Solver++(2M) 250 0.9 0.1 1.4 1.0
 50 0.7 0.3 1.4 1.2
@@
 50 0.5 0.5 1.4 1.4
-LSUN Bedroom 256 $\times$ 256 DPM-Solver++(2M) 10 0.5 0.5 1.0 1.0
+LSUN Bedroom $\times256$ DPM-Solver++(2M) 10 0.5 0.5 1.0 1.0
 5 0.5 0.5 1.0 1.2
-UniPC 10 0.5 0.5 1.0 1.2
+UniPC 10 0.5 0.5 1.0 1.5
 5 0.3 0.7 1.0 1.0
-FFHQ 64 $\times$ 64 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.4
+FFHQ $\times64$ DPM-Solver++(2M) 20 0.5 0.5 1.7 1.4
 10 0.5 0.5 1.6 1.4
-10 0.5 0.5 1.6 1.5
-iPNDM 50 0.5 0.5 1.2 1.5
-20 0.5 0.5 1.4 1.4
-10 0.5 0.5 1.5 1.4
-CIFAR10 32 $\times$ 32 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.0
-20 0.5 0.5 1.8 2.4
-10 0.5 0.5 1.6 3.0
-iPNDM 50 0.7 0.3 1.6 1.2
+UniPC(3M) 20 0.5 0.5 1.7 1.4
 10 0.5 0.5 1.6 1.6
-10 0.5 0.5 1.6 3.2
-Figure 8: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with SDE-DPM-Solver++(2M).
-Figure 9: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with PNDM.
-Figure 10: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with DPM-Solver++(2M).
-Figure 11: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with DPM-Solver++(2M).
-Figure 12: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with UniPC.
-Figure 13: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating FFHQ 64 $\times$ 64 with DPM-Solver++(2M).
+iPNDM(3M) 20 0.5 0.5 1.6 1.2
+10 0.5 0.5 1.6 1.2
+iPNDM(4M) 20 0.5 0.5 1.4 1.4
+10 0.5 0.5 1.6 1.4
+CIFAR10 $\times32$ DPM-Solver++(2M) 20 0.5 0.5 1.8 1.2
+10 0.5 0.5 1.7 1.2
+UniPC(3M) 20 0.5 0.5 2.3 1.0
+10 0.5 0.5 1.9 1.5
+iPNDM(3M) 20 0.5 0.5 1.7 1.4
+10 0.5 0.5 1.7 1.4
+iPNDM(4M) 20 0.5 0.5 1.6 1.6
+10 0.5 0.5 1.6 1.6
+Figure 7: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse $\times256$, used with SDE-DPM-Solver++(2M) during sampling.
+Figure 8: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse $\times256$, used with PNDM during sampling.
+Figure 9: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse $\times256$, used with DPM-Solver++(2M) during sampling.
+Figure 10: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Bedroom $\times256$, used with DPM-Solver++(2M) during sampling.
+Figure 11: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Bedroom $\times256$, used with UniPC during sampling.
 Appendix E Implementation Details
-E.1 Unet Model
+E.1 U-Net Architecture
 The hyperparameters of the U-Net model are listed in Table 13.
-For latent-space diffusion models, we used the same values as the ImageNet 64 $\times$ 64 in ADM [52].
-For pixel-space diffusion models, we used the same values as LSUN 256 $\times$ 256 in ADM [52].
+For latent-space diffusion models, we use the same settings as those for ImageNet $\times64$ in ADM (Dhariwal and Nichol, 2021).
+For pixel-space diffusion models, we adopt the same settings as those for LSUN $\times256$ in ADM.
 Table 13: Hyperparameters of UNet model.
@@
 Dropout 0.1 0.1
-E.2 Hyperparameters for Training
-The hyperparameters for training the diffusion models are summarized in Table 9.
-We saved ten checkpoints for each model.
-The FID score with 10K samples was evaluated for all checkpoints, and the results are reported with the checkpoint with the best FID score.
+E.2 Training Hyperparameters
+The training hyperparameters are summarized in Table 14.
+We save ten checkpoints for each model.
+FID scores are computed using 10K samples for all checkpoints, and we report the results from the checkpoint with the best FID.
 Table 14: Hyperparameters for training diffusion models. Same configuration was used for training latent-space and pixel-space diffusion models.
@@
 Batch size 32 $\times$ 8
-E.3 Noise Schedule for Training and Sampling
+The training times for each model configuration are provided in Table 15.
+When using CRS to optimize the training schedule, training time increases by 20–30% compared to conventional noise schedules, due to the overhead of computing $v_{x}(\alpha)$ during training.
+In practice, however, we observe that the noise schedule stabilizes early; for example, on LSUN Horse $\times256$, $v_{x}(\alpha)$ shows minimal change beyond 10 epochs.
+This suggests that freezing the schedule after a small number of epochs or reducing the frequency of $\overline{D_{x}^{2}}$ updates can mitigate training overhead without degrading performance.
+Developing a more efficient implementation remains an important direction for future work.
+Table 15: Training time of diffusion models.
+Diffusion type Dataset GPUs Epochs Training schedule Training time
+Latent space LSUN Church A100 $\times$ 8 1000
+Linear
+VDM++
+CRS-$v_{\mathrm{FID}}$
+85 hours
+CRS-$v_{\varepsilon}$
+CRS-$v_{x}$
+105 hours
+LSUN Bedroom A100 $\times$ 8 100
+CRS-$v_{\mathrm{FID}}$
+190 hours
+CRS-$v_{\varepsilon}$
+CRS-$v_{x}$
+245 hours
+ImageNet A100 $\times$ 8 500
+Linear
+VDM++
+CRS-$v_{\mathrm{FID}}$
+430 hours
+CRS-$v_{\varepsilon}$
+CRS-$v_{x}$
+525 hours
+Pixel space LSUN Horse H100 $\times$ 8 50
+Linear
+Shifted cosine
+VDM++
+170 hours
+CRS-$v_{\varepsilon}$
+CRS-$v_{x}$
+215 hours
+LSUN Bedroom H100 $\times$ 8 50
+CRS-$v_{x}$
+325 hours
+E.3 Noise Schedules for Training and Sampling
 The conventional noise schedules used in this study are summarized in Table 16.
-When using a CRS-based training schedule, a common hyperparameter setting of $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.0)$ is used (details are described in Appendix C.2).
-When using a CRS-based sampling schedule, the default setting is $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.01)$.
-An exception was made when using CRS-$v_{x}+v_{\mathrm{FID}}$, where the hyperparameters are tuned according to the procedure described in Appendix D.
-For the experiments using pretrained EDMs, we used $(\alpha_{\max},\alpha_{\min})=(0.999998,0.0125)$, which is equivalent to the boundary condition of EDM: $(\sigma_{\min},\sigma_{\max})=(0.002,80)$.
-The boundary condition of EDM is transformed to $\alpha_{\max}$ and $\alpha_{\min}$ using Eq.
-(16).
 Table 16: Conventional noise schedules evaluated in this study. $T_{s}$ denotes number of timesteps for sampling.
@@
 $T=1000,\tilde{\beta}_{\min}=10^{-4},\tilde{\beta}_{\max}=0.02.$
-Sampling $\left\{\alpha_{\mathrm{linear}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\}$
+[…]
 Shifted cosine Training
@@
 Sampling
-$\left\{\tilde{\alpha}_{\mathrm{shifted}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\},$
+[…]
 $\tilde{\alpha}_{\mathrm{shifted}}(t)=\alpha_{\min}+(\alpha_{\max}-\alpha_{\min})\alpha_{\mathrm{shifted}}(t),$
@@
 $p(\lambda)\propto\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\omega(\lambda)\left\|\bm{\varepsilon}_{\theta}(\bm{x},\tilde{\alpha}(\lambda))-\bm{\varepsilon}\right\|_{2}^{2}\right],$
-[…]
+$\omega(\lambda)=\begin{cases}\underset{\lambda}{\operatorname{max}}\ \tilde{\omega}(\lambda),&\lambda<\underset{\lambda}{\operatorname{argmax}}\ \tilde{\omega}(\lambda)\\ \tilde{\omega}(\lambda),&\lambda\geq\underset{\lambda}{\operatorname{argmax}}\ \tilde{\omega}(\lambda),\end{cases},$
 $\tilde{\omega}(\lambda)=\mathcal{N}(\lambda;\lambda_{\mu},\lambda_{\sigma}^{2})(e^{-\lambda}+c^{2}),$
@@
 $\sigma_{\max}=80,\sigma_{\min}=0.002,\rho=7.$
-Appendix F Detailed Results
-We present the detailed experimental results and confirm the superiority of CRS for optimizing both training and sampling schedules.
+Appendix F Detailed Experimental Results
+We provide detailed experimental results to confirm the effectiveness of CRS for optimizing both training and sampling schedules.
 F.1 Results on Pixel-Space Diffusion Models
-LSUN Horse 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=250$ are listed in Table 11.
-To verify the generality with respect to samplers, we used three different samplers: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M).
-As shown in the rows shaded in gray, multiple evaluation metrics, including FID score, are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule, regardless of the sampler.
-Table 12 presents the dependence on the sampling schedules at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$.
-When using SDE-DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ improved upon all four evaluation metrics at $\mathrm{NFE}=250$, achieving a new state-of-the-art FID score.
-When using PNDM and DPM-Solver++(2M), there was little difference between sampling schedules at $\mathrm{NFE}=250$.
-However, at $\mathrm{NFE}=50$, the performance became more sensitive to the sampling schedule, and CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the best FID score.
-LSUN Bedroom 256 $\times$ 256: Since only the FID score is reported in the main part of this study, we include sFID, precision, and recall in Table 13 for reference.
+LSUN Horse $\times256$: The results showing the dependence on training and sampling schedules at $\mathrm{NFE}=250$ are presented in Table 17.
+To assess the generality across samplers, we evaluated three different ones: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M).
+As shown in the gray-shaded rows, using CRS-$v_{x}$ for training consistently improved multiple evaluation metrics, including FID, regardless of the sampler used.
+Figure 13 shows the evolution of the FID score with 10k samples during training, evaluated using DPM-Solver++ (2M) with $\mathrm{NFE}=250$.
+In this experiment, we compare CRS-$v_{x}$ with shifted cosine, which was the strongest-performing baseline training schedule.
+The results demonstrate that using CRS-$v_{x}$ as the training schedule not only improves the final FID score but also leads to faster convergence throughout training.
+Table 18 presents the effect of different sampling schedules at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$.
+When using SDE-DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ improved all four metrics at $\mathrm{NFE}=250$, achieving a new state-of-the-art FID score.
+With PNDM and DPM-Solver++(2M), differences among sampling schedules were minor at $\mathrm{NFE}=250$; however, at $\mathrm{NFE}=50$, performance became more sensitive to the sampling schedule.
+In this low-NFE regime, CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the best FID score.
+LSUN Bedroom $\times256$: Since only FID scores are reported in the main text (Table 2), we include sFID, precision, and recall in Table 19 for reference.
+Table 20 compares the FID scores of CRS-$v_{x}+v_{\mathrm{FID}}$ and EDM sampling under various values of the hyperparameter $\rho$.
+While the default setting for EDM uses $\rho=7$, tuning $\rho$ can further improve its performance.
+However, even with the optimal $\rho$, EDM sampling did not outperform CRS-$v_{x}+v_{\mathrm{FID}}$ in terms of FID.
+Figure 12: Evolution of the FID score with 10k samples during training on LSUN Horse $\times256$ in the pixel-space diffusion model.
+Using CRS-$v_{x}$ as the training schedule not only improves the final FID score but also leads to faster convergence compared to shifted cosine.
+Evaluation is performed using DPM-Solver++(2M) with $\mathrm{NFE}=250$.
+Figure 13: Evolution of the FID score with 10k samples during training on LSUN Church $\times256$ in the latent-space diffusion model.
+CRS-$v_{x}$ achieves a slightly better best FID score and consistently lower FID scores before reaching the best value compared to the linear schedule.
+Evaluation is performed using DPM-Solver++(2M) with $\mathrm{NFE}=30$.
 F.2 Results on Latent-Space Diffusion Models
-LSUN Church 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 14.
-To verify the generality with respect to samplers, we used five different samplers: Stochastic DDIM, SDE-DPM-Solver++(2M), DDIM, PNDM, and DPM-Solver++(2M).
-As shown in the rows shaded in gray, FID scores are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule across all samplers except for SDE-DPM-Solver++(2M).
-The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 15 and 16, respectively.
-CRS-$v_{x}$ or CRS-$v_{\varepsilon}$ achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
-LSUN Bedroom 256 $\times$ 256: The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 17 and 18, respectively.
-CRS-$v_{x}$ is used for optimizing the training schedule.
-CRS achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=30$ and $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
-FID scores evaluated at small NFEs are listed in Table 19.
-Compared with the results reported in DPM-Solver-v3 [26], the better FID scores are achieved by using CRS for optimizing both the training and sampling schedules at $\mathrm{NFE}=10$ and $\mathrm{NFE}=20$.
-FID, sFID, precision, and recall are listed in Table 20.
-CRS based noise schedules consistently outperform the linear schedule.
-ImageNet 256 $\times$ 256: We demonstrated the superiority of CRS in class-conditional image generation.
-The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 21.
-As shown in the rows shaded in gray, CRS-$v_{\mathrm{FID}}$ was the best for optimizing training schedule for ImageNet, while CRS-$v_{x}$, which consistently showed strong performance on other datasets, was the second-best.
-The reason CRS-$v_{\mathrm{FID}}$ performed better on ImageNet may be attributed to the fact that the feature model used to compute FID is trained on ImageNet classification.
-The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Tables 22 and 23, respectively.
-We used CRS-$v_{x}$ for optimizing training schedule, which outperformed conventional noise schedules with significant margin on ImageNet and exhibited the best performance across a wide range of datasets.
-CRS achieved the best FID score in all cases except when using PNDM at $\mathrm{NFE}=20$ and DPM-Solver++(2M) at $\mathrm{NFE}=30$, and multiple evaluation metrics are simultaneously improved in most cases.
-F.3 Results on Pretrained EDM
-FFHQ 64 $\times$ 64: The evaluation results are listed in Table 24.
-When using DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ outperforms both EDM and AYS.
-Using iPNDM further improves FID scores, and CRS-$v_{x}+v_{\mathrm{FID}}$ consistently performs better than EDM.
-CIFAR10 32 $\times$ 32: The evaluation results on class-conditional image generation are listed in Table 25.
-CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperforms EDM, and significant performance gains are observed at small NFEs.
+LSUN Church $\times256$: The results showing the dependence on training and sampling schedules at $\mathrm{NFE}=30$, using stochastic and deterministic samplers, are summarized in Tables 21 and 22, respectively.
+To evaluate generality across samplers, we tested five different ones: Stochastic DDIM, SDE-DPM-Solver++(2M), DDIM, PNDM, and DPM-Solver++(2M).
+As shown in the gray-shaded rows, CRS-$v_{x}$ consistently improved FID scores across all samplers except for SDE-DPM-Solver++(2M).
+Figure 13 shows the evolution of the FID score with 10k samples during training, evaluated using DPM-Solver++ (2M) with $\mathrm{NFE}=30$.
+In this experiment, we compare CRS-$v_{x}$ with linear, which was the strongest-performing baseline training schedule.
+Both training schedules reach their best FID scores at around 400 epochs, with CRS-$v_{x}$ achieving a slightly better best FID score.
+Furthermore, CRS-$v_{x}$ also showed consistently lower FID scores before reaching the best value, indicating its effectiveness in terms of convergence speed as well.
+The effects of different sampling schedules are shown separately for stochastic and deterministic samplers in Tables 23 and 24, respectively.
+CRS-$v_{x}$ or CRS-$v_{\varepsilon}$ achieved the best FID scores in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=20$.
+In most cases, multiple evaluation metrics—including sFID, precision, and recall—were simultaneously improved.
+LSUN Bedroom $\times256$: The results showing the dependence on sampling schedules using stochastic and deterministic samplers are presented in Tables 25 and 26, respectively.
+The training schedule was fixed to $v_{x}.
+CRS$ achieved the best FID scores in all cases, except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=30$ and $\mathrm{NFE}=20$.
+In most cases, multiple evaluation metrics were simultaneously improved.
+FID scores evaluated at small NFE settings are listed in Table 27.
+Compared to the results reported in DPM-Solver-v3 (Zheng et al., 2023b), better FID scores are consistently achieved by using CRS to optimize both training and sampling schedules across all NFE values.
+As with other evaluations in this paper, we set $\xi=1$ by default for CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, except at $\mathrm{NFE}=5$, where $\xi$ was manually tuned to further improve performance.
+In this experiment, we observed that the importance of tuning $\xi$ increases as the NFE decreases.
+We attribute this behavior to the discretization error introduced by the continuous-time approximation used in Eq.
+(7).
+A comprehensive set of evaluation metrics, including FID, sFID, precision, and recall, is reported in Table 28.
+CRS-based noise schedules consistently outperform the linear schedule across all metrics, except for recall at $\mathrm{NFE}=20$.
+ImageNet $\times256$: We demonstrated the effectiveness of CRS in class-conditional image generation on ImageNet.
+It is important to note that our experiments do not use classifier-free guidance (CFG), and therefore the FID scores are naturally higher than those obtained under CFG-enabled settings.
+However, as shown in Table 29, our model achieves better FID score than the LDM-4 results reported in the original LDM paper (Rombach et al., 2022).
+These observations indicate that our ImageNet results represent a reasonable level of performance for class-conditional generation without CFG, and provide a meaningful basis for evaluating the impact of CRS in this setting.
+The results showing the dependence on training and sampling schedules at $\mathrm{NFE}=30$, using both stochastic and deterministic samplers, are summarized in Tables 30 and 31, respectively.
+As shown in the gray-shaded rows, CRS-$v_{\mathrm{FID}}$ achieved the best performance for training schedule optimization, while CRS-$v_{x}$, which consistently performed well on other datasets, was the second-best.
+This result may be attributed to the fact that the feature extractor used in FID computation is pre-trained on ImageNet, which may advantage CRS-$v_{\mathrm{FID}}$ in this specific setting.
+The impact of different sampling schedules using stochastic and deterministic samplers is shown in Tables 32 and 33, respectively.
+Here, the training schedule was fixed to CRS-$v_{x}$, which outperformed conventional noise schedules by a significant margin on ImageNet and delivered strong results across a wide range of datasets.
+CRS achieved the best FID scores in all cases, except when using PNDM at $\mathrm{NFE}=20$ and DPM-Solver++(2M) at $\mathrm{NFE}=30$.
+In most settings, CRS-based schedules also led to improvements in other evaluation metrics, including sFID, precision, and recall.
 Appendix G Generated Samples
-Generated images are shown in the following figures:
-  • LSUN Horse 256 $\times$ 256 in the pixel-space diffusion model: Figs. 14 and 15.
-  • LSUN Bedroom 256 $\times$ 256 in the pixel-space diffusion model: Figs. 16 and 17.
-  • LSUN Church 256 $\times$ 256 in the latent-space diffusion model: Fig. 18.
-  • LSUN Bedroom 256 $\times$ 256 in the latent-space diffusion model: Fig. 19.
-  • ImageNet 256 $\times$ 256 in the latent-space diffusion model: Fig. 20.
-All models are trained using CRS-$v_{x}$ for optimizing training schedule.
-Table 11: Dependence on training and sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models at $\mathrm{NFE}=250$.
-To verify generality with respect to samplers, we used three different samplers.
-Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules.
-Bold values indicate best training or sampling schedules for each metric.
-Note that "cosine" refers to shifted cosine in this table.
+Generated samples are shown in the following figures:
+  • LSUN Horse $\times256$ in the pixel-space diffusion model: Figs. 14 and 15.
+  • LSUN Bedroom $\times256$ in the pixel-space diffusion model: Figs. 16 and 17.
+  • LSUN Church $\times256$ in the latent-space diffusion model: Fig. 18.
+  • LSUN Bedroom $\times256$ in the latent-space diffusion model: Fig. 19.
+  • ImageNet $\times256$ in the latent-space diffusion model: Fig. 20.
+All models are trained using CRS-$v_{x}$ for the training schedule optimization.
+Table 17: Performance of pixel-space diffusion models on LSUN Horse $\times256$ at $\mathrm{NFE}=250$, evaluated under various training and sampling schedules.
+To assess generality across samplers, we used three different samplers: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M).
+Gray-shaded rows indicate results for different training schedules (with fixed sampling), while unshaded rows show results for different sampling schedules (with fixed training).
+Bold values indicate the best performance for each metric.
+Note that "cosine" refers to the shifted cosine schedule.
 Training Sampling Metrics
@@
 CRS-$v_{x}+v_{\mathrm{FID}}$ 2.19 5.56 0.67 0.57
-Table 12: Dependence on sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models.
-CRS-$v_{x}$ was used for optimizing training schedule.
-Bold values indicate best sampling schedules for each metric.
+Table 18: Performance of pixel-space diffusion models on LSUN Horse $\times256$ under different sampling schedules.
+Training schedule was fixed to CRS-$v_{x}$.
+Bold values indicate the best sampling schedule for each evaluation metric.
 Sampling Metrics
@@
 CRS-$v_{x}+v_{\mathrm{FID}}$ 2.26 5.90 0.66 0.57
-Table 13: Evaluation results on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion model. CRS-$v_{x}$ was used for optimizing training schedule.
+Table 19: Evaluation results on LSUN Bedroom $\times256$ in pixel-space diffusion model. CRS-$v_{x}$ was used for optimizing training schedule.
 Sampling Metrics
@@
 CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 11.26 0.21 0.44
-Table 14: Dependence on training and sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$.
-To verify generality with respect to samplers, we used five different samplers.
-Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules.
-Bold values indicate best training or sampling schedules for each metric.
+Table 20: Comparison of FID scores for CRS-$v_{x}+v_{\mathrm{FID}}$ and EDM sampling on LSUN Bedroom $\times256$ in the pixel-space diffusion model.
+Performance of EDM is reported under various values of the hyperparameter $\rho$.
+Even with the optimal $\rho$, EDM does not outperform CRS-$v_{x}+v_{\mathrm{FID}}$.
+Sampler Sampling schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$
+DPM-Solver++(2M) EDM: $\rho=1$ 190.32 102.10
+EDM: $\rho=3$ 30.24 13.12
+EDM: $\rho=5$ 22.97 7.92
+EDM: $\rho=7$ 23.72 4.73
+CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 4.88
+UniPC EDM: $\rho=1$ 178.24 90.30
+EDM: $\rho=3$ 23.45 4.43
+EDM: $\rho=5$ 55.06 3.96
+EDM: $\rho=7$ 82.52 4.94
+CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 3.30
+Table 21: Performance of latent-space diffusion models on LSUN Church $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using stochastic samplers.
+Two samplers were used: Stochastic DDIM and SDE-DPM-Solver++(2M).
+Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training).
+Bold values indicate the best result for each evaluation metric.
 Training Sampling Metrics
@@
 CRS-$v_{x}$ 3.20 9.14 0.62 0.58
+Table 22: Performance of latent-space diffusion models on LSUN Church $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using deterministic samplers.
+Three samplers were used: DDIM, PNDM, and DPM-Solver++(2M).
+Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training).
+Bold values indicate the best result for each evaluation metric.
+Training Sampling Metrics
+Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
 DDIM Linear Linear 6.18 13.16 0.59 0.51
@@
 CRS-$v_{x}$ 3.59 9.51 0.59 0.60
-Table 15: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with stochastic samplers.
-CRS-$v_{x}$ was used for optimizing training schedule.
-Bold values indicate best sampling schedules for each metric.
+Table 23: Performance of latent-space diffusion models on LSUN Church $\times256$ using stochastic samplers, evaluated under different sampling schedules.
+The training schedule was fixed to CRS-$v_{x}$.
+Bold values indicate the best sampling schedule for each evaluation metric.
 Sampling Metrics
@@
 CRS-$v_{x}$ 3.57 9.55 0.62 0.56
-Table 16: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with deterministic samplers.
-CRS-$v_{x}$ was used for optimizing training schedule.
-Bold values indicate best sampling schedules for each metric.
+Table 24: Performance of latent-space diffusion models on LSUN Church $\times256$ using deterministic samplers, evaluated under different sampling schedules.
+The training schedule was fixed to CRS-$v_{x}$.
+Bold values indicate the best sampling schedule for each evaluation metric.
 Sampling Metrics
@@
 CRS-$v_{x}$ 4.00 9.44 0.58 0.59
-Table 17: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with stochastic samplers.
-CRS-$v_{x}$ was used for optimizing training schedule.
-Bold values indicate best sampling schedules for each metric.
+Table 25: Performance of latent-space diffusion models on LSUN Bedroom $\times256$ using stochastic samplers, evaluated under different sampling schedules.
+The training schedule was fixed to CRS-$v_{x}$.
+Bold values indicate the best sampling schedule for each evaluation metric.
 Sampling Metrics
@@
 CRS-$v_{x}$ 3.07 6.34 0.56 0.52
-Table 18: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with deterministic samplers.
-CRS-$v_{x}$ was used for optimizing training schedule.
-Bold values indicate best sampling schedules for each metric.
+Table 26: Performance of latent-space diffusion models on LSUN Bedroom $\times256$ using deterministic samplers, evaluated under different sampling schedules.
+The training schedule was fixed to CRS-$v_{x}$.
+Bold values indicate the best sampling schedule for each evaluation metric.
 Sampling Metrics
@@
 CRS-$v_{x}$ 3.00 5.63 0.54 0.55
-Table 19: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs.
-Bold values indicate best FID scores at each NFE.
-Results for rows other than "Ours" are taken from previously published work [26].
+Table 27: FID scores on LSUN Bedroom $\times256$ in latent-space diffusion models at small NFE settings.
+Bold values indicate the best FID score at each NFE.
+Results for rows other than "Ours" are taken from Zheng et al.
+(2023b).
+As with other evaluations in this paper, we set $\xi=1$ by default for CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, except at $\mathrm{NFE}=5$, where $\xi$ was manually tuned to further improve performance.
+In this experiment, we observed that the importance of tuning $\xi$ increases as the NFE decreases.
+We attribute this behavior to the discretization error introduced by the continuous-time approximation used in Eq.
+(7).
 Sampling FID $\downarrow$
 Model Sampler schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$ $\mathrm{NFE}=20$
-LDM [57] DPM-Solver++ Linear 18.59 3.63 3.16
+LDM DPM-Solver++ Linear 18.59 3.63 3.16
 UniPC 12.24 3.56 3.07
-DPM-Solver-v3 [26] 7.54 3.16 3.05
+DPM-Solver-v3 Zheng et al. (2023b) 7.54 3.16 3.05
 Ours iPNDM Linear 10.81 3.71 2.64
@@
 CRS-$v_{x}$ 8.07 3.45 2.50
-Table 20: Detailed results evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs. Bold values indicate best sampling schedules for each metric.
+CRS-$v_{\varepsilon}$: $\xi=1.4$ 7.64
+CRS-$v_{x}$: $\xi=2.2$ 6.96
+Table 28: Detailed evaluation results on LSUN Bedroom $\times256$ in latent-space diffusion models at small NFE settings. Bold values indicate the best sampling schedule for each metric.
 Sampling Metrics
@@
 CRS-$v_{x}$ 8.07 10.43 0.45 0.48
-Table 21: Dependence on training and sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$.
-To verify generality with respect to samplers, we used five different samplers.
-Rows shaded in gray represent dependence on training schedule, while unshaded rows represent dependence on sampling schedules.
-Bold values indicate best training or sampling schedules for each metric.
+CRS-$v_{\varepsilon}$: $\xi=1.4$ 7.64 9.23 0.45 0.52
+CRS-$v_{x}$: $\xi=2.2$ 6.96 8.65 0.45 0.49
+Table 29: Comparison between our trained model and LDM-4 (Rombach et al., 2022) on ImageNet under class-conditional generation without CFG.
+Nparams denotes the number of parameters in the U-Net model.
+Our model, trained with CRS-$v_{x}$ as the training schedule, achieves a better FID score than LDM-4 despite having fewer U-Net parameters.
+Model Training schedule Sampling schedule Sampler NFE FID($\downarrow$) Nparams
+LDM-4 Linear Linear Stochastic DDIM 250 10.56 400M
+Ours CRS-$v_{x}$ 9.67 296M
+Table 30: Performance of latent-space diffusion models on ImageNet $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using stochastic samplers.
+Two samplers were used: Stochastic DDIM and SDE-DPM-Solver++(2M).
+Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training).
+Bold values indicate the best result for each evaluation metric.
 Training Sampling Metrics
@@
 CRS-$v_{x}$ 10.56 5.32 0.63 0.63
+Table 31: Performance of latent-space diffusion models on ImageNet $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using deterministic samplers.
+Three samplers were used: DDIM, PNDM, and DPM-Solver++(2M).
+Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training).
+Bold values indicate the best result for each evaluation metric.
+Training Sampling Metrics
+Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
 DDIM Linear Linear 17.33 6.79 0.55 0.62
@@
 CRS-$v_{x}$ 13.33 5.13 0.59 0.65
-Table 22: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with stochastic samplers.
-CRS-$v_{x}$ was used for optimizing training schedule.
-Bold values indicate best sampling schedules for each metric.
+Table 32: Performance of latent-space diffusion models on ImageNet $\times256$ using stochastic samplers, evaluated under different sampling schedules.
+The training schedule was fixed to CRS-$v_{x}$.
+Bold values indicate the best sampling schedule for each evaluation metric.
 Sampling Metrics
@@
 CRS-$v_{x}$ 11.64 5.26 0.63 0.62
-Table 23: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with deterministic samplers.
-CRS-$v_{x}$ was used for optimizing training schedule.
-Bold values indicate best sampling schedules for each metric.
+Table 33: Performance of latent-space diffusion models on ImageNet $\times256$ using deterministic samplers, evaluated under different sampling schedules.
+The training schedule was fixed to CRS-$v_{x}$.
+Bold values indicate the best sampling schedule for each evaluation metric.
 Sampling Metrics
@@
 CRS-$v_{x}$ 14.16 5.09 0.59 0.64
-Table 24: Evaluation results on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM.
-Bold values indicate best sampling schedules for each combination of sampler and NFE.
-FID scores on AYS is taken from their original paper [16].
-Sampling FID $\downarrow$
-Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
-DPM-Solver++(2M) EDM 5.57 3.02 2.58
-AYS 5.43 3.29 2.62
-CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
-iPNDM EDM 4.60 2.62 2.49
-CRS-$v_{x}+v_{\mathrm{FID}}$ 3.87 2.53 2.44
-Table 25: Evaluation results for CIFAR10 32 $\times$ 32 on class-conditional image generation using the pretrained model provided in EDM.
-Bold values indicate best sampling schedules for each combination of sampler and NFE.
-Sampling FID $\downarrow$
-Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
-DPM-Solver++(2M) EDM 4.12 2.20 1.94
-CRS-$v_{x}+v_{\mathrm{FID}}$ 3.03 2.07 1.89
-iPNDM EDM 2.51 1.93 1.89
-CRS-$v_{x}+v_{\mathrm{FID}}$ 2.33 1.92 1.89
-Figure 14: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=250$ ($\mathrm{FID}=2.03$).
-We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
-SDE-DPM-Solver++(2M) was used as sampler.
-Figure 15: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=50$ ($\mathrm{FID}=2.04$).
-We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
-PNDM was used as sampler.
-Figure 16: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=10$ ($\mathrm{FID}=3.30$).
-We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
-UniPC was used as sampler.
-Figure 17: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=5$ ($\mathrm{FID}=14.02$).
-We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
-DPM-Solver++(2M) was used as sampler.
-Figure 18: Samples of LSUN Church 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=3.89$).
-We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively.
-DPM-Solver++(2M) was used as sampler.
-Figure 19: Samples of LSUN Bedroom 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=2.71$).
-We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively.
-PNDM was used as sampler.
-Figure 20: Samples of ImageNet 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ with class condition ($\mathrm{FID}=10.83$).
+Figure 14: Generated samples of LSUN Horse $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=250$ ($\mathrm{FID}=2.03$).
+The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
+SDE-DPM-Solver++(2M) was used as the sampler.
+Figure 15: Generated samples of LSUN Horse $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=50$ ($\mathrm{FID}=2.04$).
+The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
+PNDM was used as the sampler.
+Figure 16: Generated samples of LSUN Bedroom $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=10$ ($\mathrm{FID}=3.30$).
+The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
+UniPC was used as the sampler.
+Figure 17: Generated samples of LSUN Bedroom $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=5$ ($\mathrm{FID}=14.02$).
+The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively.
+DPM-Solver++(2M) was used as the sampler.
+Figure 18: Generated samples of LSUN Church $\times256$ using a latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=3.89$).
+The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively.
+DPM-Solver++(2M) was used as the sampler.
+Figure 19: Generated samples of LSUN Bedroom $\times256$ using a latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=2.71$).
+The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively.
+PNDM was used as the sampler.
+Figure 20: Generated samples of ImageNet $\times256$ using a latent-space diffusion model at $\mathrm{NFE}=20$ with class condition ($\mathrm{FID}=10.83$).
 Six classes were randomly selected, and each row corresponds to one class.
-We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively.
-SDE-DPM-Solver++(2M) was used as sampler.
+The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively.
+SDE-DPM-Solver++(2M) was used as the sampler.